A Bibliography of Publications of Sir Ronald Aylmer Fisher

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

18 October 2022 Version 1.06

Abstract This bibliography records publications of John von Neumann (1903–

1957).

Title word cross-reference

2 × 2 [Fis45e]. 6 × 6 [FY34]. 9 × 9 [Fis41c]. W̄ [Edw90a]. χ^{2} [Fis22f, Fis24b, Fis28h]. d [Fis41a]. F [Aro40, Aro41]. g [Qui21]. n^{−1} [Fis25b]. P [Fis22f]. r × c [CaFJ93, MP86a, MP86b]. S [FY44]. z [Aro40, Aro41].

-distribution [Aro40]. -fastperiodische [FY44]. -strongly [FY44].

1 [BFP51, PFSB25]. 114 [Fis55a]. 11th [Tuk52]. 12th [Edw93]. 16 [Fis31e]. 18 [Fis57b]. 1918/19 [NP76]. 1924 [Hea73]. 1932 [Fis32g]. 1945 [FdB47]. 1950 [Fis50b]. 1951 [Fis52b]. 1962 [Ano62]. 1996 [Efr98].

2 [PBFN28, PBFN31]. 2012 [Ano12]. 21st [Efr98, Efr00, RS00].

3 [FBC39, PFSB25, PBFN28].

4 [BF43, PBFN31]. 40.00 [Sen91]. 480 [Fis35a].

50th [Ano12]. 52 [NP76].

643 [CaFJ93].

8 [Fis32g]. 8th [RS00].

91 [Fis51a].

abnormalities [Fis35f, Fis35g]. above [Fis22h, Fis26h]. Absolute [Fis12b]. Abundance [Fis43f, FF28, Fis37d]. academic [Fis35j]. acclimization [Fis38i]. Accuracy [Fis20b, FTM22]. Aclasis [PFSB25]. Acquired [Fis32d]. Activities [Cor64]. Actuarial [Fis27a]. Adaptation [Fis34a]. adaptations [Fis33f]. adapted [Fis47d]. adaptées [Fis47d]. addendum [Fis47a]. Address [Fis38d]. Adelaide [Edw93, May14]. advance [Fis37e]. advanced [Fis43a]. Advantage [GV99]. advantageous [Fis37e]. advocacy [Hal07]. against [Fis33e]. ageing [FM57]. Agouti} [Fis49d]. Agreed [Ber03]. agreement [Fis23d]. Agricultural [B.58, Fis35d, Fis36b, FY38, FY43, FY48, FY49, FY53a, FY53b, FY57a, FY57b, FY63, K.39, Tip53, Toc58, Wet65, Fis51a, Str90]. Agriculture [Fis20c]. ailments [Fis35m]. algebraically [Fis59e]. Algorithm [CaFJ93, MP86a, MP86b, CaFJ93]. Alleles [Fis47e, Fis61a, Fis62f]. Allelomorphs [Fis32f, Fis46b, Fis47a]. Allied [BF43]. Allowance [Fis43b]. allowances [Ayl21, Fis27b, Fis31a, Fis32c, Fis36d, Fis43c]. alternatives [Fis43e]. altitude [Fis38i]. American [Fis28c, TF27]. Amiable [Har08]. Amongst [Fis30i]. Amount [PL12, Fis34b]. Analysis [Bar92, Bas80a, Bas80b, Bin89, FO24, Fis29h, Fis38a, Fis47b, Fis54d, Fis54e, Hin80, Kem93, Kem80, Lan80b, Lin80, Rao64, Rub80, And96, Ben90, DF37, Fis13, Fis21b, FG37, Fis40d, Fis55a, GGA11, Sen91]. anecdote [Bar95]. animals [Fis33f]. anniversary [Ano12]. Anomalies [BFP51, PBFN28, PBFN31]. Answer [Fis51a, Fis55a, MF48]. Answers [MF48]. Anthropoid [FFH39]. Anti [Fis45a]. Anti-Hr [Fis45a]. antigens [Fis36c]. Apes [FFH39]. Aphis [DF22]. Apparent [Fis43f]. Appendix [Fis28a, Fis34c, Ano15, SYC^{+}17]. Application [BWF33]. Applications [Fis13, Fis25a, Fis31f, GGA11, RS00, Hal90]. applied [Fis36g]. Appreciation [FH80, Mat51, Owe62, Yat82]. approach [Fis41a]. Approximation [Tsu02]. archippus [Fis34n]. Architect [Rao64]. Archive [Ano13]. Argument [Sei92b, Zab92, Fis35k, Rya80]. arithmetic [FD48]. Arrangement [FW30, Fis92a, Spe92, Fis26a]. arrangements [BF36].

Arthur [Fis50b]. Article [Ney56, Fis21b]. ascertainment [Fis34f]. Asexual [Mus12]. Aspects [A.52, Fis50b, Mar45, Ney34, MF45, MF51]. Assay [Fis49a]. Assays [Fis35b]. Assess [Fis59d, Spi59]. assez [FD48]. asymptotic [Fis41a]. Ataxia [BCF39, FBC39, Fis36c]. Atrophy [BF35, PBFN31]. Attempts [Fis59d, Spi59]. Australia [Cor64, Lud05, May04]. Automatic [FO24]. Average [Fis41b]. Aylmer [Ano62, Ano64, Edw92, IBM^{+}63, Ken63, Ken70, Mah38, Mah64a, Moo07b, Owe62, Rao64, Rao07, Rao08, Ski07, YM63, You62].

B [Bin89, Fis26e, Sar95]. B.S [Fis35a]. background [NP76]. Bacterial [TF27, FTM22, Fis34c, TG34]. Balance [FO24, Ros17]. barley [Fis29d]. Bartlett [Fis37b, Fis57b]. Based [Tsu02]. Basic [KJ92a]. Basilarchia [Fis34n]. basis [O’D90]. Bateson [Fis52b]. Baye [Sei92b]. Bayes [Ald08, Dal91, Dal99, Fis26b, Fis62d]. Bayesian [Bar87, Cam95, Fis60b, How99]. Bearing [Fis30i, Fis32a]. became [Sti06]. Been [May04, Fis36k, Fis08]. before [Hal90, Ull99]. Behrens [Bar95, Fis41a, FH56, Yat64a]. belonging [Stu89]. Bennett [K.73, Sen91, All86, Edw87, Hea73, Kem93]. Berkson [Fis43d]. Bessel [FA25]. beta [BS07]. Between [Fis24b, Mus12, Cam95, Fis19a, Fis23d, Fis25c, Fis37d, Fis38i, Fis46e, Fis47b, Len06, MS66, NP76, Pea68, Pea70, Pea01, Stu89, Stu92]. Binomial [BF53a, Fis54d, Fis54e, BF53b, Fis41f]. Bioassay [Bli64]. Biography [Fre83]. Biological [B.58, BF53a, DF22, Fis29g, Fis35i, FY38, FY43, FY48, Fis49a, FY49, FY53a, FY53b, FY57a, FY57b, FY63, K.39, Tip53, Toc58, Wet65, Fis31a]. Biologist [Wil64]. Biology [MS07, TS15]. Biomathematics. [Fis35a]. Biometric [Bil14, Fin64b]. biometrical [Fis24a]. Biometricians [Fis33h]. Biometrics [Fre83]. Biometrika [Bar56, Fis16, Fis56c, TC01]. Biometry [Fis30a, Fis48a, Fis64a, Mor02, Fis11]. birds [Fis37d]. Birth [Fis27a, Fis35i, Edw92, Fis28c, Fis28g, Fis36d, Fis43c]. birth-rate [Fis28c, Fis28g, Fis43c]. birth-rates [Fis36d]. blocks [Fis40b, Fis41g]. Blood [FT39, FV39, FF43, Fis57a, Rac64, ACF^{+}49, Fis51d, Fis53g]. Blood-Group [FF43, ACF^{+}49, Fis51d]. Blood-Groups [FV39]. blossom [Stu89]. Blue [PBFN28]. Blütezeit [Stu89]. Boas [FG37]. Bone [PFSB25, PBFN28]. Book [A.52, All86, Ano45, Ans55, B.50, B.58, Bar57, Bin89, Cal87, Dis91, Edw87, Fis35a, Gri82, H.31, Hea67, Hea73, Hot27, I.26, I.29, I.40, J.79, Jow56, Jow57, K.39, K.73, Kan81, Kem93, Lur72, Mar80, Pea75, Pea51, Sen91, Smi66, Spi59, Tip53, Toc58, Wet65, Y.39, Yat79, Yat82]. Bowling [RS00]. Box [Gri82, Cal87, J.79, Kan81, Yat79]. Brachydactyly [BFP51]. Breakthroughs [KJ92a, KJ92b, KJ97]. Britain [Ayl21, Fis28k, FT39, FR46]. British [Fis37d]. Broadbalk [Fis21c]. bulk [Dem93]. Burnside [Fis23a].

C [AF44]. Calculation [Fis22f, Fis43b, Fis47a, Fis46b]. calculations [Fis51d]. Cambridge [Edw90b, Edw03]. campaign [Ayl21]. can [Fis26d]. Cancer [Fis58j, Fis58k, Fis59d, Spi59, Fis59b]. Capillary [Fis28e, Fis26e]. Card [Fis28a, Fis34j]. Cards [Fis24e]. carried [Fis40a]. Case [Fis35b]. causes [Fis19b, Fis28c]. Celebration [Edw92]. cells [Fis34c, TG34]. cent [FM36b]. Centenary [Sar92, Edw92]. Century [RS00, Efr98, Efr00, Rup07, Sal01]. Cepæa [FD34]. Cereals [Fis27e]. Certain [Fis30c, NP28]. Chair [TS15]. Chance [GSP^{+}89]. Changed [GSP^{+}89]. Changes [TF27]. Changing [Fis35i]. Character [Fis37a]. characteristics [Fis38i]. Characters [Fis32d]. Charles [Fis35a]. chemical [Fis33e]. chi [Cam95]. chi-square [Cam95]. Chiasma [Fis64b, Fis48d]. children [Fis28k, Fis34d]. Christianity [Fis55c]. chromosome [FLO47]. Cigarette [Fis57c]. Cigarette-smoking [Fis57c]. Cigarettes [Fis58k, Fis59b]. civilised [Fis22b]. civilizations [Fis26d]. class [Fis50a]. classical [Leh11]. classification [Fis62e]. Cleido [PFSB25]. Cleido-cranial [PFSB25]. Climatology [FH28]. Cline [Fis50c]. close [Fis53c]. clustering [GGA11]. Co [Fis35a]. Coefficient [Fis15c, Fis21a, Fis24c, Fis28f, Fis36h, SYC^{+}17]. Coefficients [Fis22c, Fis28b, Fis62g]. coherent [FNS84]. Coincidences [Fis24e]. Collected [Ben74, Hea73, K.73, Pea75]. Collegiate [Fis20c]. collide [Hal10]. colony [FF47]. colour [Fis30j]. comb [Fis38b]. Combinatorial [Fis51b, Fis41h]. Combining [MF48]. commemoration [Ano12]. Comment [Bar89, Bar56, Fis57b, Hin80, Kem80, Lan80b, Lin80, Pla89, Rub80]. Commentary [MS66]. Comments [Edi90, SR14, Pea68, Pea70, Pea01]. Commission [Fis49e, Fis49f]. Common [Fis35o]. communities [Fis22b]. Comparative [FH28]. comparison [DF37, Fis39a]. Competition [TS15]. complete [FB49]. Completely [Fis41c, Fis43e]. Complex [Fis35c]. compounded [Fis41e]. Concept [Ano52, Fis42b, Sti73]. Concepts [Sti99b, Fis33a]. Concerning [Fis28a]. Conclusions [Fis48b]. Conditions [Fis24b, FF47]. Confidence [Fis62h]. confounding [Fis41k, Fis42c, Fis43e]. connected [Fis41h]. Connecting [Fis28h]. conscience [Fis22b]. Consequences [SR14, Fis35p, Fis41j]. considerations [Fis22h]. Construction [Wis27]. Consultancy [Fre83]. Consultant [Ben91]. contemporary [Fis32c]. content [Fis23c]. Contingency [CaFJ93, Fis22f, MP86a, MP86b]. Contribution [Fis35c, Fis35d, Fis54a, Sti76]. Contributions [Bli64, FT96, Fin64b, Fis33b, Fis50g, Fis50h, Rac64, Bar63, Cro90, O’D90, Pie90, Str90, Tho90, Pea51]. controversies [Stu92]. Controversy [Fis59d, Len06, PB10, Pro92, SR14, Spi59, Fis49c, F^{+}08, How99, Pie90, Ski00, Stu89]. cooperate [Hal10]. Correcting [TS15]. correction [Fis26e, Fis41c]. Correlation [Fis15c, Fis21a, FM22, Fis24c, Fis28b, Fis28f, FH28, SYC^{+}17, Fis19a, Fis62g, MS66, NP76]. Correspondence [All86, Bar92, Ben83, Ben90, Fis26c, Kem93, Sen91, Pea68, Pea70, Pea01]. Costs [ZM08]. Could [Ber03]. Covariance [Fis47b, DF37]. Cox [Hal10].

Cram. [Fis34n]. cranial [PFSB25]. Craniometry [Fis36h]. Creasy [Fis54a]. Creating [Von94]. creation [Leh11]. Creative [Fis50b, A.52]. Crest [Fis34e]. Criteria [NP28]. Criterion [Fis12b]. Criticism [Fis43d, Fis15a]. Criticisms [Wel56, Fis54f, Fis58i]. critique [Pil07]. Croonian [Fis53a]. Crop [EF29, Fis21c, FM23, Fis27e, Fis24g]. Cross [Fis62h, Fis22i]. cross-over [Fis22i]. Cross-Product [Fis62h]. Crossing [FD34, Fis53c, Fis54b]. Crossing-over [FD34, Fis53c]. cubes [Fis43e]. Cuénot [Fis15a]. Cult [ZM08]. Cumulants [CF38, FC60]. Curve [Fis36a]. Curves [Fis12b]. customs [Stu89]. cyclic [Fis41g].

D [Fis62f, Yat82]. Daily [TF27]. Danausplexippus [Fis34n]. Dangers [Fis57c]. d’arithmétique [FD48]. Darwin [All86, Edw87, Ben83, Fis58c]. Darwinian [Fis22a]. Darwinism [Fis47f, Fis47g]. Data [Bas80a, Bas80b, BF53a, Fis23b, Hin80, Kem80, Lan80b, Lin80, Rub80, Fis22d, Fis23c, Fis36c, Fis36g, FG37, Fis47a, Fis46d, FB49, GGA11, Bin89]. Death [Lud05]. decay [Fis26d]. Deduced [Fis21a]. defect [Fis24d]. defectives [Fis34d]. degli [Fis62b]. Degree [FIT32]. del [Fis62b]. Deliver [Fis32g]. della [Fis62b]. Deming [Dem93]. density [FTM22]. Departure [Fis30h]. dependent [DF37]. Depopulation [Fis29c]. Derivation [FW31]. Derivative [FA25]. Derivatives [Fis31f]. Design [Cra36, Dis91, Fis35e, Fis37f, Fis38a, Fis42e, Fis47i, Fis47c, Fis49h, Fis51e, Fis60c, Fis62c, Fis66, Fis71, FBY90, Hea67, Pea79, Sei92a, Ull99, Yat64b, Box80, Fis45f, Fis65a, Hal02, Pre90]. detection [Fis35f, Fis35g, Fis62a]. Determination [Fis27e, Fis62d]. Determined [Fis50c]. Determining [Fis20b]. Development [Fis33b, Fis47c, PFSB25, Fis39g, Rya80]. Developments [Leh90]. Deviates [Fis31f]. Deviations [Ano15, Fis50e, Fis64c]. diagonal [Fis41h]. Diaphysial [PFSB25]. did [Edw97]. Difference [FF43, Fis46e, Fis62a]. differences [Fis53c]. Different [Fis43f, Ney34, FM23, Fis40b]. differential [Fis28c, FB49]. Differentiation [Fis61b]. Diffusion [Fis50c]. Digital [Ano13, BFP51]. Discontinuous [Fis58a]. discovery [Sti73]. Discrepancey [Fis24b]. Discrepancy [Fis28h]. discriminant [Fis40e]. Discussion [Fis35d, Fis54a, Fis54e, WTRS^{+}36, Fis35c, Sti76]. disease [PBFN31]. Diseases [BF35, BCF39, BF43, FB34, PBFN28, PBFN31]. Disorders [PFSB25]. Dispersion [Fis53e]. Distribution [Ano15, Aro41, BF53a, Fis15c, Fis22c, Fis24c, Fis24f, Fis25a, Fis28f, FT28, Fis30b, Fis30h, Fis43f, Gup60, Ken48, KJ92b, SYC^{+}17, Wis27, Aro40, BS07, Fis39e, Fis41f, Fis62g]. Distributions [CF38, Fis29a, Fis35o, FC60, Fis40d]. Dominance [Fis28d, Fis28l, Fis30c, Fis34k, Fis35h, Fis38b, Fis22e, Fis31b, Fis34e, Fis34i, Fis90]. dominant [Fis35f]. dominula [FF47]. Double [FF55, Fis43b, Fis62a]. Dr [Fis35l]. Dr. [Fis23a, Fis28a, Fis43d]. draft [Sar95]. dressed [Fis21c]. Dressings [Fis27e]. drift [Tur87]. Drill [Fis36b]. due [Fis34e]. Duplex [FM47]. Dynamics [Fis35i]. Dystosis [PFSB25]. Dystrophies [BF35, BCF39, BF43, FB34]. Dystrophy [BF43].

Early [Ull99, Pea68, Pea70, Pea01]. East [Fis35p]. ecological [Tur87]. economic [Fis32c]. Economics [TS15]. Ed [Fis35a, Sen91, Tuk52]. Eddington [Fis50b]. Edgeworth [Pra76]. Edinburgh [NP76]. edition [Fis35a]. Editor [Fis31d, Fis62f]. Editorial [Ano15]. Effect [Fis28a, FF50, PL12, Fis27b, Fis29d, Fis34f, Fis41b, Fis55a]. effects [Fis31a]. Efficiency [Pra76]. efficient [BF53b]. eggs [Fis37d]. Eigen [Mus12]. elimination [Fis24d]. Embleton [Bin89]. Eminent [Rup07]. emphasis [Pie90]. Empire [GSP^{+}89]. Enchondromata [PFSB25]. Ending [F^{+}08]. englischen [Stu89]. English [Stu89, Stu92]. Enigma [Fis34j]. enough [Pil07]. Enumeration [Fis62e]. enumerations [Fis41h, Fis50a]. Environment [Tab08]. equation [Fis25e, Har08, Fis25e]. equations [Fis39e]. Eratosthenes [Fis29e]. Ereporenten [FY44]. erporents [FY44]. Errata [Ano60]. Error [Fis20b, Fis21a, Fis24f, Fis31f, Wis27, ZM08]. errors [Fis23a, Fis55a]. especially [Fis33f]. esperimenti [Fis62b]. Essays [HHF54, HHF58]. Estadística [Fis55b]. Estimated [Fis31f, Fis34c, TG34]. estimating [FTM22]. Estimation [Fis25f, Fis54a, Pra76, Sei92a, FB28, Fis34f, Fis34m, Fis38e, Fis38g, Fis39d, Fis40a, FB49, I.40]. Ethnology [FT40]. Eugence [RS00]. Eugenics [All86, Ben83, Fis26d, Fis35j, Ayl21, Fis17, Moo07a, Edw87]. Eugenist [Fis14]. Euler [Ull99]. Eulerian [Ull99]. evaluating [Fis51d]. Evaluation [FA25, Fis34m, Ski00]. Ever [Fis33g]. Ever-sporting [Fis33g]. Everyday [GSP^{+}89]. Evidence [Fis59d, Fis33e, Spi59]. Evolution [Fis12a, Fis28d, Fis30a, Fis30c, Fis31e, Fis32a, HHF54, HHF58, Cro90, Fis15b, Fis22a, Fis22b, Fis31b]. Evolutionary [Kar92, Sar92, Fis32b, Moo95]. Exact [CaFJ93, MP86a, MP86b, Tsu02, Cam95, Fis59e]. Examination [Fis20b, Fis21c, Fis25c, Fis40b, Fis59e]. example [DF37]. examples [Fis62d]. excess [Fis41b]. Exchange [Inm94, PFI94]. Existence [TF27]. Exophthalmic [Mar45, MF45]. Exostoses [PFSB25]. Expansion [Fis25b, Fis53b]. Expectation [Fis50e, Fis64c]. Experience [Ber43]. Experiences [Wil64]. Experiment [Fis34k, Fis62b]. Experimental [Bas80a, Bas80b, Dis91, Fis27e, Fis47c, FBY90, Hin80, Kem80, Lan80b, Lin80, Pea79, Rub80, Ull99, Fis41d, Fis45f, Fis54b, Fis62d, Hal02, Pil07, Pre90]. Experimentation [FT96, Fis31c, Fis35d, Fis38c, Fis52a, You51]. Experiments [FW30, Fis35c, Fis35e, Fis36b, Fis37f, Fis38a, Fis42e, Fis47i, Fis49h, Fis51e, Fis60c, Fis62c, Fis66, Fis71, Fis92a, Men65, Sei92a, Spe92, Yat64b, Box80, Fis26a, Fis41k, Fis42c, Fis65a, EF29, Cra36, Hea67, Smi66]. Explain [PB10]. exploratory [GGA11]. Extensions [Fis60b]. Extent [May04]. Eye [PBFN28, PBFN31].

F [B.58, Fis37a, K.39, Toc58, Wet65]. F. [Fis35c, Pra76]. F.R.A.S [Fis35a]. F.R.C.P. [Fis45b]. F.R.S [Fis35a, Owe62]. F.S.S [Owe62]. faba [DF22]. Factor [Fis74, Fis39g, FM40]. Factorial [Fis38a, Fis41k, Fis42c]. Factors [Fis33d, Fis43e, Fis46d]. Fairly [Edw03, Edw90b]. faith [Moo07a]. Familial [Mar45, MF45, MF51]. families [Fis34b]. Family [Fis32c, Ayl21, Fis27b, Fis31a, Fis36d, Fis43c]. fastperiodische [FY44]. feathered [Fis38b]. February [Ano12]. feet [Fis38b]. Feldman [Fis35a]. Fertility [Fis32g, Fis35l]. FEXACT [CaFJ93, MP86a, MP86b]. fiduciaires [Fis48b]. Fiducial [Fis39c, Sei92b, Yat64a, Zab92, Fis33a, Fis35k, Fis37b, Rya80, Fis48b]. Field [FW30, Fis92a, Spe92, Fis26a, Fis34c, Fis51a, TG34]. Fienberg [Yat82]. fifth [Fis49d]. Fifty [Bar65, FM36b]. figures [Fis28c]. Fisher [All86, Ano62, Ano12, Ano13, B.58, Bin89, Cal87, Edw87, Edw93, Efr98, Gri82, Hea73, Hot27, J.79, K.39, K.73, Kan81, Kem93, Owe62, Pea75, Pro92, Ski00, Stu89, Toc58, Wet65, Yat79, Yat82, A.52, Ald95, Ald97, Ald05, Ald08, Ald13, And96, Ano15, Ano45, Ano60, Ano64, Ans55, Aro40, Aro41, Ayl21, B.50, Bar63, Bar87, Bar89, Bar90, Bar95, Bar56, Bar57, Bar65, Bas80a, Bas80b, Ben74, Ben83, Ben90, Ben91, Ber03, Ber43, Bil14, Bli64, BS07, Box78, Box80, Box87, Box05, Cam95, CaFJ93, Con92, Cor64, Cra36, Cro90, Dis91, Edi90, Edw90a, Edw90b, Edw92, Edw94, Edw97, Edw03, Efr98, Efr00, FH80, FT96, Fie97, Fin64b, Fis28a, Fis31d, Fis46a, F^{+}08, Fre83, GV99, Gei92, Gre03, Gro30, H.31, Hal02, Hal07, Hal10, Har08]. Fisher [Hea67, Hea03, Hin80, Hot51, How99, I.26, I.29, I.40, Inm94, IBM^{+}63, Jef74, Joh87, Jow56, Jow57, K.73, KM13, Kar92, Kem80, Ken48, Ken63, Ken70, Lan80a, Lan80b, Leh90, Leh93, Leh11, Len06, Lin80, Lin90, Lud05, Lur72, Mah38, Mah64a, Mah64b, Mat51, Mat64, May04, May14, MP86a, MP86b, Moo95, Moo07b, Moo07a, MS66, Mor02, Mus12, Ney56, NP76, O’D90, PL12, Pea75, Pea79, Pea92, Pea51, Pea68, Pea70, PFI94, Pea01, Pie90, Pil07, PB10, Pla89, Pra76, Pre90, Qui21, Rac64, Rao64, Rao92, Rao00, Rao07, Rao08, Ros17, Rub80, Rya80, SR14, Sar95, Sav76, Sei92a, Sei92b, Ski07, Smi66, SYC^{+}17, Spe92, Spi59, Sti73, Sti76, Sti05, Sti06, Str90, Stu89, Stu92, Tab08, TS15, Tho90, Tip53, Tsu02, Tuk52, Tur87, Ull99, Wel56, Wil64]. Fisher [Y.39, YM63, Yat64b, You62, Zab89a, Zab89b, Zab92, Kem93, Mar80, Sen91, Bar92]. Fisherian [You51]. Fit [Fis22c, Moo07a]. Fitness [GV99]. Fitting [BF53a, Fis12b, Fis36a, BF53b, Fis47a]. focus [Ski00]. fold [Fis41d]. Following [Joh87]. food [Fis29b]. Forces [Fis28e, Fis26e, Fis39f]. Forestry [Fis42a]. Foreword [Fin64a]. form [Fis41j]. Formation [Fis64b, Fis48d, Fis59e]. Forms [FT28]. formula [Fis39d]. Formulae [Fis22c, FW31, Fis26e]. formulated [Fis21b]. Formulation [Fis51b]. FORTRAN [MP86a, MP86b]. found [Fis40d]. Foundations [Fis92b, Gei92, KJ92a, Sti99a, Fis22j]. Founder [Rao92, Rao00]. Founders [Sar92]. four [Fis41d]. four-fold [Fis41d]. Fourfold [Fis26b]. Fourth [Fis50b]. fowls [Fis34e]. Fragility [PBFN28]. Frank [Tip53]. French [Fis25e, Fis38e, Fis47d, Fis48b, FD48]. Frequencies [FF43, FR46, Fis50c, Fis34f, Fis43b, Fis47a, Fis46b]. Frequency [Fis12b, Fis15c, FT28, Fis39g, FB49]. Friedreich [Fis36c]. Fulfilled [May04]. fuller [Fis54g]. Function [FA25, Fis31f, Fis34b]. Functions [Fis24f, Fis40e, FY44]. Fundamental [FT96, Fis42a, GV99, Edw90a]. Funktionen [FY44]. Further [Fis28e, Fis28l, Fis41a]. Future [Fis36h, RS00, Fis28g].

G [Fis34c]. G. [Fis45b]. Galton [Edw93, Fis39b, Fis40c]. Gamma [Gup60]. Gene [Fis30b, FR46, FM47, Fis50c, Fis34e, Fis41b, Fis47a, FF47]. General [Fis20c, Fis28f]. generalizing [FNS84]. Generation [Fis61a]. genes [Fis22i, Fis37e]. genesis [Fis19c, Fis20a, Fis22d]. Genetic [Fis64b, Fis27c, Fis32b, Fis48d, Tur87]. Genetical [Fis30d, FIT32, Fis58b, Fis59c, Fis99, H.31, Tho90]. Geneticist [Edw03, Edw90b]. Genetics [Fis30e, Fis32a, FM48, Mat64, Sar92, Cro90, Edw90b, Fis48c, Fis49c, Fis50a, Fis52b, Fis53a, Fis57a, Fis57d, Pie90, Ski00, Tur87, Edw03, Fis31e]. Genotype [Tab08, Fis43b]. geometrical [Fis40d]. geometry [Fis13]. German [FY44, Fis56b, Stu89]. Gertrude [Hal10]. given [Fis26e]. giving [Fis43e]. Goitre [Mar45, MF45, MF51]. Goodness [Fis22c]. Gosset [Box87, Pea68, Pea70, Pea01]. grain [Fis21c]. Great [Fis28k, FT39]. Green [RS00]. Gregor [Smi66, Pie90]. Grey [Fis34c]. Griffin [Fis35a]. Group [FF43, ACF^{+}49, FS48, Fis49d, Fis51d, Fis53g]. grouping [Fis55a]. Groups [FV39, Rac64, FT39, Fis41k, Fis42c, Fis57a]. Growth [Ken48, Fis21b]. Guide [Ald13]. Guinness [Box87]. Gupta [Ano60].

H [All86, Edw87, Fis34c, K.73, Kem93, Sen91]. H. [Moo95]. Haines [Fis26e]. Haldane [Fis36g, Sar95]. Half [Fis36b]. Half-Drill [Fis36b]. Harmonic [Fis29h, Fis40d]. Harold [How99]. Having [FC60, DF37]. health [Fis26g]. Herbert [Fis32g]. Hereditary [BCF39, BFP51, FBC39, Mar45, Fis35m, MF45, MF51, PFSB25, PBFN31]. Heredity [All86, Ben83, Edw87, Fis24a, Pie90]. hernia [Fis34e]. Herring [BWF33]. Heterogeneity [Fis36c]. heterostylism [Fis35p]. heterozygotes [FB28, Fis62a]. high [Fis38i]. high-altitude [Fis38i]. higher [Fis43e]. Hinckley [Yat82]. Historical [Rya80, Von94, Stu89]. historische [Stu89]. History [Bar89, Hal98, KP77, KDH87, PK70, Pea01, Pla89, Sti99b, Stu92, Ull99, Zab89a, Zab89b, Dal91, Dal99, Hal90, Pea68, Sti73]. Hogben [Tab08, TS15]. homostyle [Fis49g]. Hopes [Fis14]. house [FLO47, FS48, Fis53f]. house-mouse [Fis53f]. Hr [Fis45a]. Human [BF35, BCF39, BF43, BFP51, Fis32g, FB34, FBC39, PFSB25, PBFN28, Rac64, Fis19b, Fis53g, Fis57d, Fis61e, PBFN31]. hundred [TC01]. Hunt [FdB47]. Hybridisation [Men65, Smi66]. Hypotheses [Leh93]. Hypothesis [Fis24b, Fis23d].

icarus [DFF40]. Ideal [Fis28e, Fis26e]. Ideas [KDH87, KGM87]. II [BF35, Fis22h, FM23, Fis26h, KP77, KGM87, SYC^{+}17]. III [BCF39, Fis24g]. Illustrated [Gri82]. Illustrations [Fis31f]. Impact [Hot51]. Imperfect [PL12]. importance [Fis50a]. Impossibility [FNS84]. inattendu [FD48]. Inbreeding [Fis49i, Fis49j, Fis65b, Fis54g, Fis59e, B.50]. Including [All86, Ben83]. Income [Fis28g, Fis36d]. Income-tax [Fis28g, Fis36d]. incomplete [Fis40b, Fis41g]. increasing [Fis29d]. Indefinitely [Fis15c]. Independent [MF48, DF37, FY44]. Indeterminism [Fis34g]. Indian [Fis45c]. induction [Fis33e, Fis55e]. Inductive [Fis35n]. Inference [Bar92, Fis35n, Fis39c, Fis56d, Fis56e, Fis59f, Fis60b, Fis62c, Fis73a, FBY90, Kem93, Len06, Mar80, NP28, Ben90, Con92, Fis34h, Fis35k, Fis36i, Fis62b, Fis65a, Sen91, Bar57, Dis91, Jow57]. inferenza [Fis62b]. Influence [FT40, Fis24g]. influencing [Fis39g]. Information [PL12, Fis34b, Fis34h]. Inheritance [BF35, BCF39, BF43, BFP51, FIT32, Fis32d, Fis33c, Fis33d, FB34, FG37, FBC39, FM42, Fis58a, PFSB25, PBFN28, Fis19a, Fis35l, Fis41j, Fis43b, FM43, Fis47h, Fis62e, NP76, PBFN31, MS66]. Inquiry [Ano52]. insects [Fis33f]. Inspired [Fre83]. Institute [Fis45c]. integral [Fis25b, Fis25e, Fis41a]. intégrale [Fis25e]. Integrals [Fis31f]. Intensity [Fis36e]. Interaction [Tab08, Fis51a]. internal [Fis38b]. International [Bil14]. Interpolated [Wis27]. Interpretation [Fis22f, Fis31c, FIT32, NP28, Fis26h, Fis41d]. Interpretations [How99]. Interval [Fis54a]. interwar [Ayl21]. Introduction [Fie97, Gei92, Pea92, Spe92]. Inverse [Bar89, Fis30f, Fis32e, Jef74, Pla89, Zab89a, Zab89b, Dal91, Dal99, Edw97, Fie97, Fis33a]. inversion [Fis45d]. Invited [Efr98]. involved [Fis22h]. Ireland [Fis28k]. Italian [Fis62b]. IV [Fis27e].

J [All86, Bin89, Edw87, Fis22g, K.73, Kem93, Moo95, Sen91, Stu89]. J. [Fis35d, Sar95, Stu89, Stu92]. Jeffreys [Ber03, How99, Lan80a]. Joan [Cal87, Gri82, J.79, Kan81, Yat79]. Jobs [ZM08]. John [Gri82]. Journal [Fis57b]. July [Ano12]. junction [Fis59e]. Junctions [Fis54g]. June [Ano12, Fis32g]. Justice [ZM08].

Karl [Fis37c, BS07, Dal91, Dal99, Edw94, Inm94, Pea68, Pea70, PFI94, Pea01, Sar95]. Keynes [Con92, Fis22g]. Knibbs [May04]. Known [Edw03, Fis24f, FC60, Edw90b].

L [Fis45b, FD34]. L. [MF45, FF47]. Laboratory [Fis39b, Fis40c]. Lady [Sal01]. Lancelot [Tab08, TS15]. Land [FD34]. Laplace [Sti73]. Large [Fis15c]. Largest [FT28]. Last [Bar65]. Later [Leh90]. Latin [Ull99, FY34, Fis41h]. Lauterbach [Fis25c]. Law [Fis50b, A.52]. laws [FD48]. leaden [Fis53f]. Leading [Ken48]. Learning [Mar80]. Leber [PBFN31]. Lecture [Fis50b, Fis52b, Edw93, Efr98, Fis32g, Fis53a, May04]. Legend [TS15]. Length [FM48, FM43]. Leonard [All86, Edw87, Ben83]. Lepidoptera [DFF40, DFF49, FF28, Fis33e]. lethality [FM40]. Lethals [Fis35q]. Letter [Fis31d, Fis62f]. level [Fis38i]. Levine [Fis45a]. Lewis [Bin89, ACF^{+}49, Fis62f]. Life [Cal87, Fis58c, GSP^{+}89, Gre03, J.79, Kan81, Lud05, Owe62, Rao07, Rao08, Yat79, Box78, Gri82, Sar95]. light [Fis28c]. Likelihood [Ald95, Fis28h, Fis34o, Pra76, Ald97, Fie97, Fis32e, Fis34h, Fis41e]. Likeness [Fis36h]. Limiting [FT28]. Limits [Fis62h]. Lindley [Fis60b]. linear [Fis39e, FY44]. Linkage [Fis35m, Fis51b, FB28, Fis34b, Fis34m, Fis35f, Fis35g, Fis36c, FM36a, Fis36g, Fis39d, Fis46d, Fis47h, FS48, FB49, Fis49b, Fis49d, Fis53f, Fis53c]. links [Con92]. Linn. [Fis34n, DF22]. Lives [Rup07, ZM08]. location [Fis22i]. Locus [FF55]. Logarithmic [Ken48]. Logic [Fis35n, Fis62c, Fis34h, Fis62b, Fis65a]. logica [Fis62b]. logical [Fis45d]. lois [FD48]. Lond [Fis35a]. London [Edw90b, Edw03, TS15]. Ltd [Fis35a]. Lukacs [RS00]. Lung [Fis58k]. Lutheran [ACF^{+}49]. Lysenko [Fis48f]. Lythrum [Fis35p, FM40, Fis41j, FM42, FM43, FM47, Fis49b, FF55].

M [Fis22g, Fis35a]. M. [Fis25e, Fis37b, Fis54a]. M.D [Fis35a, Fis45b]. M.R.C.P [Fis35a]. Made [Wel56]. Making [Ald95, Gan82, Ald97]. Man [Fis48f, FG37]. Management [Fis42a]. manners [Stu92]. Manslau [Fis35l]. manurial [FM23]. March [Fis31e]. Martin [MF45]. mass [Ros17]. Mathematical [Fis20b, Fis34o, Fis35o, Fis42a, Fis50g, Fis50h, Fis58d, Fis59a, Fis59g, Fis60a, Fis92b, Gei92, Hal98, Pea51, Bar63, Fis22j, Sti06]. Mathematics [Fis20c, Fis30e, Fis30g, Fis33c, Fis38c, Von94]. Maximum [Ald95, FH28, Fis28h, Pra76, Ald97]. Maximum- [FH28]. Mean [Fis20b, GV99, Edw97, Fis61d]. Means [FO24, Fis46e, DF37, Fis22i]. Measure [Fis28h]. Measurement [Fis36e]. measurements [Fis25c, Fis36j, Fis37d, Fis38h]. Measures [Fis24b, Fis30h]. Mechanical [FO24]. Medical [B.58, FY38, FY43, FY48, FY49, FY53a, FY53b, FY57a, FY57b, FY63, K.39, Tip53, Toc58, Wet65]. Medicine [Bli64]. melanism [Fis33e]. Member [FT28]. Memorial [Fis50b, PBFN28, You62, PBFN31]. Memoriam [Ano64]. Memories [Mah64b]. Mendel [Smi66, Edw93, Fis36k, Fis08, F^{+}08, Har08, Pie90, Pil07, PB10]. Mendelian [Fis31e, MS66, NP76, Fis19a, Fis33d]. Mendelism [Fis11, Mor02]. mental [Fis24d, Fis34d]. Mesoscopic [GGA11]. Meteorology [Fis28b]. Method [Ano45, Fis24e, Fis28h, Fis29f, Fis47b, Fis74, Ney34, BS07, FTM22, FG37, Fis37c, Fis62d, TG34, Fis34c]. Methoden [Fis56b]. méthodes [Fis47d]. Methodologies [RS00]. Methodology [Bar65, KJ92b]. Methods [Ans55, Bar57, BWF33, Dis91, Fis20b, Fis25d, Fis28j, Fis30k, Fis32h, Fis34l, Fis36f, Fis38f, Fis41i, Fis42a, Fis44, Fis46c, Fis48e, Fis50f, Fis54c, Fis56d, Fis56e, Fis56b, Fis57d, Fis58h, Fis59f, Fis63, Fis67, Fis70a, Fis70b, Fis73b, Fis73a, FBY90, Fis92c, Gro30, Hot27, I.26, I.29, Jow56, Jow57, Lur72, Mat51, Pea92, Sti99b, Tuk52, Y.39, You51, Fis21b, Fis34f, Fis47d, Fis52b, Fis55e]. Mice [Fis32f, Fis50d, FM36a, FM36b, Fis46d, Fis49d]. Mid [FM40, Fis41j]. Milk [FB31]. mimicry [Fis27c]. Minimum [FH28]. Minimum-Correlation [FH28]. mistaken [Fis34n]. Model [Fis61a, Leh90, Mus12, PB10, Fis34n]. Modelling [Mor02]. Models [Len06]. Modern [BWF33, Fis48c, Pea79, Rao92, Von94, Rao00]. Modes [Ken48]. Modification [Fis28i, Fis32b]. moment [BS07]. Moments [CF38, Fis29a, Fis30h, Fis37c]. Monarch [Fis34n]. Morgan [FdB47].

Mortality [Fis30i]. Mosaic [Fis30j]. moth [FF47]. mouse [Fis30j, FLO47, FS48, Fis53f]. Mr [Fis60b]. Muller [Moo95]. Multiple [Fis28f, Fis51b, DF37, Fis36j, Fis38h, Fis54b, PFSB25]. Multivariate [Rao64, And96]. Muscular [BF35, BCF39, BF43, FB34]. mutants [Fis39g]. Mutation [FRT44]. Mutations [Fis28i, Fis30b, Fis22a, Fis34a].

N [Bin89]. naively [FNS84]. narrate [Ros17]. Nations [Fis49f]. Natural [A.52, Ben83, Fis30e, Fis30d, Fis30i, Fis34g, Fis50b, Fis58b, Fis58f, Fis59g, Fis59c, Fis99, H.31, WTRS^{+}36, FF47, Fis54f, Fis58d, Fis58g, Fis58i, Fis59a, Fis60a, All86, Edw87]. Nature [Fis58e, Lan80a, Inm94, PFI94]. Negative [BF53a, BF53b, Fis41f]. nella [Fis62b]. nemoralis [FD34]. Nervous [BF35, BCF39, BF43, FB34]. nesting [Fis37d]. Nettleship [PBFN28, PBFN31]. networks [GGA11]. Neyman [Stu89, Bar95, Ber03, FT96, Fis35d, Fis57b, Leh90, Leh93, Leh11, Len06, SR14, Stu89, Stu92]. Nile [Fis46a, KM13]. nitrogen [EF29, Fis23c]. No [Bar56, Fis57b, Sen91, Fis56c]. Nodular [Mar45, MF45, MF51]. Non [FM40, Fis39e]. Non-lethality [FM40]. non-linear [Fis39e]. Normal [Fis30h, Fis31f, Fis61d]. Normality [Fis30h]. notation [ACF^{+}49]. Note [BF53b, Fis23a, FA25, Fis28e, Fis30j, Fis39c, Fis43d, Fis46b, Ney56, Wel56, Fis23c, Fis29d, NP76, MF45, Fis47a, FB49]. Notes [Fis28l, Fis57b, Gre03, Rac64, Fis26c, Pea68, Pea70, Pea01]. notion [Fis45d]. November [Edw93, Fis50b]. Number [Fis33d, Fis47e, Fis39g]. Numbers [TF27, Fis34c, Fis41h, TG34]. Numerical [FA25].

obituary [Fis37a, Fis45b, FdB47]. objections [Fis27c]. Observation [Fis20b, Fis24b, Fis23a, Fis23d]. Observations [Fis46e]. obtained [Fis39e]. Occurrence [FM47]. Oenothera [Fis61b]. Official [Fis27a]. offspring [FB28, Fis39g, Fis59e]. OH [RS00]. Old [Fis34j]. One [Leh93, TC01]. op [Fis38i]. optic [PBFN31]. Order [Gup60]. Organensis [Fis61b]. Origin [Fis28l, Tab08]. Orthogonal [Fis41c, Fis43e]. Other [Fis31f]. our [Fis28g]. overproduction [Fis29b]. Oxalis [FM48]. Oxford [Sen91, Fis32g, Tur87].

P [Fis34c, Fis53g]. pages [Sen91]. Panaxia [FF47]. Paper [Bar56, Fis35c, Fis35d, Fis54a, Efr98, Fis22h, Fis23a, Fis26h, MF45, MS66, NP76]. Papers [Ano15, Fis49c, Hea73, K.73, KP77, Pea75, SYC^{+}17, Ben74, Dem93]. Paradoxical [Fis23b]. parameters [Fis33a]. Paraplegia [BCF39, FBC39]. Paratettix [Fis39f]. parent [Fis59e]. Part [Fis47b, BF35, BCF39, BF43, BFP51, FBC39, NP28, PFSB25, PBFN28, PBFN31]. Partial [Fis24c, Fis36g]. Partitions [FW31]. Pasteurised [FB31]. Pattern [FW31]. Patterns [FW31]. Pearson [Bar56, BS07, Cam95, Dal91, Dal99, Edw94, Fis37c, Fis56c, Inm94, Leh93, Len06, Mor02, Pea68, Pea70, PFI94, Pea01, Sar95]. Percentile [FC60]. Performing [CaFJ93]. periodic [FY44]. Periodical [Fis26g]. Peroneal [BF35]. Personal [Mah64b, Har08]. Perspective [Von94, Cam95]. Ph.D [Fis45b]. phenomena [Fis32b]. philosophical [Ski00, Mar80].

Philosophy [MS07, Von94]. Physics [Von94]. piano [Fis62b]. picture [Stu92]. pied [Fis46d]. pigment [Fis38b]. pile [Fis38b]. Pink [FF55]. pioneers [Hal10]. Place [Fis62c, Fis65a]. plan [Fis62b]. Plant [Men65, Smi66, Fis21b, Fis49b]. Plants [Fis30i, FM47]. plating [FTM22]. Play [Fis34j]. Playing [Fis24e]. Plot [Fis31c]. point [Fis37b, FB49]. Points [FC60]. Poisson [Fis50e, Fis64c]. policy [Fis28g]. Polydactyly [Fis34k, Fis50d, Fis53f]. Polygene [Fis42b]. Polymorphic [Fis30c]. Polymorphism [Fis58f, Fis58g]. Polyommatus [DFF40]. Polyploid [FM42, Fis41j]. polysomic [Fis43b, Fis47h, Fis62e]. Population [Fis15c, Fis29c, Fis35i, Fis53a, Fis61b, Ken48, Fis27b, Fis34b, Fis57a, Ski00, Fis49e]. Populations [Mor02, DF37, DFF40, DFF49, FTM22, Fis31e, Fis39f]. portrayal [Stu89]. Positive [Fis17]. possibility [FM36b]. Possible [Fis28i, Fis61b, Fis40b]. possibly [Fis39a]. potash [EF29]. potato [EF29, FM23]. Poultry [Fis34k, Fis35h, Fis38b]. Powers [Fis25b, Fis43e]. Pp [Fis35a, Gri82]. practical [Fis35j]. Practice [Fis27d]. preadaptation [Fis15a]. Precision [Fis46e, BF36, Fis39d, Fis40e]. preference [Fis15b]. Preferences [Fis28a]. preliminary [Fis29d, Fis49d]. Present [WTRS^{+}36]. presented [Efr98]. presents [Ano12]. Presidential [Fis38d]. Press [Sen91]. Price [Sen91]. Primula [Fis49g]. Principles [Fis27d, Fis31c, Fis42a]. priori [Fis62d]. Probabilistic [KDH87, KGM87]. probabilité [FD48]. probabilities [Fis41e, Fis62d]. Probability [Bar89, Fis34h, Fis57e, Fis58e, Fis59g, GSP^{+}89, Jef74, KP77, PK70, Pea01, Pla89, Von94, Yat64a, Zab89a, Zab89b, Con92, Dal91, Dal99, Edw97, Fie97, Fis30f, Fis32e, Fis33a, Fis37b, Fis40d, FD48, Fis58d, Fis59a, Fis60a, Hal90, How99, Lan80a, Pea68, Sti73, Fis22g]. Probable [Fis21a]. Probit [Fis35b]. problem [Bar95, Fis26d, Fis40b, Fis40d, Fis41e, Fis49b, KM13, Fis46a]. Problems [BWF33, Fis35d, Mar80, Fis36j, Fis41g]. process [HHF54, HHF58]. Product [Fis29a, Fis62h, Fis39d]. Production [Fis33g]. professor [Edw90b, Ber43, Edw03, Fis28d, Fis34i, Fis37c, Fis62f, Mah38, Mah64a]. prognosis [Fis35m]. Program [May04]. progress [Fis45f]. Progressive [BF35, BF43]. Properties [Fis31f, Fis34o]. Property [Fis28h]. proportion [Fis40a]. Proposed [Fis60b]. Protective [Fis33f]. Pseudohypertrophic [BF43]. Psychical [Fis29f]. Psychological [Fis28a]. Publications [Cra36, Fis31d, Gro30]. Purposes [NP28]. Purposive [Ney34].

quality [Fis51a]. Quantitative [FIT32, Fis33d, Fis64b, DFF40, DFF49, Fis21b, Fis48d]. quantities [Fis39g]. quantity [Fis34h, Fis51a]. Quelques [Fis38e]. Query [Fis51a, Fis55a]. Question [MF48]. Questions [MF48].

R [A.52, All86, Ano45, Ano12, B.50, Bar92, Ber43, Cal87, Dis91, Edw87, Efr98, H.31, Hea73, I.26, I.29, I.40, K.39, K.73, Kem93, Lur72, Mar80, Pea75, Pea51, Sen91, Sti76, Stu89, Toc58, Tuk52, Wet65, Y.39, Yat82]. R.

[Ald95, Ald97, Ald08, Ald13, And96, Ano15, Ano13, Aro40, Aro41, Ayl21, Bar87, Bar89, Bar65, Ben74, Ben83, Ben90, Ben91, Bli64, BS07, Box78, Box80, Box05, Cra36, Edw90b, Edw94, Edw03, Efr98, Efr00, FH80, Fie97, Fis31d, Fre83, Gre03, Gro30, Hal02, Hal07, Hea03, Hot27, Hot51, How99, Inm94, Joh87, Kar92, Ken48, Lud05, Mah64b, Mat51, Mat64, Moo95, Moo07a, MS66, NP76, Pea79, Pea68, Pea70, PFI94, Pea01, Pla89, Pra76, Pre90, Pro92, Rao92, Rao00, Rya80, Sar95, Sav76, Sei92a, Sei92b, Ski00, SYC^{+}17, Stu89, Stu92, Tab08, TS15, Tho90, Tur87, Ull99, Wil64, Zab89a, Zab89b, Zab92, Cal87, Gri82, J.79, Kan81, Yat79, Yat82]. Race [Ano52, BWF33]. Racial [Fis36h]. Rainfall [FM22, Fis23b, Fis24g]. raised [Fis37b]. Random [Fis26f, Tur87, Fis45d]. Randomisation [Fis34j]. Randomization [Bas80a, Bas80b, Hin80, Kem80, Lan80b, Lin80, Rub80, Hal07]. randomized [Hal02]. Range [Fis42a]. Ranked [PL12]. Ranking [PL12]. Rare [Fis30b]. rate [Fis28c, Fis28g, Fis43c]. Rates [Fis35i, Fis36d]. rather [FD48]. Ratio [Fis62h, Fis22e, Fis34c, Fis61d, Fis90, TG34]. Ratios [Fis30b, Fis22i]. Raw [FB31]. Reaction [FRT44]. reactions [Fis47a]. Reading [Fre83]. Reappraisal [Sar92]. reasoning [Fis61e]. rebates [Fis28g]. recessive [Fis35g]. recessives [Fis40a]. recherche [Fis47d]. Recognisable [Yat64a]. Recombination [Fis64b, FM36b, Fis48d, Fis62a]. Reconsidering [FT96]. Records [Fis27a, Fis34b]. Recurrent [Fis28i]. rediscovered [Fis36k, Fis08]. Reduction [FW30, FF55, Fis43b]. refereeing [NP76]. Reference [Fis34k, FF28, Fis46d, Fis61c]. referring [Fis33a]. refinement [Fis61e]. Regression [Ald05, Fis22c]. Rejoinder [Bas80b, Zab89b]. Related [Aro41, Aro40, Fis39g]. relating [Fis23c]. Relation [Fis47b, Mus12, Fis22b, Fis31c, Fis37d, Fis38i, Fis41k, Fis42c]. relationship [Cam95]. Relative [GV99]. relatives [Fis19a, MS66, NP76]. Religion [Rup07]. Remark [CaFJ93]. remarks [Fis21b, Fis38e]. remarques [Fis38e]. Reminiscence [Box05]. renaissance [Fis47f, Fis47g]. Reply [Ber43, Fis28d, Fis62f]. report [Fis49e]. Representative [Ney34]. reproduction [DF22]. Rereading [Sav76, Sti76]. Research [Ano45, Ans55, Fis25d, Fis28j, Fis29g, Fis29f, Fis30k, Fis32h, Fis34l, Fis36f, Fis38f, FY38, Fis41i, FY43, Fis44, Fis46c, Fis48e, FY48, FY49, Fis50f, FY53a, FY53b, Fis54c, Fis56b, FY57a, FY57b, Fis58h, Fis63, FY63, Fis67, Fis70a, Fis70b, Fis73b, Fis92c, Fre83, Gro30, Hot27, I.26, I.29, Jow56, Lur72, Mat51, May04, Pea92, Tuk52, Y.39, Fis47d, O’D90, B.58, K.39, Tip53, Toc58, Wet65]. resemblance [Fis25c]. Response [Fis28i, EF29, FM23]. result [FD48]. résultat [FD48]. Results [Ano52, FW30, Fis31c, Fis34k, Pil07]. Retrospect [Fis54f, Fis58i]. Retrospective [Bar90, Lin90]. Review [A.52, All86, Ano45, Ans55, B.50, B.58, Bar92, Bar57, Bin89, Cal87, Dis91, Edw87, Fis22g, Fis31e, Fis35a, Gri82, H.31, Hea67, Hea73, Hot27, I.26, I.29, I.40, J.79, K.39, K.73, Kan81, Kem93, Lur72, Mar80, Pea75, Pea51, Sen91, Smi66, Spi59, Toc58, Tuk52, Wet65, Y.39, Yat79, Yat82, Pre90]. Reviews [Cra36, Fis42d, Fis51c, Fis55d, Fis56a, Gro30, Jow56, Jow57, Tip53]. Revisited [Qui21]. revisiting [Pil07]. Revolution [KDH87, KGM87, You51]. Revolutionized [Sal01]. Rh [FR46]. Rhesus [FRT44, Fis46b, Fis47a, Fis74]. rise [Stu92]. Rivals [Pea79]. Role [Ben91, Fis62b]. Romanovsky [Fis25e]. Ronald [Ano62, Ans55, B.58, Bar56, Bar57, Bil14, Edw92, Edw93, Fin64b, Hea67, IBM^{+}63, KM13, Mah38, Mah64a, Ney56, Owe62, Rao64, Ski07, Smi66, Spi59, Wel56, Yat64b, You62, Ano64, Hal10, Jow56, Jow57, Ken63, Ken70, Moo07b, Pil07, Rao07, Rao08, Ros17, Sti06, Tip53, YM63]. rose [Fis38b]. Rosy [FF55]. Rothamsted [Fis24g, Fis33b]. rott [DFF40]. Rough [Tsu02]. Roy [NP76]. Royal [Ano12, Fis49e]. rumicis [DF22]. ruolo [Fis62b].

S [Fis31e, Fis37b, Pea68, Pea70, Pea01, Sar95, Yat82, Tab08]. salicaria [FM40, Fis41j, FM42, FM43, FM47, Fis49b, FF55]. Sample [Fis21a, FT28, Fis40a]. sampled [Fis34b]. Samples [Ano15, Box87, Fis15c, Fis30h, PL12, SYC^{+}17, Fis39a, Fis61d]. Sampling [FT96, Fis28f, Fis29a, Fis31f, Fis42a, Fis49f, Fis61c, Ney34, Fis39e]. Sc.D. [Owe62]. Scandinavian [FT40]. School [TS15, Stu89, Stu92, Tur87]. Schule [Stu89]. Science [Fis20c, Fis33b, Fis55c, GSP^{+}89, Rup07, Sal01]. Sciences [Fis59g, KGM87, Fis58d, Fis59a, Fis60a]. Scientific [Bar57, Dis91, Fis56d, Fis56e, Fis59f, Fis61e, Fis62c, Fis73a, Fis74, FBY90, Jow57, Fis47d, Fis55e, Fis62b, Fis65a]. scientifica [Fis62b]. scientifique [Fis47d]. scientist [Box78, Cal87, J.79, Kan81, Yat79, Gri82]. Sclerotics [PBFN28]. Scoring [Fis24e, Fis46d]. Scottish [FT40]. sea [Fis38i]. sea-level [Fis38i]. Second [Edw93, Fis35a]. secondary [MF51]. Sediments [FO24]. Seidenfeld [Mar80]. Selected [All86, Bar92, Ben83, Ben90, Edi90, Kem93, Sen91]. Selection [All86, Ben83, Edw87, Fis30e, Fis30d, Fis30i, Fis32g, Fis33g, Fis34g, Fis50c, Fis58b, Fis58f, Fis59c, Fis99, Ney34, WTRS^{+}36, Fis49g, Fis54f, Fis58g, Fis58i, O’D90, Ros17, H.31]. Selective [Fis36e, Fis39f, Fis35p]. Self [Fis47e, Fis61a, Fis62f]. Self-Sterility [Fis47e, Fis61a, Fis62f]. selfed [FB28]. Sequence [Fis26f]. Sequential [Fis52a]. Series [Fis32f, Fis50e, Fis64c, Ken48, KP77]. Serum [Fis45a]. Set [PL12, Tsu02, Fis61c]. Sets [Yat64a]. Several [Fis24f]. Sewall [Fis28d, FF50, Pro92, Ski00, Ros17]. Sex [FF43, Fis53c, FF28, Fis36g, FLO47, Fis62a, Moo95]. sexual [Fis15b, Fis22b, O’D90]. Sheltering [Fis35q]. Sheppard [Fis37a]. shifting [Ros17]. Short [FM47]. Short-style [FM47]. shown [Fis37d]. Sieve [Fis29e]. Significance [Bar56, Ber43, Fis29h, Fis35i, Fis35o, Fis43d, Fis50e, FH56, Fis64c, Joh87, MF48, ZM08, Fis36g, Fis40d, Fis41a, FB49, Fis56c, Moo95]. silicate [Fis29d]. similarity [Fis40d]. Simpler [FW31]. simultaneous [Fis34m, Fis62g]. single [Fis34e]. Sir [Edw93, Ano62, Bar56, Bil14, Edw92, Fin64b, IBM^{+}63, KM13, Ney56, Owe62, Rao64, Ski07, Wel56, Yat64b, You62]. Sittengemälde [Stu89]. situation [Fis32c]. Small [Ano15, Box87, Fis21a, SYC^{+}17]. Smallest [FT28]. Smoking [Fis58j, Fis59d, Fis57c, Spi59]. Snail [FD34].

Soc [NP76]. Social [Fis32g, Fis35i, TS15]. Society [Fis12a, Ano12, Bil14]. sodium [Fis29d]. Soil [Fis28e, TF27, Fis26e]. soils [Fis34c, TG34]. Solution [BWF33, Fis25e, Fis41e]. solutions [Fis40b, Fis41g]. solve [Fis26d]. Some [BWF33, Fis14, Fis21b, Fis34k, Fis41h, Fis59d, Fis60b, Fis62d, Fre83, Ken48, Mah64b, Pea79, Rac64, Spi59, Wel56, Wil64, Fis27c, Fis39e, Fis38e, Pea68, Pea70, Pea01]. sometimes [Hal10]. Sons [Gri82]. Sort [Fis48f]. Soviet [Fis49c]. Space [FM57]. Spanish [Fis55b]. Spastic [BCF39, FBC39]. Special [Fis34k, Fis22b, Fis46d, Pie90]. Species [FF26, Fis30c, Fis43f, FF28]. Specification [CF38, Leh90]. Spencer [Fis32g]. sphere [Fis53e]. Spherical [Bin89]. Spontaneous [FM47, Fis36c]. sporting [Fis33g]. spread [FF47]. Square [Fis20b, Ull99, Cam95, Fis41h]. Squares [Ull99, FY34, Fis41c]. squeezed [FNS84]. Stage [Fis39g]. Standard [Ano15, Fis51d, ZM08]. Standpoint [Fis59c]. Stanford [Sti99a]. Stanley [Fis50b]. State [RS00, WTRS^{+}36]. Statement [Fis28a]. states [FNS84, Fis35i]. Statistical [Ano12, Bar65, Ben90, Ben91, BWF33, Fis22h, Fis23c, Fis23d, Fis25d, Fis25f, Fis26h, Fis28j, Fis29f, FW30, Fis30k, Fis31c, Fis32h, Fis34l, Fis35d, Fis35r, Fis36f, Fis38f, FY38, Fis38g, Fis41i, FY43, Fis44, Fis45c, Fis46c, Fis48e, FY48, Fis49f, FY49, Fis50f, Fis52b, FY53a, FY53b, Fis54c, Fis55e, Fis56d, Fis56e, FY57a, FY57b, Fis58h, Fis59f, Fis63, FY63, Fis67, Fis70a, Fis70b, Fis73b, Fis73a, FBY90, Fis92c, Gro30, I.40, Inm94, Jow56, Jow57, Len06, Mar80, Mat51, NP28, PFI94, PB10, Sei92a, Sti99b, Tip53, Tsu02, Tuk52, ZM08, Con92, Fis25c, Fis27c, Fis35k, Fis38i, Fis38h, Hal10, Fis47d, Fis56b, Pea92, Hot27, Ano45, Ans55, B.58, Bar92, Bar57, Bin89, Dis91, I.26, I.29, K.39, Kem93, Lur72, Sen91, Toc58, Wet65, Y.39]. Statistician [Hea03, Sti06]. Statisticians [Edi90, Fis33h, Gan82]. Statistics [Fin64b, Fis24f, Fis29g, FIT32, Fis33b, Fis42a, Fis50g, Fis50h, Fis53b, Fis53d, Fis92b, Gup60, Hal98, Hot51, KP77, KJ92a, KJ92b, KJ97, PK70, Pea01, Pil07, Rao92, RS00, Sal01, Sti99a, Sti99b, Wil64, Bar63, Fis22j, Fis38e, Fis39e, Fis43a, Fis59b, Hal90, Leh11, Pea68, Rao00, Sti73, Str90, Stu89, Stu92, Tho90, Fis55b, Pea51, Gei92]. Statistik [Stu89]. statistique [Fis38e]. statistiques [Fis47d]. Statistische [Fis56b]. Sterilisation [BFG^{+}30]. Sterility [Fis47e, Fis61a, Fis62f]. Steven [Fis40d]. Stocks [Fis33g]. Stratified [Ney34]. Streit [Stu89]. strength [Fis53g]. Strip [Fis36b]. strongly [FY44]. Student [Ano15, Fis39h, SYC^{+}17]. Students [Fis20c, Fis25a, Fis25b]. studied [FG37]. Studies [EF29, Fis21c, FM23, Fis24g, Fis27e, KP77, Pea68, PK70, Pea01, Sti73, DF22, Fis35m]. Study [Aro41, FIT32, Fis74, Aro40, DFF40, DFF49, Fis24a, Fis54b]. Style [FM48, Fis41j, FM43, FM47]. Style-Length [FM48]. Sub [Yat64a, Fis49f]. Sub-Commission [Fis49f]. Sub-Sets [Yat64a]. Subroutine [MP86a, MP86b]. Substitute [GV99]. substitution [Fis41b]. sufficiency [Sti73]. Supplement [Fis38i]. supplied [Fis34b]. supposed [BF36]. supposition [Fis19a, MS66, NP76]. Surnames [FV39]. surveys [Fis26g]. Survivors [Fis35b]. suspicious [Har08]. Symphalangism [BFP51]. Symposium [RS00]. System [Fis36b, Fis43e, Fis46d, Fis49g, Fis51d]. systematic [BF36, Fis22i]. systems [ACF^{+}49].

T [Bin89, Mar80]. Table [Fis26b, Sti99b]. Tables [B.58, Bar56, CaFJ93, Fis22f, FH28, Fis33h, FY38, FY43, Fis45e, FY48, FY49, FY53a, FY53b, FH56, FY57a, FY57b, FY63, K.39, MP86a, MP86b, Tip53, Toc58, Wet65, Wis27, Fis35l, Fis41a, Fis41d, Fis22h, Fis26h, Fis56c]. Taste [FFH39]. Taste-testing [FFH39]. Tasting [Sal01]. tax [Fis28g, Fis36d]. taxonomic [Fis36j]. Taylor [Fis45b]. Tea [Sal01]. Test [Bar56, Bas80a, Bas80b, CaFJ93, Fis45e, FH56, Hin80, Kem80, Lan80b, Lin80, MP86a, MP86b, NP28, Rub80, Tsu02, Yat64a, BF36, Cam95, FM36a, Fis40d, Fis41a, FB49, Fis49d, Fis56c]. Testing [Ber03, Fis46e, Leh93, FFH39]. Tests [Ber43, Fis24e, Fis29h, Fis35o, Fis35r, Fis36g, Fis43d, Fis51b, Inm94, Joh87, MF48, PFI94, Fis23d, Fis40a]. tetrasomic [Fis49b]. texanus [Fis39f]. their [BWF33, Hal90]. Theorem [Fis26b, GV99, Ald08, Edw90a, Sei92b]. theorems [Fis41h]. Theoretical [Fis43f, Fis92b, Gei92, Fis22j, Fis41j, Fis49g]. Theories [Fis32a, Leh93, Ros17]. Theory [B.50, FO24, Fis25f, Fis30d, Fis47c, Fis49i, Fis49j, Fis58b, Fis64b, Fis65b, Fis99, H.31, I.40, Kar92, KJ92a, Leh93, Ski00, WTRS^{+}36, Fis27c, Fis34i, Fis35p, Fis38g, Fis41k, Fis42c, Fis43a, Fis47h, Fis48d, Fis54g, Fis54f, Fis58i, O’D90]. therapeutic [AF44]. Third [FIT32]. Thomas [Dal91, Dal99, FdB47]. Thornton [Fis34c]. those [FW31]. thought [Fis61e]. Thoughts [Fre83]. three [FB49]. three-point [FB49]. thyrotoxicosis [MF51]. time [Stu89, Stu92]. Together [Fis31f]. Top [Fis27e]. toxic [MF51]. Trans [NP76]. Transformations [Fis54d, Fis54e]. transmission [Fis59e]. travel [FM57]. Treasury [BF35, BCF39, BF43, BFP51, FB34, FBC39, PFSB25, PBFN28, PBFN31]. Treatise [Fis22g]. Treatment [Fis27a, Fis38i]. tri [Fis30j]. tri-colour [Fis30j]. Trials [Fis27d, Fis51a]. Triplet [Fis28k]. true [Bar87]. Tuberculins [Fis49a]. twelfth [FS48]. Twentieth [Rup07, Sal01]. Twentieth-century [Rup07]. Twice [Edw03, Edw90b]. twins [Fis19c, Fis20a, Fis22d, Fis25c]. Two [Fis28l, FW31, Fis34o, Fis46e, Leh93, Ney34, Fis43e, Fis61d, Hal10]. Two-Way [FW31]. Type [BF35, Fis28i]. Types [BF43, Fis41h].

unabhängigen [FY44]. Uncertain [Fis36i, Fis34h]. Underworld [Fis57e]. Undulated} [Fis49d]. Unequal [Fis46e, DF37, Fis39a]. unexpected [FD48]. unfit [BFG^{+}30]. United [Fis49f, Fis35i]. University [Edw93, RS00, Sen91]. unknown [Fis26i, Fis33a, Fis61d]. Unordered [MP86a, MP86b]. unrelated [Fis40a]. Unsigned [Fis42d]. upon [Fis34f]. USA [RS00]. Use [AF44, NP28, Fis32e, Fis34m, Fis36j]. Used [Fis35o]. using [Fis62a]. utilization [Fis38h].

V [Fis25e, Yat82]. valuation [Fis23c]. Value [Fis27e, Wis27]. Values [Fis15c, Fis62a]. Variability [FF26, DF37, Fis19b, FF28, Fis37d]. variable [Fis45d]. Variance [Fis54d, Fis54e, FG37, Fis39g, Fis55a, Fis61d]. variances [Fis39a]. variates [DF37]. Variation [EF29, Fis21c, FM23, Fis27e, Fis24g, Fis53g]. varieties [DF22, FM23]. Various [Fis54d, Fis54e]. vector [Fis13]. Verification [FM36b]. VI [EF29]. viability [FB49]. ’Viceroy [Fis34n]. vices [Fis22b]. Vicia [DF22]. Views [Leh90]. viii [Fis35a]. Vitamin [AF44]. Vol [Fis57b, KDH87, KGM87]. Volume [Hea73, K.73, KP77, KJ92a, KJ92b, KJ97, Pea75, PBFN28, PBFN31, PFSB25, PBFN28, PBFN31].

W [Fis35a]. W. [Dem93, Fis26e, Fis37a, Pea68, Pea70, Pea01]. Wagner [Fis35l]. Wagner-Manslau [Fis35l]. wave [Fis37e]. Way [FW31]. Weekly [FM22]. weighted [Fis61d]. Welch [Fis57b]. Well [Edw03, Fis24f, Edw90b]. Well-Known [Edw03, Edw90b]. wheat [Fis24g]. Which [Fis24b]. Whole [Fis47b]. wholly [Fis40a]. Wild [Fis28i, Fis39f, Fis49b, Fis61b]. Wiley [Gri82]. Wissenschaft [Fis56b]. without [Fis34e]. Work [Gre03, Mat64, Owe62, Fis36k, Fis37a, Fis08, Rao07, Rao08]. Workers [Fis25d, Fis28j, Fis30k, Fis32h, Fis34l, Fis36f, Fis38f, Fis41i, Fis44, Fis46c, Fis48e, Fis50f, Fis54c, Fis56b, Fis58h, Fis63, Fis67, Fis70a, Fis70b, Fis73b, Fis92c, Mat51, Pea92, Tuk52, Ano45, Ans55, Gro30, I.26, I.29, Jow56, Lur72, Y.39, Hot27]. world [Ros17]. Wright [FF50, Fis28d, Fis31e, Fis34i, Har08, Mus12, Pro92, Ros17, Ski00].

xii [Gri82]. xviii [Sen91]. XX [Pea68, Pea01]. XXXII [Sti73].

Yates [B.58, Fis35c, K.39, Tip53, Toc58, Wet65]. Years [Bar65, TC01]. Yield [Fis27d, Fis21c, Fis24g, Fis29d]. Yielding [Fis24f]. York [Gri82].

Zero [Fis35b]. zwischen [Stu89].

References

A:1952:BRBc

[A.52] R. L. A. Book review: Creative Aspects of Natural Law by R. A. Fisher. Philosophy of Science, 19(4):350–351, October 1952. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

           URL http://www.jstor.org/stable/185409.

Andresen:1949:NLL

[ACF^{+}49] H. Andresen, S. T. Callender, R. A. Fisher, R. Grubb, W. T. J. Morgan, A. E. Mourant, M. M. Pickles, and R. R. Race. A notation for the Lewis and Lutheran blood-group systems. Nature, 163 (4145):580–581, April 9, 1949. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL https://www.nature.com/

articles/163580a0.

Atkins:1944:TUV

[AF44] W. R. G. Atkins and R. A. Fisher. The therapeutic use of vitamin C. Journal of the Royal Army Medical Corps, 83(5):251–252, November 1944. CODEN JRAMAI. ISSN 0035-8665 (print), 2052-0468 (electronic).

Aldrich:1995:RFM

[Ald95] John Aldrich. R. A. Fisher and the making of maximum likelihood, 1912–22. Discussion Papers in Economics and Econometrics 9504, Department of Economics, University of Southampton,

Southampton, UK, 1995. URL http://eprints.soton.ac.uk/
id/eprint/33249.

Aldrich:1997:RFM

[Ald97] John Aldrich. R. A. Fisher and the making of maximum likelihood 1912–1922. Statistical Science, 12(3):162–176, August 1997. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1030037906.

Aldrich:2005:FR

[Ald05] John Aldrich. Fisher and regression. Statistical Science, 20(4): 401–417, November 2005. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1137076660.

Aldrich:2008:RFB

[Ald08] John Aldrich. R. A. Fisher on Bayes and Bayes’ theorem. Bayesian Analysis, 3(1):161–170, March 2008. CODEN ???? ISSN 1931-6690 (print), 1931-6690 (electronic). URL http://ba.

stat.cmu.edu/journal/2008/vol03/issue01/aldrich.pdf;
http://projecteuclid.org/euclid.ba/1340370565.

Aldrich:2013:GRF

[Ald13] John Aldrich. A guide to R. A. Fisher. University of Southampton Web site, 2013.

Allen:1986:BRB

[All86] Garland E. Allen. Book review: Natural Selection, Heredity, and Eugenics; Including Selected Correspondence of R. A. Fisher with Leonard Darwin and Others by J. H. Bennett. Isis, 77(1):168–169, March 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/232557.

Anderson:1996:RFM

[And96] T. W. Anderson. R. A. Fisher and multivariate analysis. Statistical Science, 11(1):20–34, February 1996. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1032209662.

Anonymous:1915:EDS

[Ano15] Anonymous. Editorial: On the distribution of the standard deviations of small samples: Appendix I. To papers by “Student” and R. A. Fisher. Biometrika, 10(4):522–529, May 1915. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2331839.

Anonymous:1945:BRB

[Ano45] Anonymous. Book review: Statistical Method for Research Workers, by R. A. Fisher. Journal of the Royal Statistical Society, 108 (1/2):235, ???? 1945. ISSN 0952-8385. URL http://www.jstor.

org/stable/2981200.

Anonymous:1952:RCR

[Ano52] Anonymous. The Race Concept: Results of an Inquiry. UNESCO,

Paris, France, 1952. 108 pp. URL https://unesdoc.unesco.org/
ark:/48223/pf0000073351.

Anonymous:1960:EFG

[Ano60] Anonymous. Errata to Fisher and Gupta. Technometrics, 2(4): 523–524, November 1960. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/ stable/1266462. See [FC60, Gup60].

Anonymous:1962:SRA

[Ano62] Anonymous. Sir Ronald Aylmer Fisher, 1890–1962. Journal of the Royal Statistical Society. Series A (General), 125(4):668, ???? 1962. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2982658.

Anonymous:1964:MRA

[Ano64] Anonymous. In memoriam: Ronald Aylmer Fisher, 1890–1962. Biometrics, 20(2):i–iii, June 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2528390.

Anonymous:2012:RSS

[Ano12] Anonymous. The Royal Statistical Society presents: A 50th anniversary commemoration of R. A. Fisher (17 February 1890– 29 July 1962) 7 June 2012. Conference Program from the University of Southampton, UK., 2012. URL http://www.

economics.soton.ac.uk/staff/aldrich/fisherguide/RSS_R%
20A%20_Fisher_meeting.pdf.

Anonymous:2013:RFD

[Ano13] Anonymous. R. A. Fisher Digital Archive. The University of Ade-

laide Web site, 2013. URL http://digital.library.adelaide.
edu.au/dspace/handle/2440/3860.

Anscombe:1955:BRB

[Ans55] F. J. Anscombe. Book review: Statistical Methods for Research Workers, by Ronald A. Fisher. Journal of the Royal Statistical Society. Series A (General), 118(4):486–487, ???? 1955. CODEN

JSSAEF. ISSN 0035-9238. URL http://www.jstor.org/stable/
2342720.

Aroian:1940:SRF

[Aro40] Leo Avedis Aroian. A study of R. A. Fisher’s z-distribution and the related F distribution. Ph.D. thesis, University of Michigan, Ann Arbor, MI, USA, 1940. 26 pp. URL https://search.proquest.

com/docview/301790576.

Aroian:1941:SRF

[Aro41] Leo A. Aroian. A study of R. A. Fisher’s z distribution and the related F distribution. Annals of Mathematical Statistics, 12(4): 429–448, December 1941. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177731681.

Aylward:2021:RFE

[Ayl21] Alex Aylward. R. A. Fisher, eugenics, and the campaign for family allowances in interwar Britain. British Journal for the History of Science, 54(4):485–505, December 2021. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL https:

//www.cambridge.org/core/journals/british-journal-for-
the-history-of-science/article/ra-fisher-eugenics-and-

the-campaign-for-family-allowances-in-interwar-britain/
BD28704590AC49EAD4B6FC8F4471C455.

B:1950:BRBa

[B.50] M. S. B. Book review: The Theory of Inbreeding, by R. A. Fisher. Journal of the Royal Statistical Society. Series A (General), 113 (2):249–250, ???? 1950. CODEN JSSAEF. ISSN 0035-9238. URL

http://www.jstor.org/stable/2981045.

B:1958:BRBc

[B.58] W. R. B. Book review: Statistical Tables for Biological, Agricultural & Medical Research, by Ronald A. Fisher; F. Yates. The Incorporated Statistician, 8(4):193–194, July 1958. CODEN ????

ISSN 1466-9404. URL http://www.jstor.org/stable/2986384.

Bartlett:1956:CSR

[Bar56] M. S. Bartlett. Comment on Sir Ronald Fisher’s paper: “On a Test of Significance in Pearson’s Biometrika Tables (No. 11)”. Journal of the Royal Statistical Society. Series B (Methodological), 18(2): 295–296, ???? 1956. CODEN JSTBAJ. ISSN 0035-9246. URL

http://www.jstor.org/stable/2983717. See [Fis56c] and com-

ment [Fis57b].

Bartlett:1957:BRB

[Bar57] M. S. Bartlett. Book review: Statistical Methods and Scientific Inference by Ronald A. Fisher. Biometrika, 44(1/2):293–295, June 1957. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2333276.

Barnard:1963:FCM

[Bar63] G. A. Barnard. Fisher’s contributions to mathematical statistics. Journal of the Royal Statistical Society. Series A (General), 126 (1):162–166, ???? 1963. CODEN JSSAEF. ISSN 0035-9238.

Bartlett:1965:RFL

[Bar65] M. S. Bartlett. R. A. Fisher and the last fifty years of statistical methodology. Journal of the American Statistical Association, 60 (310):395–409, June 1965. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2282678.

Barnard:1987:RFT

[Bar87] G. A. Barnard. R. A. Fisher — a true Bayesian? International Statistical Review = Revue Internationale de Statistique, 55 (2):183–189, August 1987. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.jstor.org/

stable/1403194.

Barnard:1989:RFH

[Bar89] G. A. Barnard. [R. A. Fisher on the history of inverse probability]: Comment. Statistical Science, 4(3):258–260, August 1989. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1177012490.

Barnard:1990:FR

[Bar90] George Barnard. Fisher: a retrospective. Chance, 3(1):22–28, Winter 1990. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Barnard:1992:RSI

[Bar92] George A. Barnard. Review of Statistical Inference and Analysis: Selected Correspondence of R. A. Fisher. Statistical Science, 7(1):5–12, February 1992. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1177011440.

Barnard:1995:NBF

[Bar95] G. A. Barnard. Neyman and the Behrens–Fisher problem — an anecdote. Probability and Mathematical Statistics, 15(??):67–71, ???? 1995. ISSN 0208-4147 (print), 2300-8113 (electronic).

Basu:1980:RAEa

[Bas80a] D. Basu. Randomization analysis of experimental data: The Fisher randomization test. Journal of the American Statistical Association, 75(371):575–582, September 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2287648.

Basu:1980:RAEb

[Bas80b] D. Basu. Randomization analysis of experimental data: The Fisher randomization test rejoinder. Journal of the American Statistical Association, 75(371):593–595, September 1980. CODEN JSTNAL.

ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2287654.

Bell:1939:THI

[BCF39] Julia Bell, E. Arnold Carmichael, and R. A. Fisher, editors. The Treasury of Human Inheritance: Nervous Diseases and Muscular Dystrophies. Part III. On Hereditary Ataxia and Spastic Paraplegia. Cambridge University Press, Cambridge, UK, 1939. 141–281 pp. LCCN ????

Bennett:1974:CPR

[Ben74] J. H. Bennett, editor. Collected papers of R. A. Fisher. University of Adelaide, Adelaide, South Australia, Australia, 1974. ISBN 0-909688-00-1 (vol. 1), 0-909688-01-X (vol. 2), 0-909688-02-8 (vol. 3), 0-909688-03-6 (vol. 4), 0-909688-04-4 (vol. 5). 604 or 668 or 575 (vol. 1), ??? (vol. 2), 560 (vol. 3), 668 (vol. 4), 575 (vol. 5) pp.

LCCN QA276.A12 F57. URL https://wellcomelibrary.org/

item/b18032357. Volume 1: 1912–1924. Volume 2: 1925–1931. Volume 3: 1932–1936. Volume 4. 1937-1947. Volume 5. 1948–1962.

Bennett:1983:NSH

[Ben83] J. H. Bennett, editor. Natural Selection, Heredity, and Eugenics: including Selected Correspondence of R. A. Fisher with Leonard Darwin and Others. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1983. ISBN 0-19-858177-7. vii + 306 pp. LCCN QH31.F56 F57 1983.

Bennett:1990:SIA

[Ben90] J. H. Bennett, editor. Statistical inference and analysis: Selected Correspondence of R. A. Fisher. Oxford Science Publications. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1990. ISBN 0-19-855552-0. xviii + 380 pp.

Bennett:1991:RFR

[Ben91] J. H. Bennett. R. A. Fisher and the role of a statistical consultant. Journal of the Royal Statistical Society. Series A (Statistics in Society), 154(3):443–445, ???? 1991. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2983153.

Berkson:1943:ETS

[Ber43] Joseph Berkson. Experience with tests of significance: a reply to Professor R. A. Fisher. Journal of the American Statistical Association, 38(222):242–246, June 1943. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.

org/stable/2279546.

Berger:2003:CFJ

[Ber03] James O. Berger. Could Fisher, Jeffreys and Neyman have agreed on testing? Statistical Science, 18(1):1–32, February 2003. CO-DEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1056397485.

Bell:1935:THI

[BF35] Julia Bell and Ronald Aylmer Fisher, editors. The Treasury of Human Inheritance: Nervous Diseases and Muscular Dystrophies. Part II On the Peroneal Type of Progressive Muscular Atrophy. Cambridge University Press, Cambridge, UK, 1935. 69–139 pp. LCCN ????

Barbacki:1936:TSP

[BF36] S. Barbacki and R. A. Fisher. A test of the supposed precision of systematic arrangements. Annals of Eugenics, 7(2):189– 193, September 1936. CODEN ???? ISSN 2050-1420 (print),

2050-1439 (electronic). URL https://onlinelibrary.wiley.
com/doi/abs/10.1111/j.1469-1809.1936.tb02138.x.

Bell:1943:THI

[BF43] Julia Bell and R. A. Fisher, editors. The Treasury of Human Inheritance: Nervous Diseases and Muscular Dystrophies. Part 4: On Pseudohypertrophic and Allied Types of Progressive Muscular Dystrophy. Cambridge University Press, Cambridge, UK, 1943. 284–381 pp. LCCN ????

Bliss:1953:FNB

[BF53a] C. I. Bliss and R. A. Fisher. Fitting the negative binomial distribution to biological data. Biometrics, 9(2):176–200, June 1953. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/3001850.

Bliss:1953:NEF

[BF53b] C. I. Bliss and R. A. Fisher. Note on the efficient fitting of the negative binomial. Biometrics, 9(2):197–200, June 1953. CODEN BIOMB6. ISSN 0006-341x (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/3001850.

Blacker:1930:SU

[BFG^{+}30] C. P. Blacker, R. A. Fisher, R. A. Gibbons, R. Langdon-Down, J. A. Ryle, and C. Collyer. Sterilisation of the unfit. The Lancet (London, England), 216(5577):161, 1930. CODEN LANCAO. ISSN 0140-6736 (print), 1474-547x (electronic).

Bell:1951:THI

[BFP51] Julia Bell, Ronald Aylmer Fisher, and Karl Pearson, editors. Treasury of Human Inheritance. On Hereditary Digital Anomalies. Part 1. On Brachydactyly and Symphalangism. Cambridge University Press, Cambridge, UK, 1951. 31 + 12 pp. LCCN ????

Billard:2014:SRF

[Bil14] Lynne Billard. Sir Ronald A. Fisher and the International Biometric Society. Biometrics, 70(2):259–265, June 2014. CODEN BIOMB6. ISSN 0006-341x (print), 1541-0420 (electronic).

Bingham:1989:BRB

[Bin89] Christopher Bingham. Book review: Statistical Analysis of Spherical Data by N. I. Fisher; T. Lewis; B. J. J. Embleton. Technometrics, 31(4):494–495, November 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1270013.

Bliss:1964:RFC

[Bli64] C. I. Bliss. R. A. Fisher’s contributions to medicine and bioassay. Biometrics, 20(2):273–285, June 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://

www.jstor.org/stable/2528396.

Box:1978:RFL

[Box78] Joan Fisher Box. R. A. Fisher, the life of a scientist. Wiley series in probability and mathematical statistics. John Wiley, New York, NY, USA, 1978. ISBN 0-471-09300-9, 0-471-83898-5 (paperback). xii + 512 + 13 pp. LCCN QA29.F57 B68.

Box:1980:RFD

[Box80] Joan Fisher Box. R. A. Fisher and the design of experiments, 1922–1926. The American Statistician, 34(1):1–7, February 1980. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/stable/2682986.

Box:1987:GGF

[Box87] Joan Fisher Box. Guinness, Gosset, Fisher, and small samples. Statistical Science, 2(1):45–52, February 1987. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177013437.

Box:2005:RRF

[Box05] Joan Fisher Box. A reminiscence of R. A. Fisher. The American Statistician, 59(4):312–314, November 2005. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Bowman:2007:BDM

[BS07] K. O. Bowman and L. R. Shenton. The beta distribution, moment method, Karl Pearson and R. A. Fisher. Far East Journal of Theoretical Statistics, 23(2):133–164, ???? 2007. ISSN 0972-0863.

Buchanan-Wol:1933:SMS

[BWF33] H. J. Buchanan-Wol and R. A. Fisher. Some modern statistical methods: their application to the solution of herring race problems. ICES Journal of Marine Science, 8(1):7–8, April 1933. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL

http://academic.oup.com/icesjms/article/8/1/7/682457.

Clarkson:1993:RAF

[CaFJ93] Douglas B. Clarkson, Yuan an Fan, and Harry Joe. A remark on Algorithm 643: FEXACT: An algorithm for performing Fisher’s exact test in r × c contingency tables. ACM Transactions on Mathematical Software, 19(4):484–488, December 1993. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (elec-

tronic). URL http://www.acm.org/pubs/citations/journals/
toms/1993-19-4/p484-clarkson/. See [MP86a].

Calder:1987:BRB

[Cal87] Allan B. Calder. Book review: R. A. Fisher: The Life of a Scientist, by R. A. Fisher; Joan Fisher Box. Journal of the Royal Statistical Society. Series D (The Statistician), 36(1):60–62, ???? 1987. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/stable/2988278.

Camilli:1995:RBF

[Cam95] Gregory Camilli. The relationship between Fisher’s exact test and Pearson’s chi-square test: a Bayesian perspective. Psychometrika, 60(2):305–312, ???? 1995. CODEN PSMIA3. ISSN 0033-3123 (print), 1860-0980 (electronic).

Cornish:1938:MCS

[CF38] E. A. Cornish and R. A. Fisher. Moments and cumulants in the specification of distributions. Revue de l’Institut international de statistique = Review of the International Statistical Institute, 5(4): 307–320, January 1938. ISSN 0373-1138.

Conniffe:1992:KPS

[Con92] D. Conniffe. Keynes on probability and statistical inference and the links to Fisher. Cambridge Journal of Economics, 16(??):475– 489, ???? 1992.

Cornish:1964:FAA

[Cor64] E. A. Cornish. Fisher’s activities in Australia, 1959–62. Biometrics, 20(2):372–373, June 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2528405.

Craig:1936:RPRa

[Cra36] C. C. Craig. Recent publications: Reviews: The Design of Experiments, by R. A. Fisher. American Mathematical Monthly, 43 (3):180–181, March 1936. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Crow:1990:FCG

[Cro90] James F. Crow. Fisher’s contributions to genetics and evolution. Theoretical Population Biology, 38(3):263–275, ???? 1990. CO-DEN TLPBAQ. ISSN 0040-5809 (print), 1096-0325 (electronic).

Dale:1991:HIP

[Dal91] Andrew I. Dale. A history of inverse probability: from Thomas Bayes to Karl Pearson, volume 16 of Studies in the history of mathematics and physical sciences. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1991. ISBN 0-387-97620-5 (New York), 3-540-97620-5 (Berlin). xx + 495 pp. LCCN QA279.5 .D35 1991.

Dale:1999:HIP

[Dal99] Andrew I. Dale. A history of inverse probability: from Thomas Bayes to Karl Pearson. Studies in the history of mathematics and physical sciences. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 1999. ISBN 0-387-98807-6. xxiv + 670 pp.

LCCN QA279.5 .D35 1999. URL http://www.loc.gov/catdir/
enhancements/fy0816/99018596-d.html;   http://www.loc.
gov/catdir/enhancements/fy0816/99018596-t.html.

Deming:1993:WED

[Dem93] W. Edwards (William Edwards) Deming. W. Edwards Deming papers, 1795–1994 (bulk 1930–1993). Library of Congress collection, 1993. 58,000 items. 164 containers plus 1 classified and 2 oversize. 70 linear feet. Includes correspondence with Sir Ronald Aylmer Fisher.

Davidson:1922:BSA

[DF22] J. Davidson and R. A. Fisher. Biological studies of Aphis rumicis linn. reproduction on varieties of Vicia faba. Annals of Applied Biology, 9(2):135–142, June 1922. CODEN AABIAV. ISSN 0003-4746 (print), 1744-7348 (electronic). URL https:

//onlinelibrary.wiley.com/doi/abs/10.1111/j.1744-7348.
1922.tb05943.x.

Day:1937:CVP

[DF37] B. Day and R. A. Fisher. The comparison of variability in populations having unequal means. an example of the analysis of covariance with multiple dependent and independent variates. Annals of Eugenics, 7(4):333–348, June 1937. CODEN ???? ISSN 2050-1420 (print), 2050-1439 (electronic). URL https://onlinelibrary.

wiley.com/doi/abs/10.1111/j.1469-1809.1937.tb02151.x.

Dowdeswell:1940:QSP

[DFF40] W. H. Dowdeswell, R. A. Fisher, and E. B. Ford. The quantitative study of populations in the Lepidoptera I. Polyommatus icarus rott. Annals of Eugenics, 10(1):123–136, January 1940. CODEN ???? ISSN 2050-1420 (print), 2050-1439 (electronic).

URL https://onlinelibrary.wiley.com/doi/abs/10.1111/j.
1469-1809.1940.tb02242.x.

Dowdeswell:1949:QSP

[DFF49] W. H. Dowdeswell, R. A. Fisher, F.R.S., and E. B. Ford, F.R.S. The quantitative study of populations in the Lepidoptera. Heredity, 3(1), April 1, 1949. CODEN HDTYAT. ISSN 0018-067X (print), 1365-2540 (electronic). URL https://www.nature.com/

articles/hdy19493.

Disney:1991:BRB

[Dis91] John Disney. Book review: Statistical Methods, Experimental Design and Scientific Inference, by R. A. Fisher. Journal of the Royal Statistical Society. Series D (The Statistician), 40(4):462–463, ???? 1991. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/stable/2348748.

Editors:1990:CSS

[Edi90] Editors. Comments from selected statisticians on Fisher. Chance, 3(1):29–32, Winter 1990. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Edwards:1987:BRB

[Edw87] A. W. F. Edwards. Book review: Natural Selection, Heredity, and Eugenics, by R. A. Fisher; Leonard Darwin; J. H. Bennett. Journal of the Royal Statistical Society. Series A (General), 150 (2):168–169, ???? 1987. CODEN JSSAEF. ISSN 0035-9238. URL

http://www.jstor.org/stable/2981635.

Edwards:1990:FFT

[Edw90a] A. W. F. Edwards. Fisher, W̄, and the fundamental theorem. Theoretical Population Biology, 38(3):276–284, ???? 1990. CODEN TLPBAQ. ISSN 0040-5809 (print), 1096-0325 (electronic).

Edwards:1990:RFT

[Edw90b] A. W. F. Edwards. R. A. Fisher: twice professor of genetics: London and Cambridge or “a fairly well-known geneticist”. Biometrics, 46(4):897–904, 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Edwards:1992:CCB

[Edw92] A. W. F. Edwards. Celebration of the centenary of the birth of Sir Ronald Aylmer Fisher, 1990. Historia Mathematica, 19(1):81– 82, February 1992. CODEN HIMADS. ISSN 0315-0860 (print),

1090-249X (electronic). URL http://www.sciencedirect.com/
science/article/pii/031508609290061F.

Edwards:1993:MGF

[Edw93] A. W. F. Edwards. Mendel, Galton, Fisher (Second Sir Ronald Fisher lecture, University of Adelaide, 12th November 1992). Australian Journal of Statistics, 35(2):129–140, 1993. CODEN AU-JSA3. ISSN 0004-9581.

Edwards:1994:RFK

[Edw94] A. W. F. Edwards. R. A. Fisher on Karl Pearson. Notes and Records of the Royal Society of London, 48(1):97–106, January 1994. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (elec-

tronic). URL http://www.jstor.org/stable/531422.

Edwards:1997:WDF

[Edw97] A. W. F. Edwards. What did Fisher mean by “inverse probability” in 1912–1922? Statistical Science, 12(3):177–184, August 1997. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1030037907.

Edwards:2003:RFT

[Edw03] A. W. F. Edwards. R. A. Fisher: Twice Professor of Genetics: London and Cambridge, or ‘a fairly well-known geneticist’. Journal of the Royal Statistical Society. Series D (The Statistician), 52(3): 311–318, ???? 2003. CODEN ???? ISSN 0039-0526 (print), 1467-

9884 (electronic). URL http://www.jstor.org/stable/4128206.

Eden:1929:SCV

[EF29] T. Eden and R. A. Fisher. Studies in crop variation. VI. Experiments on the response of the potato to potash and nitrogen. Journal of Agricultural Science, 19(2):201– 213, April 1929. CODEN JASIAB. ISSN 0021-8596 (print),

1469-5146 (electronic). URL https://www.cambridge.org/
core/journals/journal-of-agricultural-science/article/
studies-in-crop-variation-vi-experiments-on-the-response-
of-the-potato-to-potash-and-nitrogen/3D858B32BEDB50DBE2341A9559C86319.

Efron:1998:RFC

[Efr98] Bradley Efron. R. A. Fisher in the 21st century (invited paper presented at the 1996 R. A. Fisher Lecture). Statistical Science, 13(2):95–122, May 1998. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1028905930.

Efron:2000:RFC

[Efr00] Bradley Efron. R. A. Fisher in the 21st century. In Rao and Székely [RS00], pages 109–144. ISBN 0-8247-9029-4 (hardcover).

LCCN QA276.16 .S844 2000. URL http://www.loc.gov/catdir/
enhancements/fy0647/99057753-d.html.

Franklin:2008:EMF

[F^{+}08] Allan Franklin et al., editors. Ending the Mendel–Fisher controversy. University of Pittsburgh Press, Pittsburgh, PA, USA, 2008. ISBN 0-8229-4319-0 (hardcover), 0-8229-5986-0 (paperback). x + 330 pp. LCCN QH428 .E53 2008. URL http:

//www.loc.gov/catdir/enhancements/fy0805/2007041848-b.
html;   http://www.loc.gov/catdir/enhancements/fy0805/
2007041848-d.html;   http://www.loc.gov/catdir/toc/ecip082/
2007041848.html.

Fisher:1925:NNE

[FA25] R. A. Fisher and P. R. Ansell. Note on the numerical evaluation of a Bessel function derivative. Proceedings of the London Mathematical Society. First Series, 24(??):54–56, ???? 1925. CODEN PLMTAL. ISSN 0024-6115 (print), 1460-244X (electronic).

Fisher:1928:ELO

[FB28] R. A. Fisher and Bhai Balmukand. The estimation of linkage from the offspring of selfed heterozygotes. Journal of Genetics, 20(??):79–92, July 1928. CODEN JOGNAU. ISSN 0022-1333 (print), 0973-7731 (electronic). URL https://link.springer.

com/article/10.1007/BF02983317.

Fisher:1931:PRM

[FB31] R. A. Fisher and S. Bartlett. Pasteurised and raw milk. Nature, 127(3207):591–592, April 18, 1931. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL https://www.nature.

com/articles/127591c0.

Fisher:1934:THI

[FB34] R. A. Fisher and Julia Bell, editors. Treasury of Human Inheritance: Nervous Diseases and Muscular Dystrophies, volume 4 of Eugenics Laboratory Memoirs. Cambridge University Press, Cambridge, UK, 1934. 67 + 12 pp.

Fisher:1949:ELD

[FB49] R. A. Fisher and Norman T. J. Bailey. The estimation of linkage with differential viability. I. Note on the test of significance for differential viability in frequency data from a complete three-point test. Heredity, 3(2):215–219, August 1, 1949. CODEN HDTYAT. ISSN 0018-067X (print), 1365-2540 (electronic). URL https://

www.nature.com/articles/hdy194913.

Fisher:1939:THI

[FBC39] R. A. Fisher, Julia Bell, and E. Arnold Carmichael, editors. The Treasury of Human Inheritance: Part 3: On Hereditary Ataxia and Spastic Paraplegia. Cambridge University Press, Cambridge, UK, 1939. 142–281 + xix–xxxii pp. LCCN ????

Fisher:1990:SME

[FBY90] R. A. (Ronald Aylmer) Fisher, John Henry Bennett, and Frank Yates. Statistical Methods, Experimental Design, and Scientific Inference. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1990. ISBN 0-19-852229-0. xxxii + 362, 248, 180 pp. LCCN

QA276. URL http://catdir.loc.gov/catdir/enhancements/
fy0602/90006726-d.html;   http://catdir.loc.gov/catdir/
enhancements/fy0602/90006726-t.html;   http://www.gbv.
de/dms/bowker/toc/9780198522294.pdf;   http://www.zentralblatt-
math.org/zmath/en/search/?an=0705.62003.

Fisher:1960:PPD

[FC60] Ronald A. Fisher and E. A. Cornish. The percentile points of distributions having known cumulants. Technometrics, 2(2):209– 225, May 1960. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1266546.

See errata [Ano60].

Fisher:1934:CLS

[FD34] R. A. Fisher and C. Diver. Crossing-over in the land snail Cepæa nemoralis, L. Nature, 133(3370):834–835, June 2, 1934. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL

https://www.nature.com/articles/133834b0.

Fisher:1948:RAI

[FD48] Ronald Aylmer Fisher and Daniel Dugué. Un résultat assez inattendu d’arithmétique des lois de probabilité. (French) [A rather unexpected result of arithmetic of the laws of probability]. Comptes rendus de l’Académie des sciences, Paris, 227:1205–1206, 1948. ISSN 0001-4036 (print), 2419-6304 (electronic).

Fisher:1947:THM

[FdB47] R. A. Fisher and G. R. de Beer. Thomas Hunt Morgan, 1866–1945 (obituary). Obituary Notices of the Royal Society, 5(15):451–466, February 1947. ISSN 1479-571X (print), 2053-9118 (electronic).

URL https://www.jstor.org/stable/769094.

Fisher:1926:VS

[FF26] R. A. Fisher and E. B. Ford. Variability of species. Nature, 118 (2971):515–516, October 9, 1926. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL https://www.nature.

com/articles/118515a0.

Fisher:1928:VSL

[FF28] R. A. Fisher and E. B. Ford. The variability of species in the Lepidoptera with reference to abundance and sex. Transactions of the Entomological Society of London, 76(2):367–379, January 1928. CODEN TRYEAK. ISSN 2053-2520 (print), 2056-810X

(electronic). URL https://onlinelibrary.wiley.com/doi/10.
1111/j.1365-2311.1929.tb01411.x.

Fisher:1943:SDB

[FF43] R. A. Fisher and J. A. Fraser Roberts. A sex difference in bloodgroup frequencies. Nature, 151(3840):640–641, June 5, 1943. CO-DEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

URL https://www.nature.com/articles/151640a0.

Fisher:1947:SGN

[FF47] R. A. Fisher and E. B. Ford. The spread of a gene in natural conditions in a colony of the moth Panaxia dominula L. Heredity, 1(2):143–174, October 1, 1947. CODEN NATUAS. ISSN 0018-067X (print), 1365-2540 (electronic). URL https://www.nature.

com/articles/hdy194711.

Fisher:1950:SWE

[FF50] R. A. Fisher and E. B. Ford. The “Sewall Wright Effect”. Heredity, 4(1):117–119, April 1, 1950. CODEN HDTYAT. ISSN 0018-067X (print), 1365-2540 (electronic). URL https://www.nature.com/

articles/hdy19508.

Fisher:1955:DRR

[FF55] R. A. Fisher and V. C. Fyfe. Double reduction at the rosy, or pink, locus in Lythrum salicaria. Nature, 176(4494):1176, December 17, 1955. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (elec-

tronic). URL https://www.nature.com/articles/1761176a0.

Fisher:1939:TTA

[FFH39] R. A. Fisher, E. B. Ford, and Julian Huxley. Taste-testing the anthropoid apes. Nature, 144(3652):750, October 28, 1939. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL

https://www.nature.com/articles/144750a0.

Fisher:1937:IMB

[FG37] R. A. Fisher and H. Gray. Inheritance in man; Boas’ data studied by the method of analysis variance. Annals of Eugenics, 8 (1):74–93, October 1, 1937. ISSN 2050-1420 (print), 2050-1439

(electronic). URL https://onlinelibrary.wiley.com/doi/10.
1111/j.1469-1809.1937.tb02162.x.

Fisher:1928:MMC

[FH28] R. A. Fisher and T. N. Hoblyn. Maximum- and minimumcorrelation tables in comparative climatology. Geografiska Annaler [Stockholm], 10(??):267–281, ???? 1928. ISSN 1651-3215 (print),

2001-4422 (electronic). URL https://www.jstor.org/stable/
519800.

Fisher:1956:NTB

[FH56] R. A. Fisher and M. J. R. Healy. New tables of Behrens’ test of significance. Journal of the Royal Statistical Society. Series B (Methodological), 18(2):212–216, ???? 1956. CODEN JST-

BAJ. ISSN 0035-9246. URL http://links.jstor.org/sici?
sici=0035-9246(1956)18:2<212:NTOBTO>2.0.CO;2-W&origin=
MSN;   http://www.jstor.org/stable/2983705.

Fienberg:1980:RFA

[FH80] Stephen E. Fienberg and D. V. Hinkley. R. A. Fisher: An Appreciation, volume 1 of Lecture notes in statistics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1980. ISBN 0-387-90476-X. ix + 208 pp.

LCCN QA276.16 .R18. URL http://www.loc.gov/catdir/
enhancements/fy0814/80000255-d.html;   http://www.loc.
gov/catdir/enhancements/fy0814/80000255-t.html.

Fienberg:1997:IRF

[Fie97] Stephen E. Fienberg. Introduction to R. A. Fisher on inverse probability and likelihood. Statistical Science, 12(3):161, August 1997. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1030037905.

Finney:1964:F

[Fin64a] D. J. Finney. Foreword. Biometrics, 20(2):237, June 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528391.

Finney:1964:SRF

[Fin64b] D. J. Finney. Sir Ronald Fisher’s contributions to biometric statistics. Biometrics, 20(2):322–329, June 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://

www.jstor.org/stable/2528400.

Fisher:1911:MB

[Fis11] R. A. Fisher. Mendelism and biometry, 1911. URL https://

digital.library.adelaide.edu.au/dspace/bitstream/2440/

3853/2/chap2.pdf. Presented at Second Undergraduate meeting of the Cambridge University Eugenics Society, Friday, 10 November 1911. Printed in [Ben83, pages 51–58].

Fisher:1912:ES

[Fis12a] R. A. Fisher. Evolution and society, 1912. URL https://

digital.library.adelaide.edu.au/dspace/bitstream/2440/

3853/2/chap2.pdf. Presented at Sixth Undergraduate meeting of the Cambridge University Eugenics Society, Wednesday, 13 March 1912.

Fisher:1912:ACF

[Fis12b] R. A. Fisher. On an absolute criterion for fitting frequency curves. Messenger of Mathematics, 41(??):155–160, ???? 1912.

Fisher:1913:AVA

[Fis13] R. A. Fisher. Applications of vector analysis to geometry. Messenger of Mathematics, 42(??):161–178, ???? 1913.

Fisher:1914:SHE

[Fis14] R. A. Fisher. Some hopes of a eugenist. Eugenics Review, 5(4): 309–315, January 1914. CODEN EUREAB. ISSN 0374-7573. URL

https://www.ncbi.nlm.nih.gov/pmc/articles/PMC2986993.

Fisher:1915:CPC

[Fis15a] R. A. Fisher. Cuénot on preadaptation: a criticism. Eugenics Review, 7(1):46–61, April 1915. CODEN EUREAB. ISSN 0374-

7573. URL https://www.ncbi.nlm.nih.gov/pmc/articles/
PMC2987108/.

Fisher:1915:ESP

[Fis15b] R. A. Fisher. The evolution of sexual preference. Eugenics Review, 7(3):184–192, October 1915. CODEN EUREAB.

ISSN 0374-7573. URL https://www.ncbi.nlm.nih.gov/pmc/
articles/PMC2987134.

Fisher:1915:FDV

[Fis15c] R. A. Fisher. Frequency distribution of the values of the correlation coefficient in samples from an indefinitely large population. Biometrika, 10(4):507–521, May 1915. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2331838;   https://zenodo.org/record/
1431583#.Xvzu-pZMEc4.

Fisher:1916:B

[Fis16] R. A. Fisher. Biometrika. Eugenics Review, 8(1):62–64, April 1916. CODEN EUREAB. ISSN 0374-7573. URL https://www.ncbi.

nlm.nih.gov/pmc/articles/PMC2990418/.

Fisher:1917:PE

[Fis17] R. A. Fisher. Positive eugenics. Eugenics Review, 9(3):206–212, October 1917. CODEN EUREAB. ISSN 0374-7573. URL https:

//www.ncbi.nlm.nih.gov/pmc/articles/PMC2942227/.

Fisher:1918:CBR

[Fis19a] R. A. Fisher. The correlation between relatives on the supposition of Mendelian inheritance. Transactions of the Royal Society of Edinburgh, 52(2):399–433, October 1, 1918–1919. CODEN TRSEAO. ISSN 0080-4568 (print), 2053-5945 (electronic). URL https:/

/www.biodiversitylibrary.org/page/48504075;   https://
www.cambridge.org/core/journals/earth-and-environmental-
science-transactions-of-royal-society-of-edinburgh/article/
xvthe-correlation-between-relatives-on-the-supposition-
of-mendelian-inheritance/A60675052E0FB78C561F66C670BC75DE.

Fisher:1919:CHV

[Fis19b] R. A. Fisher. The causes of human variability. Eugenics Review, 10(4):213–220, January 1919. CODEN EUREAB. ISSN 0374-

7573. URL https://www.ncbi.nlm.nih.gov/pmc/articles/
PMC2942138/.

Fisher:1919:GT

[Fis19c] R. A. Fisher. The genesis of twins. Genetics, 4(5):489–499, September 1919. CODEN GENTAE. ISSN 0016-6731 (print), 1943-2631

(electronic). URL https://www.genetics.org/content/4/5/
489;   https://www.ncbi.nlm.nih.gov/pmc/articles/PMC1200469.

Fisher:1920:GT

[Fis20a] R. A. Fisher. The genesis of twins. The Lancet (London, England), 195(5039):736–737, 1920. CODEN LANCAO. ISSN 0140-6736 (print), 1474-547x (electronic).

Fisher:1920:MEM

[Fis20b] R. A. Fisher. A mathematical examination of the methods of determining the accuracy of an observation by the mean error, and by the mean square error. Monthly Notices of the Royal Astronomical Society, 80(8):758–770, June 1920. CODEN MNRAA4. ISSN 0035-8711 (print), 1365-2966 (electronic). URL https://

ui.adsabs.harvard.edu/abs/1920MNRAS..80..758F/abstract.

Fisher:1920:MCS

[Fis20c] R. A. Fisher. Mathematics for collegiate students of agriculture and general science. Nature, 105(2631):131, April 1, 1920. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL

https://www.nature.com/articles/105131a0.

Fisher:1921:PEC

[Fis21a] R. A. Fisher. On the “probable error” of a coefficient of correlation deduced from a small sample. Metron, 1(4):1–32, ???? 1921. CO-DEN MRONAM. ISSN 0026-1424 (print), 2281-695X (electronic).

Fisher:1921:SRM

[Fis21b] R. A. Fisher. Some remarks on the methods formulated in a recent article on the quantitative analysis of plant growth. Annals of Applied Biology, 7(4):367–372, February 1921. CO-DEN AABIAV. ISSN 0003-4746 (print), 1744-7348 (electronic).

URL https://onlinelibrary.wiley.com/doi/abs/10.1111/j.
1744-7348.1921.tb05524.x.

Fisher:1921:SCV

[Fis21c] R. A. Fisher. Studies in crop variation. I. An examination of the yield of dressed grain from Broadbalk. Journal of Agricultural Science, 11(2):107–135, ???? 1921. CO-DEN JASIAB. ISSN 0021-8596 (print), 1469-5146 (electronic).

URL https://www.cambridge.org/core/journals/journal-
of-agricultural-science/article/studies-in-crop-variation-
i-an-examination-of-the-yield-of-dressed-grain-from-broadbalk/
882CB236D1EC608B1A6C74CA96F82CC3.

Fisher:1922:DEM

[Fis22a] R. A. Fisher. Darwinian evolution of mutations. Eugenics Review, 14(1):31–34, April 1922. CODEN EUREAB. ISSN 0374-

7573. URL https://www.ncbi.nlm.nih.gov/pmc/articles/
PMC2942494/.

Fisher:1922:ECC

[Fis22b] R. A. Fisher. The evolution of the conscience in civilised communities, in special relation to sexual vices. Eugenics Review, 14(3): 190–193, October 1922. CODEN EUREAB. ISSN 0374-7573. URL

https://www.ncbi.nlm.nih.gov/pmc/articles/PMC2942471/.

Fisher:1922:GFR

[Fis22c] R. A. Fisher. The goodness of fit of regression formulae, and the distribution of regression coefficients. Journal of the Royal Statistical Society, 85(4):597–612, June 1922. ISSN 0952-8385. URL

http://www.jstor.org/stable/2341124.

Fisher:1922:NDG

[Fis22d] R. A. Fisher. New data on the genesis of twins. Eugenics Review, 14 (2):115–117, July 1922. CODEN EUREAB. ISSN 0374-7573. URL

https://www.ncbi.nlm.nih.gov/pmc/articles/PMC2942516/.

Fisher:1922:DR

[Fis22e] R. A. Fisher. On the dominance ratio. Proceedings of the Royal Society of Edinburgh, 42(??):321–341, ???? 1922. CODEN PRSEAE. ISSN 0080-4541 (print), 2053-5902 (electronic).

Fisher:1922:ICT

[Fis22f] R. A. Fisher. On the interpretation of χ^{2} from contingency tables, and the calculation of P. Journal of the Royal Statistical Society, 85(1):87–94, January 1922. CODEN ???? ISSN 0952-

8385. URL http://www.jstor.org/stable/2340521;   https:
//onlinelibrary.wiley.com/doi/abs/10.1111/j.2397-2335.
1922.tb00768.x.

Fisher:1922:RTP

[Fis22g] R. A. Fisher. Review of A Treatise on Probability by J. M. Keynes. Eugenics Review, 14(1):46–50, April 1922. CODEN EU-

REAB. ISSN 0374-7573. URL https://www.ncbi.nlm.nih.gov/
pmc/articles/PMC2942496/.

Fisher:1922:SCI

[Fis22h] R. A. Fisher. Statistical considerations involved in Tables I and II of the above paper. Annals of Applied Biology, 9(2):142–145, June 1922. CODEN AABIAV. ISSN 0003-4746 (print), 1744-7348 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/abs/10.
1111/j.1744-7348.1922.tb05944.x.

Fisher:1922:SLG

[Fis22i] R. A. Fisher. The systematic location of genes by means of cross-over ratios. American Naturalist, 56(646):406–411, September/October 1922. CODEN AMNTA4. ISSN 0003-0147 (print),

1537-5323 (electronic). URL https://www.jstor.org/stable/
2456619.

Fisher:1922:MFT

[Fis22j] Ronald Aylmer Fisher. On the mathematical foundations of theoretical statistics. Philosophical Transactions of the Royal Society of London. Series A, Containing Papers of a Mathematical or Physical Character, 222(594–604):309–368, January 1, 1922. ISSN 0264-3952 (print), 2053-9258 (electronic). URL https://

royalsocietypublishing.org/doi/10.1098/rsta.1922.0009;
 https://www.jstor.org/stable/91208.

Fisher:1923:NDB

[Fis23a] R. A. Fisher. Note on Dr. Burnside’s recent paper on errors of observation. Proceedings of the Cambridge Philosophical Society. Mathematical and physical sciences, 21(??):655–658, ???? 1923. CODEN PCPSA4. ISSN 0008-1981.

Fisher:1923:PRD

[Fis23b] R. A. Fisher. Paradoxical rainfall data. Nature, 111(2788):465, April 7, 1923. CODEN NATUAS. ISSN 0028-0836 (print), 1476-

4687 (electronic). URL https://www.nature.com/articles/
111465a0.

Fisher:1923:SND

[Fis23c] R. A. Fisher. Statistical note on the data relating to nitrogen content and valuation. Journal of the Institute of Brewing, 29 (8):642–654, August 1923. CODEN JINBAL. ISSN 0046-9750 (print), 2050-0416 (electronic). URL https://onlinelibrary.

wiley.com/doi/abs/10.1002/j.2050-0416.1923.tb06645.x.

Fisher:1923:STA

[Fis23d] R. A. Fisher. Statistical tests of agreement between observation and hypothesis. Economica, 3(8):139–147, June 1923. ISSN 0013-0427 (print), 1468-0335 (electronic). URL https://www.jstor.

org/stable/2548482.

Fisher:1924:BSH

[Fis24a] R. A. Fisher. The biometrical study of heredity. Eugenics Review, 16(3):189–210, October 1924. CODEN EUREAB.

ISSN 0374-7573. URL https://www.ncbi.nlm.nih.gov/pmc/
articles/PMC2942624.

Fisher:1924:CUW

[Fis24b] R. A. Fisher. The conditions under which χ^{2} measures the discrepancey between observation and hypothesis. Journal of the Royal Statistical Society, 87(3):442–450, May 1924. ISSN 0952-

8385. URL http://www.jstor.org/stable/2341149;   https:
//onlinelibrary.wiley.com/doi/abs/10.1111/j.2397-2335.
1924.tb01691.x.

Fisher:1924:DPC

[Fis24c] R. A. Fisher. The distribution of the partial correlation coefficient. Metron, 3(3–4):329–332, ???? 1924. CODEN MRONAM. ISSN 0026-1424 (print), 2281-695x (electronic).

Fisher:1924:EMD

[Fis24d] R. A. Fisher. The elimination of mental defect. Eugenics Review, 16(2):114–116, July 1924. CODEN EUREAB. ISSN 0374-

7573. URL https://www.ncbi.nlm.nih.gov/pmc/articles/
PMC2942659/.

Fisher:1924:MSC

[Fis24e] R. A. Fisher. A method of scoring coincidences in tests with playing cards. Proceedings of the Society for Psychical Research, 34(Part 91):181–185, July 1924. CODEN PPSRA5. ISSN

0081-1475. URL http://iapsop.com/archive/materials/spr_
proceedings/spr_proceedings_v34_1924.pdf.

Fisher:1924:DYE

[Fis24f] R. A. Fisher. On a distribution yielding the error functions of several well known statistics. Proceedings of the International Congress of Mathematics, 2:805–813, 1924.

Fisher:1924:SCV

[Fis24g] R. A. Fisher. Studies in crop variation. III. The influence of rainfall on the yield of wheat at Rothamsted. Philosophical Transactions of the Royal Society of London Series B, Biological sciences, 213(402–410):89–142, January 1, 1924. CODEN PTRBAE. ISSN 0962-8436 (print), 1471-2970 (electronic). URL https://

royalsocietypublishing.org/doi/10.1098/rstb.1925.0003.

Fisher:1925:ASD

[Fis25a] R. A. Fisher. Applications of “Student’s” distribution. Metron, 5(3):90–104, ???? 1925. CODEN MRONAM. ISSN 0026-1424 (print), 2281-695x (electronic).

Fisher:1925:ESI

[Fis25b] R. A. Fisher. Expansion of ‘Student’s” integral in powers of n^{−1}. Metron, 5(3):109–120, ???? 1925. CODEN MRONAM. ISSN 0026-1424 (print), 2281-695x (electronic).

Fisher:1925:RBT

[Fis25c] R. A. Fisher. The resemblance between twins: a statistical examination of Lauterbach’s measurements. Genetics, 10(6):569–579, November 1925. CODEN GENTAE. ISSN 0016-6731 (print), 1943-

2631 (electronic). URL https://www.genetics.org/content/
10/6/569.

Fisher:1925:SMR

[Fis25d] R. A. Fisher. Statistical Methods for Research Workers. Oliver and Boyd, Edinburgh, UK; London, UK, 1925. ix + 239 pp.

Fisher:1925:SEI

[Fis25e] R. A. Fisher. Sur la solution de l’équation intégrale de M. V. Romanovsky. (French) [On the solution of M. V. Romanovsky’s integral equation]. Comptes rendus de l’Académie des sciences, Paris, 181(??):88–89, ???? 1925.

Fisher:1925:TSE

[Fis25f] R. A. Fisher. Theory of statistical estimation. Proceedings of the Cambridge Philosophical Society. Mathematical and physical sciences, 22(5):700–725, July 1925. CODEN PCPSA4. ISSN

0008-1981. URL https://www.cambridge.org/core/journals/
mathematical-proceedings-of-the-cambridge-philosophical-
society/article/theory-of-statistical-estimation/7A05FB68C83B36C0E91D42C76AB177D4.

Fisher:1926:AFE

[Fis26a] R. A. Fisher. The arrangement of field experiments. Journal of the Ministry of Agriculture Great Britain, 33(??):503–513, ???? 1926. CODEN JMINAQ. ISSN 0368-3087.

Fisher:1926:BTF

[Fis26b] R. A. Fisher. Bayes’ theorem and the fourfold table. Eugenics Review, 18(1):32–33, April 1926. CODEN EUREAB.

ISSN 0374-7573. URL https://www.ncbi.nlm.nih.gov/pmc/
articles/PMC2984620.

Fisher:1926:CN

[Fis26c] R. A. Fisher. Correspondence and notes. Quarterly journal of the Royal Meteorological Society, 52(219):250, July 1926. CODEN QJRMAM. ISSN 0035-9009 (print), 1477-870X (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/abs/10.
1002/qj.49705221904.

Fisher:1926:ECI

[Fis26d] R. A. Fisher. Eugenics — can it solve the problem of decay of civilizations? Eugenics Review, 18(2):128–136, July 1926. CODEN

EUREAB. ISSN 0374-7573. URL https://www.ncbi.nlm.nih.
gov/pmc/articles/PMC2984648/.

Fisher:1926:CFI

[Fis26e] R. A. Fisher. On the capillary forces in an ideal soil: correction of formulae given by W. B. Haines. Journal of Agricultural Science, 16(3):492–505, July 1926. CODEN JASIAB. ISSN 0021-8596 (print), 1469-5146 (electronic).

URL https://www.cambridge.org/core/journals/journal-
of-agricultural-science/article/on-the-capillary-forces-
in-an-ideal-soil-correction-of-formulae-given-by-w-b-haines/
25F9721C5B86BDC0360ED1E49AA3D417.

Fisher:1926:RS

[Fis26f] R. A. Fisher. On the random sequence. Quarterly journal of the Royal Meteorological Society, 52(219):250–??, ???? 1926. CODEN QJRMAM. ISSN 0035-9009 (print), 1477-870X (electronic).

Fisher:1926:PHS

[Fis26g] R. A. Fisher. Periodical health surveys. Journal of State Medicine, 34(??):446–449, ???? 1926. ISSN 0368-4407.

Fisher:1926:SIT

[Fis26h] R. A. Fisher. Statistical interpretation of Tables I and II of the above paper. Annals of Applied Biology, 13(2):253–255, May 1926. CODEN AABIAV. ISSN 0003-4746 (print), 1744-7348 (electronic).

URL https://onlinelibrary.wiley.com/doi/abs/10.1111/j.
1744-7348.1926.tb04269.x.

Fisher:1926:U

[Fis26i] R. A. Fisher. [unknown]. D.Sc. dissertation, Cambridge University, Cambridge, UK, 1926.

Fisher:1927:ATO

[Fis27a] R. A. Fisher. The actuarial treatment of official birth records. Eugenics Review, 19(2):103–108, July 1927. CODEN EUREAB.

ISSN 0374-7573. URL https://www.ncbi.nlm.nih.gov/pmc/
articles/PMC2987494/.

Fisher:1927:EFA

[Fis27b] R. A. Fisher. The effect of family allowances on population. In Report of the Public Conference on Family Allowances, pages 7–11. ????, ????, 1927.

Fisher:1927:SOM

[Fis27c] R. A. Fisher. On some objections to mimicry theory; statistical and genetic. Transactions of the Entomological Society of London, 75(2):269–278, December 1927. ISSN 2053-2520 (print),

2056-810X (electronic). URL https://onlinelibrary.wiley.
com/doi/abs/10.1111/j.1365-2311.1927.tb00074.x.

Fisher:1927:PPY

[Fis27d] R. A. Fisher. The principles and practice of yield trials. Nature, 120(3013):145–147, July 30, 1927. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL https://www.nature.

com/articles/120145a0.

Fisher:1927:SCV

[Fis27e] R. A. Fisher. Studies in crop variation. IV. The experimental determination of the value of top dressings with cereals. Journal of Agricultural Science, 17(4):548–562, October 1927. CODEN JASIAB. ISSN 0021-8596 (print),

1469-5146 (electronic). URL https://www.cambridge.org/
core/journals/journal-of-agricultural-science/article/

studies-in-crop-variation-iv-the-experimental-determination-
of-the-value-of-top-dressings-with-cereals/09320298DA8E0F0BAB62DD3245707DAB.

Fisher:1928:ADF

[Fis28a] R. A. Fisher. Appendix: Dr. Fisher’ statement concerning the effect of psychological card preferences. Proceedings of the Society for Psychical Research, 38(Part 109):269–271, ???? 1928. CODEN

PPSRA5. ISSN 0081-1475. URL http://iapsop.com/archive/
materials/spr_proceedings/spr_proceedings_v38_1928-29.
pdf.

Fisher:1928:CCM

[Fis28b] R. A. Fisher. Correlation coefficients in meteorology. Nature, 153 (3053):716, May 5, 1928. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL https://www.nature.com/

articles/121712a0.

Fisher:1928:DBR

[Fis28c] R. A. Fisher. The differential birth-rate: new light on causes from American figures. Eugenics Review, 20(3):183–184, October 1928. CODEN EUREAB. ISSN 0374-7573. URL https://www.ncbi.

nlm.nih.gov/pmc/articles/PMC2984765/.

Fisher:1928:EDR

[Fis28d] R. A. Fisher. The evolution of dominance; reply to Professor Sewall Wright. American Naturalist, 63(689):553–556, November/ December 1928. CODEN AMNTA4. ISSN 0003-0147 (print),

1537-5323 (electronic). URL https://www.journals.uchicago.
edu/doi/10.1086/280289.

Fisher:1928:FNC

[Fis28e] R. A. Fisher. Further note on the capillary forces in an ideal soil. Journal of Agricultural Science, 18(3):406–410, July 1928. CODEN JASIAB. ISSN 0021-8596 (print),

1469-5146 (electronic). URL https://www.cambridge.org/
core/journals/journal-of-agricultural-science/article/
further-note-on-the-capillary-forces-in-an-ideal-soil/
D34E23BD711969C5BC0FF899012876BE.

Fisher:1928:GSD

[Fis28f] R. A. Fisher. The general sampling distribution of the multiple correlation coefficient. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 121(788):654–673, December 1, 1928. CODEN PRLAAZ. ISSN 0080-4630 (print),

2053-9169 (electronic). URL https://royalsocietypublishing.
org/doi/10.1098/rspa.1928.0224.

Fisher:1928:ITR

[Fis28g] R. A. Fisher. Income-tax rebates; the birth-rate and our future policy. Eugenics Review, 20(2):79–81, July 1928. CODEN EU-

REAB. ISSN 0374-7573. URL https://www.ncbi.nlm.nih.gov/
pmc/articles/PMC2987494/.

Fisher:1928:PCM

[Fis28h] R. A. Fisher. On a property connecting the χ^{2} measure of discrepancy with the method of maximum likelihood. Atti del Congresso Internazionale dei Matematici, 6:95–100, 1928. ISSN 1013-8668.

Fisher:1928:PMR

[Fis28i] R. A. Fisher. The possible modification of the response of the wild type to recurrent mutations. American Naturalist, 62(679): 115–126, March/April 1928. CODEN AMNTA4. ISSN 0003-0147 (print), 1537-5323 (electronic). URL https://www.journals.

uchicago.edu/doi/10.1086/280234.

Fisher:1928:SMR

[Fis28j] R. A. Fisher. Statistical Methods for Research Workers. Oliver and Boyd, Edinburgh, UK; London, UK, second edition, 1928. xi + 269 pp.

Fisher:1928:TCG

[Fis28k] R. A. Fisher. Triplet children in Great Britain and Ireland. Proceedings of the Royal Society of London. Series B. Biological sciences, 102(717):286–311, January 2, 1928. CODEN PRLBA4. ISSN 0962-8452 (print), 1471-2954 (electronic). URL https://

royalsocietypublishing.org/doi/10.1098/rspb.1928.0005.

Fisher:1928:TFN

[Fis28l] R. A. Fisher. Two further notes on the origin of dominance. American Naturalist, 62(683):571–574, November/December 1928. CODEN AMNTA4. ISSN 0003-0147 (print), 1537-5323 (elec-

tronic). URL https://www.journals.uchicago.edu/doi/10.
1086/280234.

Fisher:1929:MPM

[Fis29a] R. A. Fisher. Moments and product moments of sampling distributions. Proceedings of the London Mathematical Society. Second Series, 30(3):199–238, 1929. CODEN PLMTAL. ISSN 0024-6115 (print), 1460-244x (electronic).

Fisher:1929:OF

[Fis29b] R. A. Fisher. The overproduction of food. Realist, 1(??):45–60,

???? 1929.

Fisher:1929:PD

[Fis29c] R. A. Fisher. Population and depopulation. Nature, 123(3097): 357–358, March 9, 1929. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL https://www.nature.com/

articles/123357a0.

Fisher:1929:PNE

[Fis29d] R. A. Fisher. A preliminary note on the effect of sodium silicate in increasing the yield of barley. Journal of Agricultural Science, 19(1):132–139, January 1929. CODEN JASIAB. ISSN 0021-8596 (print), 1469-5146 (electronic). URL https://www.cambridge.

org/core/journals/journal-of-agricultural-science/article/
preliminary-note-on-the-effect-of-sodium-silicate-in-increasing-
the-yield-of-barley/1C4024BBD9EE50ACA94A7BB65B8D16BA.

Fisher:1929:SE

[Fis29e] R. A. Fisher. The sieve of Eratosthenes. Mathematical Gazette, 14 (204):564–566, December 1929. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://www.jstor.

org/stable/3607073.

Fisher:1929:SMP

[Fis29f] R. A. Fisher. The statistical method in psychical research. Proceedings of the Society for Psychical Research, 39(Part 112):189–192, August 1929. CODEN PPSRA5. ISSN 0081-

1475.  URL http://iapsop.com/archive/materials/spr_
proceedings/spr_proceedings_v39_1930-31.pdf.

Fisher:1929:SBR

[Fis29g] R. A. Fisher. Statistics and biological research. Nature, 124(3120): 266–267, August 17, 1929. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL https://www.nature.com/

articles/124266a0.

Fisher:1929:TSH

[Fis29h] R. A. Fisher. Tests of significance in harmonic analysis. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 125(795):54–59, August 1, 1929. CODEN PRLAAZ. ISSN 0080-4630 (print), 2053-9169 (electronic). URL https://

royalsocietypublishing.org/doi/10.1098/rspa.1929.0151.

Fisher:1930:BE

[Fis30a] R. A. Fisher. Biometry and evolution. Nature, 126(3172):246– 247, August 16, 1930. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL https://www.nature.com/

articles/126246a0.

Fisher:1930:DGR

[Fis30b] R. A. Fisher. The distribution of gene ratios for rare mutations. Proceedings of the Royal Society of Edinburgh, 50(??):205–220, ???? 1930. CODEN PRSEAE. ISSN 0080-4541 (print), 2053-5902 (electronic).

Fisher:1930:EDC

[Fis30c] R. A. Fisher. The evolution of dominance in certain polymorphic species. American Naturalist, 64(694):385–406, September/ October 1930. CODEN AMNTA4. ISSN 0003-0147 (print), 1537-

5323 (electronic). URL https://www.journals.uchicago.edu/
doi/10.1086/280325.

Fisher:1930:GTN

[Fis30d] R. A. Fisher. The Genetical Theory of Natural Selection. Clarendon Press, Oxford, UK, 1930. 272 (est.) pp.

Fisher:1930:GMN

[Fis30e] R. A. Fisher. Genetics, mathematics, and natural selection. Nature, 126(3186):805–806, November 22, 1930. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL https://

www.nature.com/articles/126805a0.

Fisher:1930:IP

[Fis30f] R. A. Fisher. Inverse probability. Proceedings of the Cambridge Philosophical Society. Mathematical and physical sciences, 26(4):528–535, October 1930. CODEN PCPSA4. ISSN

0008-1981. URL https://www.cambridge.org/core/journals/
mathematical-proceedings-of-the-cambridge-philosophical-
society/article/inverse-probability/C9AB0A7C4566A3F9FCCEC489CA854814.

Fisher:1930:M

[Fis30g] R. A. Fisher. Mathematics. Nature, 125(3149):379, March 8, 1930. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (elec-

tronic). URL https://www.nature.com/articles/125379d0.

Fisher:1930:MDN

[Fis30h] R. A. Fisher. The moments of the distribution for normal samples of measures of departure from normality. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 130(812):16–28, December 2, 1930. CODEN PRLAAZ. ISSN 0080-4630 (print), 2053-9169 (electronic). URL https://

royalsocietypublishing.org/doi/10.1098/rspa.1930.0185.

Fisher:1930:MAP

[Fis30i] R. A. Fisher. Mortality amongst plants and its bearing on natural selection. Nature, 125(3165):972–973, June 28, 1930. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL

https://www.nature.com/articles/125972a0.

Fisher:1930:NTC

[Fis30j] R. A. Fisher. Note on a tri-colour (mosaic) mouse. Journal of Genetics, 23(??):77–81, September 1930. CODEN JOGNAU. ISSN 0022-1333 (print), 0973-7731 (electronic). URL https://link.

springer.com/article/10.1007/BF03052595.

Fisher:1930:SMR

[Fis30k] R. A. Fisher. Statistical Methods for Research Workers. Oliver and Boyd, Edinburgh, UK; London, UK, third edition, 1930. xi + 283 + 6 pp.

Fisher:1931:BEF

[Fis31a] R. A. Fisher. The biological effects of family allowances. Family Endowment Chronicle, 1(3):21–25, ???? 1931.

Fisher:1931:ED

[Fis31b] R. A. Fisher. The evolution of dominance. Biological Reviews of the Cambridge Philosophical Society, 6(4):345–368, October

1931. ISSN 1469-185X. URL https://onlinelibrary.wiley.
com/doi/10.1111/j.1469-185X.1931.tb01030.x.

Fisher:1931:PPE

[Fis31c] R. A. Fisher. Principles of plot experimentation in relation to the statistical interpretation of the results. In ????, editor, Report of a Conference on the Technique of Field Experiments Held at Rothamsted on 7 May 1931, page ?? ????, ????, 1931.

Fisher:1931:RPL

[Fis31d] R. A. Fisher. Recent publications: a letter from R. A. Fisher to the Editor. American Mathematical Monthly, 38(6):335–338, June/July 1931. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Fisher:1931:RWE

[Fis31e] R. A. Fisher. A review of Wright, S. (1931). Evolution in Mendelian populations. (Genetics, March 1931, 16(2): 97–159). Eugenics Review, 23(1):88–90, April 1931. CODEN EUREAB.

ISSN 0374-7573. URL https://www.ncbi.nlm.nih.gov/pmc/
articles/PMC2985000/pdf/eugenrev00302-0003.pdf.

Fisher:1931:SEE

[Fis31f] R. A. Fisher. The sampling error of estimated deviates, together with other illustrations of the properties and applications of the integrals and derivatives of the normal error function. British Association Mathematical Tables, 1(??):xxvi–xxxv,

???? 1931. URL https://digital.library.adelaide.edu.au/
dspace/bitstream/2440/15209/1/91.pdf.

Fisher:1932:BGT

[Fis32a] R. A. Fisher. The bearing of genetics on theories of evolution. Science Progress Twentieth Century, 27(106):273–287, October 1932.

URL https://www.jstor.org/stable/43429513.

Fisher:1932:EMG

[Fis32b] R. A. Fisher. The evolutionary modification of genetic phenomena. In ????, editor, Proceedings of the 6th International Congress on Genetics, volume 1, pages 165–172. ????, ????, 1932.

Fisher:1932:FAC

[Fis32c] R. A. Fisher. Family allowances in the contemporary economic situation. Eugenics Review, 24(2):87–95, July 1932. CODEN EU-

REAB. ISSN 0374-7573. URL https://www.ncbi.nlm.nih.gov/
pmc/articles/PMC2985161/.

Fisher:1932:IAC

[Fis32d] R. A. Fisher. Inheritance of acquired characters. Nature, 130 (3285):579, October 15, 1932. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL https://www.nature.com/

articles/130579a0.

Fisher:1932:IPU

[Fis32e] R. A. Fisher. Inverse probability and the use of likelihood. Proceedings of the Cambridge Philosophical Society. Mathematical and physical sciences, 28(3):257–261, July 1932. CODEN PCPSA4.

ISSN 0008-1981. URL https://www.cambridge.org/core/
journals/mathematical-proceedings-of-the-cambridge-philosophical-
society/article/inverse-probability-and-the-use-of-likelihood/
B3E94B37CA29899A79FE5C0CD0E6A6DD.

Fisher:1932:NSA

[Fis32f] R. A. Fisher. A new series of allelomorphs in mice. Nature, 129 (3247):130, January 23, 1932. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL https://www.nature.com/

articles/129130b0.

Fisher:1932:SSH

[Fis32g] R. A. Fisher. The Social Selection of Human Fertility. The Herbert Spencer Lecture, Deliver at Oxford, 8 June 1932. Clarendon Press, Oxford, UK, 1932. 30 pp.

Fisher:1932:SMR

[Fis32h] R. A. Fisher. Statistical Methods for Research Workers. Oliver and Boyd, Edinburgh, UK; London, UK, fourth edition, 1932. xiii + 307 pp.

Fisher:1933:CIP

[Fis33a] R. A. Fisher. The concepts of inverse probability and fiducial probability referring to unknown parameters. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, A-139(838):343–348, February 1, 1933. CODEN PRLAAZ. ISSN 0080-4630 (print), 2053-9169 (electronic). URL https://

royalsocietypublishing.org/doi/10.1098/rspa.1933.0021.

Fisher:1933:CRD

[Fis33b] R. A. Fisher. The contributions of Rothamsted to the development of the science of statistics. Annual Report Rothamsted Experimental Station, ??(??):43–50, ???? 1933.

Fisher:1933:MI

[Fis33c] R. A. Fisher. Mathematics of inheritance. Nature, 132(3348): 1012, December 30, 1933. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL https://www.nature.com/

articles/1321012a0.

Fisher:1933:NMF

[Fis33d] R. A. Fisher. Number of Mendelian factors in quantitative inheritance. Nature, 131(3307):400–401, March 18, 1933. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL

https://www.nature.com/articles/131400a0.

Fisher:1933:EAC

[Fis33e] R. A. Fisher. On the evidence against the chemical induction of melanism in Lepidoptera. Proceedings of the Royal Society B: Biological Sciences, 112(778):407–416, March 2, 1933. CODEN PRSBC7. ISSN 0080-4649 (print), 2053-9193 (elec-

tronic). URL https://royalsocietypublishing.org/doi/10.
1098/rspb.1933.0018.

Fisher:1933:PAA

[Fis33f] R. A. Fisher. Protective adaptations of animals, especially insects. Proceedings of the Entomological Society of London, 7(3):87–89, April 1933. ISSN 1472-0949 (print), 2056-8096

(electronic). URL https://onlinelibrary.wiley.com/doi/10.
1111/j.1365-3032.1933.tb00551.x.

Fisher:1933:SPE

[Fis33g] R. A. Fisher. Selection in the production of the ever-sporting stocks. Annals of Botany (Oxford, United Kingdom), os-47(4):727– 733, October 1933. CODEN ANBOA4. ISSN 0305-7364 (print),

1095-8290 (electronic). URL https://academic.oup.com/aob/
article-abstract/os-47/4/727/160862.

Fisher:1933:TSB

[Fis33h] R. A. Fisher. Tables for statisticians and biometricians. Nature, 131(3321):893–894, June 24, 1933. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL https://www.nature.

com/articles/131893a0.

Fisher:1934:AM

[Fis34a] R. A. Fisher. Adaptation and mutations. School Science Review, 59(??):294–301, ???? 1934. CODEN SSCRAD. ISSN 0036-6811.

Fisher:1934:AIS

[Fis34b] R. A. Fisher. The amount of information supplied by records of families as a function of the linkage in the population sampled. Annals of Eugenics, 6(1):66–70, October 1934. ISSN 2050-1420 (print), 2050-1439 (electronic). URL https://onlinelibrary.

wiley.com/doi/10.1111/j.1469-1809.1934.tb02107.x.

Fisher:1934:ANB

[Fis34c] R. A. Fisher. Appendix to The numbers of bacterial cells in field soils as estimated by the ratio method, by H. G. Thornton and P. H. H. Grey. Proceedings of the Royal Society of London. Series B. Biological sciences, 115(795):540–543, August 1, 1934. CODEN PRLBA4. ISSN 0962-8452 (print), 1471-2954 (elec-

tronic). URL https://royalsocietypublishing.org/doi/10.
1098/rspb.1934.0057.

Fisher:1934:CMD

[Fis34d] R. A. Fisher. The children of mental defectives. Departmental Committee on Sterilisation, Report, 1934.

Fisher:1934:CHF

[Fis34e] R. A. Fisher. Crest and hernia in fowls due to a single gene without dominance. Science, 80(2074):288–289, September 28, 1934. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (elec-

tronic). URL https://science.sciencemag.org/content/80/
2074/288.

Fisher:1934:EMA

[Fis34f] R. A. Fisher. The effect of methods of ascertainment upon the estimation of frequencies. Annals of Eugenics, 6(1):13–25, October 1934. CODEN ???? ISSN 2050-1420 (print), 2050-1439 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/abs/10.
1111/j.1469-1809.1934.tb02105.x.

Fisher:1934:INS

[Fis34g] R. A. Fisher. Indeterminism and natural selection. Philosophy of Science, 1(1):99–117, January 1934. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http:

//www.jstor.org/stable/184485;   https://www.journals.
uchicago.edu/doi/10.1086/286308.

Fisher:1934:PLQ

[Fis34h] R. A. Fisher. Probability likelihood and quantity of information in the logic of uncertain inference. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 146(856): 1–8, August 1, 1934. CODEN PRLAAZ. ISSN 0080-4630 (print),

2053-9169 (electronic). URL https://royalsocietypublishing.
org/doi/10.1098/rspa.1934.0134.

Fisher:1934:PWT

[Fis34i] R. A. Fisher. Professor Wright on the theory of dominance. American Naturalist, 68(717):370–374, July/August 1934. CODEN AM-NTA4. ISSN 0003-0147 (print), 1537-5323 (electronic). URL

https://www.jstor.org/stable/2456936.

Fisher:1934:ROE

[Fis34j] R. A. Fisher. Randomisation, and an old enigma of card play. Mathematical Gazette, 18(231):294–297, December 1934. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Fisher:1934:SRE

[Fis34k] R. A. Fisher. Some results of an experiment on dominance in poultry, with special reference to polydactyly. Proceedings of the Linnean Society of London, 147(??):71–81, ???? 1934. CODEN PLSLAR. ISSN 0370-0461 (print), 1747-2741 (elec-

tronic). URL https://hekyll.services.adelaide.edu.au/
dspace/bitstream/2440/15116/1/116.pdf.

Fisher:1934:SMR

[Fis34l] R. A. Fisher. Statistical Methods for Research Workers. Oliver and Boyd, Edinburgh, UK; London, UK, fifth edition, 1934. xiii + 319 pp.

Fisher:1934:USE

[Fis34m] R. A. Fisher. The use of simultaneous estimation in the evaluation of linkage. Annals of Eugenics, 6(1):71–76, October 1934. CODEN ???? ISSN 2050-1420 (print), 2050-1439 (electronic).

URL https://onlinelibrary.wiley.com/doi/abs/10.1111/j.
1469-1809.1934.tb02108.x.

Fisher:1934:VBA

[Fis34n] R. A. Fisher. The ’Viceroy’ (Basilarchia archippus Cram.) mistaken for its model, the Monarch (Danausplexippus Linn.). Proceedings of the Royal Entomological Society of London, 9(3):97– ??, March 1934. CODEN JENGA8. ISSN 0047-2409 (print),

2056-5313 (electronic). URL https://onlinelibrary.wiley.
com/doi/epdf/10.1111/j.1365-3032.1935.tb00014.x.

Fisher:1934:TNP

[Fis34o] Ronald Aylmer Fisher. Two new properties of mathematical likelihood. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 144(852):285–307, March 29, 1934. CODEN PRLAAZ. ISSN 0080-4630 (print), 2053-9169 (elec-

tronic). URL https://royalsocietypublishing.org/doi/10.
1098/rspa.1934.0050.

Fisher:1935:BRB

[Fis35a] R. A. Fisher. Book review: Biomathematics. By W. M. Feldman, M.D., B.S., M.R.C.P. (Lond.), F.R.A.S., F.R.S. (Ed.). Second edition. Charles Griffin and Co. Ltd., 1935. Pp. viii + 480. Annals of Eugenics, 6(2):252, June 1935. CODEN ???? ISSN 2050-1420 (print), 2050-1439 (electronic). URL https://onlinelibrary.

wiley.com/doi/abs/10.1111/j.1469-1809.1935.tb02231.x.

Fisher:1935:CZS

[Fis35b] R. A. Fisher. The case of zero survivors in probit assays. Annals of Applied Biology, 22(1):164–165, February 1935. CODEN AABIAV. ISSN 0003-4746 (print), 1744-7348 (electronic).

Fisher:1935:CDF

[Fis35c] R. A. Fisher. Contribution to a discussion of F. Yates’ paper on complex experiments. Journal of the Royal Statistical Society, 98 (1):229–231, ???? 1935. ISSN 0952-8385.

Fisher:1935:CDJ

[Fis35d] R. A. Fisher. Contribution to a discussion of J. Neyman’s paper on statistical problems in agricultural experimentation. Journal of the Royal Statistical Society, 98(1):154–157, ???? 1935. ISSN

0952-8385. URL https://www.jstor.org/stable/2342435.

Fisher:1935:DE

[Fis35e] R. A. Fisher. The Design of Experiments. Oliver and Boyd, Edinburgh, UK; London, UK, 1935. ???? pp.

Fisher:1935:DLD

[Fis35f] R. A. Fisher. The detection of linkage with ‘dominant’ abnormalities. Annals of Eugenics, 6(2):187–201, June 1, 1935. ISSN 2050-1420 (print), 2050-1439 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/10.1111/j.1469-1809.1935.
tb02227.x.

Fisher:1935:DLR

[Fis35g] R. A. Fisher. The detection of linkage with recessive abnormalities. Annals of Eugenics, 6(4):339–351, December 1935. CO-DEN ???? ISSN 2050-1420 (print), 2050-1439 (electronic).

URL https://onlinelibrary.wiley.com/doi/abs/10.1111/j.
1469-1809.1935.tb02116.x.

Fisher:1935:DP

[Fis35h] R. A. Fisher. Dominance in poultry. Philosophical Transactions of the Royal Society of London Series B, Biological sciences, 225(523):197–226, July 30, 1935. CODEN PTRBAE. ISSN 0962-8436 (print), 1471-2970 (electronic). URL https://

royalsocietypublishing.org/doi/10.1098/rstb.1935.0011.

Fisher:1935:DPS

[Fis35i] R. A. Fisher. Dynamics of population: Social and biological significance of changing birth rates in the United States. Nature, 135 (3402):46–48, January 12, 1935. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL https://www.nature.

com/articles/135046a0.

Fisher:1935:EAP

[Fis35j] R. A. Fisher. Eugenics, academic and practical. Eugenics Review, 27(2):95–100, July 1935. CODEN EUREAB. ISSN 0374-

7573. URL https://www.ncbi.nlm.nih.gov/pmc/articles/
PMC2985438/.

Fisher:1935:FAS

[Fis35k] R. A. Fisher. The fiducial argument in statistical inference. Annals of Eugenics, 6(4):391–398, December 1935. CODEN ???? ISSN 2050-1420 (print), 2050-1439 (electronic). URL https:

//onlinelibrary.wiley.com/doi/abs/10.1111/j.1469-1809.
1935.tb02120.x.

Fisher:1935:IFD

[Fis35l] R. A. Fisher. The inheritance of fertility — Dr Wagner-Manslau’s tables. Annals of Eugenics, 6(2):225–251, June 1935.

ISSN 2050-1420 (print), 2050-1439 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/10.1111/j.1469-1809.1935.
tb02230.x.

Fisher:1935:LSP

[Fis35m] R. A. Fisher. Linkage studies and the prognosis of hereditary ailments. In ????, editor, Proceedings of the International Congress on Life Assurance Medicine, page ?? ????, ????, 1935.

Fisher:1935:LII

[Fis35n] R. A. Fisher. The logic of inductive inference. Journal of the Royal Statistical Society, 98(1):39–54 + 55–82 (discussion), ????

1935. ISSN 0952-8385. URL http://www.jstor.org/stable/
2342435;   https://onlinelibrary.wiley.com/doi/abs/10.
1111/j.2397-2335.1935.tb04208.x.

Fisher:1935:MDU

[Fis35o] R. A. Fisher. The mathematical distributions used in the common tests of significance. Econometrica, 3(4):353–365, October 1935. CODEN ECMTA7. ISSN 0012-9682. URL https://www.jstor.

org/stable/1905628.

Fisher:1935:SCE

[Fis35p] R. A. Fisher. On the selective consequences of East’s (1927) theory of heterostylism in Lythrum. Journal of Genetics, 30(??):369–382, May 1935. CODEN JOGNAU. ISSN 0022-1333 (print), 0973-

7731 (electronic). URL https://link.springer.com/article/
10.1007/BF02982245.

Fisher:1935:SL

[Fis35q] R. A. Fisher. The sheltering of lethals. American Naturalist, 69(724):446–455, September 1935. CODEN AMNTA4. ISSN 0003-0147 (print), 1537-5323 (electronic). URL https://www.

journals.uchicago.edu/doi/10.1086/280618.

Fisher:1935:ST

[Fis35r] R. A. Fisher. Statistical tests. Nature, 136(3438):474, September 21, 1935. CODEN NATUAS. ISSN 0028-0836 (print), 1476-

4687 (electronic). URL https://www.nature.com/articles/
136474b0.

Fisher:1936:CF

[Fis36a] R. A. Fisher. Curve fitting. Nature, 138(3500):934, November 28, 1936. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (elec-

tronic). URL https://www.nature.com/articles/138934a0.

Fisher:1936:HDS

[Fis36b] R. A. Fisher. The half-drill strip system agricultural experiments. Nature, 138(3504):1101, December 26, 1936. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL https://

www.nature.com/articles/1381101a0.

Fisher:1936:HLD

[Fis36c] R. A. Fisher. Heterogeneity of linkage data for Friedreich’s ataxia and the spontaneous antigens. Annals of Eugenics, 7(1):17–21, June 1936. CODEN ???? ISSN 2050-1420 (print), 2050-1439 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/abs/10.
1111/j.1469-1809.1936.tb02125.x.

Fisher:1936:ITB

[Fis36d] R. A. Fisher. Income-tax and birth-rates: family allowances. The Times [London, UK], ??(??):??, April 30, 1936. ISSN 0140-0460, 0956-1382.

Fisher:1936:MSI

[Fis36e] R. A. Fisher. The measurement of selective intensity. Proceedings of the Royal Society of London. Series B. Biological sciences, 121(820):58–62, August 1, 1936. CODEN PRLBA4. ISSN 0962-8452 (print), 1471-2954 (electronic). URL https://

royalsocietypublishing.org/doi/10.1098/rspb.1936.0052.

Fisher:1936:SMR

[Fis36f] R. A. Fisher. Statistical Methods for Research Workers. Oliver and Boyd, Edinburgh, UK; London, UK, sixth edition, 1936. xiii + 319 pp.

Fisher:1936:TSA

[Fis36g] R. A. Fisher. Tests of significance applied to Haldane’s data on partial sex linkage. Annals of Eugenics, 7(2):87– 104, September 1, 1936. ISSN 2050-1420 (print), 2050-1439

(electronic). URL https://onlinelibrary.wiley.com/doi/10.
1111/j.1469-1809.1936.tb02131.x.

Fisher:1936:CRL

[Fis36h] R. A. Fisher. “The coefficient of racial likeness” and the future of craniometry. The Journal of the Royal Anthropological Institute of Great Britain and Ireland, 66(1):57–63, January/June 1936. ISSN

0307-3114. URL https://www.jstor.org/stable/2844116.

Fisher:1936:UI

[Fis36i] R. A. Fisher. Uncertain inference. Proceedings of the American Academy of Arts and Sciences, 71(4):245–258, October 1936. CO-DEN PAAAAV. ISSN 0065-6836.

Fisher:1936:UMM

[Fis36j] R. A. Fisher. The use of multiple measurements in taxonomic problems. Annals of Eugenics, 7(2):179–188, September 1936. CODEN ???? ISSN 2050-1420 (print), 2050-1439 (electronic).

URL https://onlinelibrary.wiley.com/doi/abs/10.1111/j.
1469-1809.1936.tb02137.x.

Fisher:1936:MWB

[Fis36k] R. A. Fisher, M.A., Sc.D., F.R.S. Has Mendel’s work been rediscovered? Annals of Science, 1(2):115–137, April 15, 1936. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X

(electronic). URL https://pdfs.semanticscholar.org/13e9/
a7f537d955b139a3d3dc033e48e4a53dc8be.pdf;   https://www.
tandfonline.com/doi/abs/10.1080/00033793600200111.

Fisher:1937:CSW

[Fis37a] R. A. Fisher. The character of Sheppard’s work [W. F. Sheppard. (obituary)]. Annals of Eugenics, 8(1):11–12, October 1937. ISSN 2050-1420 (print), 2050-1439 (electronic). URL https:/

/onlinelibrary.wiley.com/doi/10.1111/j.1469-1809.1937.
tb02156.x.

Fisher:1937:PRM

[Fis37b] R. A. Fisher. On a point raised by M. S. Bartlett on fiducial probability. Annals of Eugenics, 7(4):370–375, June 1937. CODEN ???? ISSN 2050-1420 (print), 2050-1439 (electronic).

URL https://onlinelibrary.wiley.com/doi/abs/10.1111/j.
1469-1809.1937.tb02154.x.

Fisher:1937:PKP

[Fis37c] R. A. Fisher. Professor Karl Pearson and the method of moments. Annals of Eugenics, 7(4):303–318, June 1937. ISSN 2050-1420 (print), 2050-1439 (electronic). URL https://onlinelibrary.

wiley.com/doi/abs/10.1111/j.1469-1809.1937.tb02149.x.

Fisher:1937:RBV

[Fis37d] R. A. Fisher. The relation between variability and abundance shown by the measurements of the eggs of British nesting birds. Proceedings of the Royal Society of London. Series B. Biological sciences, 122(826):1–26, March 3, 1937. CODEN PRLBA4. ISSN 0962-8452 (print), 1471-2954 (electronic). URL https://

royalsocietypublishing.org/doi/10.1098/rspb.1937.0006.

Fisher:1937:WAA

[Fis37e] R. A. Fisher. The wave of advance of advantageous genes. Annals of Eugenics, 7(4):355–369, June 1937. CODEN ???? ISSN 2050-1420 (print), 2050-1439 (electronic). URL https:

//onlinelibrary.wiley.com/doi/abs/10.1111/j.1469-1809.
1937.tb02153.x.

Fisher:1937:DE

[Fis37f] Ronald Aylmer Fisher. The Design of Experiments. Oliver and Boyd, Edinburgh, UK; London, UK, second edition, 1937. xi + 260 pp. LCCN HA29 .F48 1937.

Fisher:1938:DAF

[Fis38a] R. A. Fisher. The design and analysis of factorial experiments. Nature, 142(3585):90–92, July 16, 1938. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL https://www.

nature.com/articles/142090a0.

Fisher:1938:DPF

[Fis38b] R. A. Fisher. Dominance in poultry: feathered feet, rose comb, internal pigment and pile. Proceedings of the Royal Society of London. Series B. Biological sciences, 125(838):25–48, March 3, 1938. CODEN PRLBA4. ISSN 0962-8452 (print), 1471-2954 (elec-

tronic). URL https://royalsocietypublishing.org/doi/10.
1098/rspb.1938.0011.

Fisher:1938:ME

[Fis38c] R. A. Fisher. The mathematics of experimentation. Nature, 142 (3592):442–443, September 3, 1938. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL https://www.

nature.com/articles/142442a0.

Fisher:1938:PA

[Fis38d] R. A. Fisher. Presidential address. Sankhyā (Indian Journal of Statistics), 4(1):14–17, ???? 1938. CODEN SNKYA5. ISSN 0036-

4452. URL https://www.jstor.org/stable/40383882.

Fisher:1938:QRE

[Fis38e] R. A. Fisher. Quelques remarques sur l’estimation en statistique. (French) [Some remarks on estimation in statistics]. Biotypologie, 6(??):153–158, ???? 1938. CODEN BTYPA5. ISSN 0366-3892.

Fisher:1938:SMR

[Fis38f] R. A. Fisher. Statistical Methods for Research Workers. Oliver and Boyd, Edinburgh, UK; London, UK, seventh edition, 1938. xiv + 356 pp. LCCN HA29.F5 1938; QH 232; SK 850.

Fisher:1938:STE

[Fis38g] R. A. Fisher. Statistical theory of estimation. Calcutta University Readership Lectures., 1938.

Fisher:1938:SUM

[Fis38h] R. A. Fisher. The statistical utilization of multiple measurements. Annals of Eugenics, 8(4):376–386, August 1938. CO-DEN ???? ISSN 2050-1420 (print), 2050-1439 (electronic).

URL https://onlinelibrary.wiley.com/doi/abs/10.1111/j.
1469-1809.1938.tb02189.x.

Fisher:1938:SST

[Fis38i] R. A. Fisher. Supplement: On the statistical treatment op the relation between sea-level characteristics and high-altitude acclimization. Proceedings of the Royal Society of London. Series B. Biological sciences, 126(842):25–29, September 23, 1938. CODEN PRLBA4. ISSN 0962-8452 (print), 1471-2954 (elec-

tronic). URL https://royalsocietypublishing.org/doi/10.
1098/rspb.1938.0044.

Fisher:1939:CSP

[Fis39a] R. A. Fisher. The comparison of samples with possibly unequal variances. Annals of Eugenics, 9(2):174–180, June 1939. ISSN 2050-1420 (print), 2050-1439 (electronic).

Fisher:1939:GL

[Fis39b] R. A. Fisher. The Galton Laboratory. Science, 90(2341):436, November 10, 1939. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

Fisher:1939:NFI

[Fis39c] R. A. Fisher. A note on fiducial inference. Annals of Mathematical Statistics, 10(4):383–388, December 1939. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http:

//projecteuclid.org/euclid.aoms/1177732151.

Fisher:1939:PPF

[Fis39d] R. A. Fisher. The precision of the product formula for the estimation of linkage. Annals of Eugenics, 9(1):50–54, January 1939. CODEN ???? ISSN 2050-1420 (print), 2050-1439 (electronic).

URL https://onlinelibrary.wiley.com/doi/abs/10.1111/j.
1469-1809.1939.tb02196.x.

Fisher:1939:SDS

[Fis39e] R. A. Fisher. The sampling distribution of some statistics obtained from non-linear equations. Annals of Eugenics, 9(3):238–249, August 1939. CODEN ???? ISSN 2050-1420 (print), 2050-1439 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/abs/10.
1111/j.1469-1809.1939.tb02211.x.

Fisher:1939:SFW

[Fis39f] R. A. Fisher. Selective forces in wild populations of Paratettix texanus. Annals of Eugenics, 9(2):109–122, June 1939. ISSN 2050-1420 (print), 2050-1439 (electronic).

Fisher:1939:SDF

[Fis39g] R. A. Fisher. Stage of development as a factor influencing the variance in the number of offspring, frequency of mutants and related quantities. Annals of Eugenics, 9(4):406–408, December 1939. CODEN ???? ISSN 2050-1420 (print), 2050-1439 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/abs/10.
1111/j.1469-1809.1939.tb02221.x.

Fisher:1939:S

[Fis39h] R. A. Fisher. “Student”. Annals of Eugenics, 9(1):1–9, January 1939. CODEN ???? ISSN 2050-1420 (print), 2050-1439 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/abs/10.

1111/j.1469-1809.1939.tb02192.x. Obituary for William Sealy

Gossett (‘Student’).

Fisher:1940:EPR

[Fis40a] R. A. Fisher. The estimation of the proportion of recessives from tests carried out on a sample not wholly unrelated. Annals of Eugenics, 10(1):160–170, January 1940. CO-DEN ???? ISSN 2050-1420 (print), 2050-1439 (electronic).

URL https://onlinelibrary.wiley.com/doi/abs/10.1111/j.
1469-1809.1940.tb02245.x.

Fisher:1940:EDP

[Fis40b] R. A. Fisher. An examination of the different possible solutions of a problem in incomplete blocks. Annals of Eugenics, 10(1): 52–75, January 1940. CODEN ???? ISSN 2050-1420 (print),

2050-1439 (electronic). URL https://onlinelibrary.wiley.
com/doi/abs/10.1111/j.1469-1809.1940.tb02237.x.

Fisher:1940:GL

[Fis40c] R. A. Fisher. The Galton Laboratory. Science, 91(2350):44–45, January 12, 1940. CODEN SCIEAS. ISSN 0036-8075 (print),

1095-9203 (electronic). URL https://science.sciencemag.org/
content/91/2350/44.2.

Fisher:1940:SDF

[Fis40d] R. A. Fisher. On the similarity of the distributions found for the test of significance in harmonic analysis, and in Steven’s problem in geometrical probability. Annals of Eugenics, 10(1):14–17, January 1940. CODEN ???? ISSN 2050-1420 (print), 2050-1439 (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/abs/10.
1111/j.1469-1809.1940.tb02233.x.

Fisher:1940:PDF

[Fis40e] R. A. Fisher. The precision of discriminant functions. Annals of Eugenics, 10(1):422–429, January 1940. CODEN ???? ISSN 2050-1420 (print), 2050-1439 (electronic). URL https:

//onlinelibrary.wiley.com/doi/abs/10.1111/j.1469-1809.
1940.tb02264.x.

Fisher:1941:AAB

[Fis41a] R. A. Fisher. The asymptotic approach to Behrens’s integral, with further tables for the d test of significance. Annals of Eugenics, 11(1):141–172, January 1941. CODEN ???? ISSN 2050-1420 (print), 2050-1439 (electronic). URL https://onlinelibrary.

wiley.com/doi/abs/10.1111/j.1469-1809.1941.tb02281.x.

Fisher:1941:AEA

[Fis41b] R. A. Fisher. Average excess and average effect of a gene substitution. Annals of Eugenics, 11(1):53–63, January 1941. CO-DEN ???? ISSN 2050-1420 (print), 2050-1439 (electronic).

URL https://onlinelibrary.wiley.com/doi/abs/10.1111/j.
1469-1809.1941.tb02272.x.

Fisher:1941:COS

[Fis41c] R. A. Fisher. Completely orthogonal 9 × 9 squares. A correction. Annals of Eugenics, 11(1):402–403, January 1941. ISSN 2050-1420 (print), 2050-1439 (electronic). URL https://onlinelibrary.

wiley.com/doi/abs/10.1111/j.1469-1809.1941.tb02303.x.

Fisher:1941:IEF

[Fis41d] R. A. Fisher. The interpretation of experimental four-fold tables. Science, 94(2435):210–211, August 19, 1941. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL https://

science.sciencemag.org/content/94/2435/210.

Fisher:1941:LSP

[Fis41e] R. A. Fisher. The likelihood solution of a problem in compounded probabilities. Annals of Eugenics, 11(1):306–307, January 1941. CODEN ???? ISSN 2050-1420 (print), 2050-1439 (electronic).

URL https://onlinelibrary.wiley.com/doi/abs/10.1111/j.
1469-1809.1941.tb02293.x.

Fisher:1941:NBD

[Fis41f] R. A. Fisher. The negative binomial distribution. Annals of Eugenics, 11(1):182–187, January 1941. ISSN 2050-1420 (print),

2050-1439 (electronic). URL https://onlinelibrary.wiley.
com/doi/abs/10.1111/j.1469-1809.1941.tb02284.x.

Fisher:1941:NCS

[Fis41g] R. A. Fisher. New cyclic solutions to problems in incomplete blocks. Annals of Eugenics, 11(1):290–299, January 1941. CO-DEN ???? ISSN 2050-1420 (print), 2050-1439 (electronic).

URL https://onlinelibrary.wiley.com/doi/abs/10.1111/j.
1469-1809.1941.tb02291.x.

Fisher:1941:SCT

[Fis41h] R. A. Fisher. Some combinatorial theorems and enumerations connected with the numbers of diagonal types of a Latin square. Annals of Eugenics, 11(1):395–401, January 1941. CO-DEN ???? ISSN 2050-1420 (print), 2050-1439 (electronic).

URL https://onlinelibrary.wiley.com/doi/abs/10.1111/j.
1469-1809.1941.tb02302.x.

Fisher:1941:SMR

[Fis41i] R. A. Fisher. Statistical Methods for Research Workers. Oliver and Boyd, Edinburgh, UK; London, UK, eighth edition, 1941. xv + 344 pp. LCCN HA29.F5 1941; QH 232; SK 850.

Fisher:1941:TCP

[Fis41j] R. A. Fisher. The theoretical consequences of polyploid inheritance for the Mid style form of Lythrum salicaria. Annals of Eugenics, 11(1):31–38, January 1941. ISSN 2050-1420 (print), 2050-1439

(electronic). URL https://onlinelibrary.wiley.com/doi/10.
1111/j.1469-1809.1941.tb02268.x.

Fisher:1941:TCF

[Fis41k] R. A. Fisher. The theory of confounding in factorial experiments in relation to the theory of groups. Annals of Eugenics, 11(1):341–353, January 1941. CODEN ???? ISSN 2050-1420 (print), 2050-1439 (electronic). URL https://onlinelibrary.

wiley.com/doi/abs/10.1111/j.1469-1809.1941.tb02298.x.

Fisher:1942:FPM

[Fis42a] R. A. Fisher. The fundamental principles of mathematical statistics sampling methods in forestry and range management. Nature, 150(3798):196, August 15, 1942. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL https://www.nature.

com/articles/150196a0.

Fisher:1942:PC

[Fis42b] R. A. Fisher. The polygene concept. Nature, 150(3796):154, August 1, 1942. CODEN NATUAS. ISSN 0028-0836 (print), 1476-

4687 (electronic). URL https://www.nature.com/articles/
150154b0.

Fisher:1942:TCF

[Fis42c] R. A. Fisher. The theory of confounding in factorial experiments in relation to the theory of groups. Annals of Eugenics, 11(1):

341–353, January 1, 1942. ISSN 2050-1420 (print), 2050-1439

(electronic). URL https://onlinelibrary.wiley.com/doi/10.
1111/j.1469-1809.1941.tb02298.x.

Fisher:1942:UR

[Fis42d] R. A. Fisher. Unsigned reviews. The Lancet (London, England), 240(6213):377–378, 1942. CODEN LANCAO. ISSN 0140-6736 (print), 1474-547x (electronic).

Fisher:1942:DE

[Fis42e] Ronald Aylmer Fisher. The Design of Experiments. Oliver and Boyd, Edinburgh, UK; London, UK, third edition, 1942. xi + 236 pp. LCCN HA29 .F48 1942.

Fisher:1943:ATS

[Fis43a] R. A. Fisher. The advanced theory of statistics. Nature, 152(3859): 431–432, October 16, 1943. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL https://www.nature.com/

articles/152431a0.

Fisher:1943:ADR

[Fis43b] R. A. Fisher. Allowance for double reduction in the calculation of genotype frequencies with polysomic inheritance. Annals of Eugenics, 12(1):169–171, January 1943. ISSN 2050-1420 (print), 2050-1439 (electronic). URL https://onlinelibrary.

wiley.com/doi/10.1111/j.1469-1809.1943.tb02320.x.

Fisher:1943:BRF

[Fis43c] R. A. Fisher. The birth-rate and family allowances. Agenda, 2(??): 124–133, ???? 1943.

Fisher:1943:NDB

[Fis43d] R. A. Fisher. Note on Dr. Berkson’s criticism of tests of significance. Journal of the American Statistical Association, 38(221): 103–104, March 1943. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2279319.

Fisher:1943:SCF

[Fis43e] R. A. Fisher. A system of confounding for factors with more than two alternatives, giving completely orthogonal cubes and higher powers. Annals of Eugenics, 12(1):283–290, January 1, 1943. ISSN 2050-1420 (print), 2050-1439 (electronic).

URL https://onlinelibrary.wiley.com/doi/abs/10.1111/j.
1469-1809.1943.tb02332.x.

Fisher:1943:TDA

[Fis43f] R. A. Fisher. A theoretical distribution for the apparent abundance of different species. Journal of Animal Ecology, 12(??):54–58, ???? 1943. CODEN JAECAP. ISSN 0021-8790 (print), 1365-2656 (electronic).

Fisher:1944:SMR

[Fis44] R. A. Fisher. Statistical Methods for Research Workers, volume 5 of Biological Monographs and Manuals. Oliver and Boyd, Edinburgh, UK; London, UK, ninth edition, 1944. xv + 350 pp. LCCN HA29.F5 1944; QH 232; SK 850.

Fisher:1945:AHS

[Fis45a] R. A. Fisher. Anti-Hr serum of Levine. Nature, 155(3940):543, May 5, 1945. CODEN NATUAS. ISSN 0028-0836 (print), 1476-

4687 (electronic). URL https://www.nature.com/articles/
155543a0.

Fisher:1945:GTM

[Fis45b] R. A. Fisher. G. L. Taylor, M.D., Ph.D., F.R.C.P. (obituary). British Medical Journal, 1:463–??, ???? 1945. CODEN BMJOAE. ISSN 0007-1447.

Fisher:1945:ISI

[Fis45c] R. A. Fisher. The Indian Statistical Institute. Nature, 156(3972): 722, December 15, 1945. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL https://www.nature.com/

articles/156722a0.

Fisher:1945:LIN

[Fis45d] R. A. Fisher. The logical inversion of the notion of the random variable. Sankhyā (Indian Journal of Statistics), 7(2):129–132, November 1945. CODEN SNKYA5. ISSN 0036-4452. URL https:

//www.jstor.org/stable/25047836.

Fisher:1945:NTT

[Fis45e] R. A. Fisher. A new test for 2 × 2 tables. Nature, 156(3961): 388, September 29, 1945. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL https://www.nature.com/

articles/156388a0.

Fisher:1945:RPE

[Fis45f] R. A. Fisher. Recent progress in experimental design. In ????, editor, L’application du calcul des probabilités. Colloque tenu à Genève, 12-15 juillet 1939, Collection Scientifique, pages 19–31. Institut International de Coopération Intellectuelle, Paris, 1945.

Fisher:1946:FPN

[Fis46a] R. A. Fisher. Fisher’s “Problem of the Nile”. Nature, 158(4013): 453, September 28, 1946. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL https://www.nature.com/

articles/158453a0.

Fisher:1946:NCF

[Fis46b] R. A. Fisher. Note on the calculation of the frequencies of rhesus allelomorphs. Annals of Eugenics, 13(1):223–224, January 1, 1946. CODEN ???? ISSN 2050-1420 (print), 2050-1439 (electronic).

URL https://onlinelibrary.wiley.com/doi/abs/10.1111/j.
1469-1809.1946.tb02362.x.

Fisher:1946:SMR

[Fis46c] R. A. Fisher. Statistical Methods for Research Workers, volume 5 of Biological monographs and manuals. Oliver and Boyd, Edinburgh, UK; London, UK, tenth edition, 1946. xiv + 354 pp. LCCN HA29; QH 230.

Fisher:1946:SSL

[Fis46d] R. A. Fisher. A system of scoring linkage data, with special reference to the pied factors in mice. American Naturalist, 80(794):568– 578, September/October 1946. CODEN AMNTA4. ISSN 0003-0147 (print), 1537-5323 (electronic). URL https://www.jstor.

org/stable/2457895.

Fisher:1946:TDB

[Fis46e] R. A. Fisher. Testing the difference between two means of observations of unequal precision. Nature, 158(4020):713–714, November 16, 1946. CODEN NATUAS. ISSN 0028-0836 (print), 1476-

4687 (electronic). URL https://www.nature.com/articles/
158713b0.

Fisher:1946:FGF

[Fis47a] R. A. Fisher. The fitting of gene frequencies to data on Rhesus reactions; with addendum. Note on the calculation of the frequencies of Rhesus allelomorphs. Annals of Eugenics, 13(1):150–155, 223–224, January 1, 1946–1947. ISSN 2050-1420 (print), 2050-1439

(electronic). URL https://onlinelibrary.wiley.com/doi/10.
1111/j.1469-1809.1946.tb02352.x.

Fisher:1947:ACM

[Fis47b] R. A. Fisher. The analysis of covariance method for the relation between a part and the whole. Biometrics, 3(2):65–68, June 1947. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/3001641.

Fisher:1947:DTE

[Fis47c] R. A. Fisher. Development of the theory of experimental design. In ????, editor, Proceedings of the International Statistical Conferences, volume 3, pages 434–439. ????, ????, 1947.

Fisher:1947:MSA

[Fis47d] R. A. Fisher. Les méthodes statistiques adaptées à la recherche scientifique. (French) [Statistical methods adapted to scientific research]. Presses Universitaires de France, Paris, France, 1947. ix + 325 pp.

Fisher:1947:NSS

[Fis47e] R. A. Fisher. Number of self-sterility alleles. Nature, 160(4075): 797–798, December 6, 1947. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL https://www.nature.com/

articles/160797a0.

Fisher:1947:RDa

[Fis47f] R. A. Fisher. The renaissance of Darwinism. The Listener (BBC), 37(??):1001–??, ???? 1947.

Fisher:1947:RDb

[Fis47g] R. A. Fisher. The renaissance of Darwinism, 1947. Radio broadcast.

Fisher:1947:TLP

[Fis47h] R. A. Fisher. The theory of linkage in polysomic inheritance. Philosophical Transactions of the Royal Society of London Series B, Biological sciences, 233(594):55–87, June 19, 1947. CODEN PTRBAE. ISSN 0962-8436 (print), 1471-2970 (elec-

tronic). URL https://royalsocietypublishing.org/doi/10.
1098/rstb.1947.0006.

Fisher:1947:DE

[Fis47i] Ronald Aylmer Fisher. The Design of Experiments. Oliver and Boyd, Edinburgh, UK; London, UK, fourth edition, 1947. xi + 240 pp. LCCN HA29 .F48 1947.

Fisher:1948:B

[Fis48a] R. A. Fisher. Biometry. Biometrics, 4(3):217–219, September 1948. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/3001567.

Fisher:1948:CFF

[Fis48b] R. A. Fisher. Conclusions fiduciaires. (French) [Fiducial conclusions]. Annales de l’Institut Henri Poincaré, 10:191–213, 1948. CODEN AIHPA2. ISSN 0365-320x (print), 2400-4855 (electronic).

URL http://www.numdam.org/item?id=AIHP_1948__10_3_191_
0.

Fisher:1948:MG

[Fis48c] R. A. Fisher. Modern genetics. British Science News, 1(10):2–4, ???? 1948. CODEN BSNEAC. ISSN 0366-3442.

Fisher:1948:QTG

[Fis48d] R. A. Fisher. A quantitative theory of genetic recombination and chiasma formation. Biometrics, 4(1):1–13, March 1948. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/3001694.

Fisher:1948:SMR

[Fis48e] R. A. Fisher. Statistical Methods for Research Workers. Hafner Publishing Company, New York, NY, USA, tenth edition, 1948. xiv + 354 pp. LCCN HA29; QH 230.

Fisher:1948:WSM

[Fis48f] R. A. Fisher. What sort of man is Lysenko? The Listener (BBC), 40(??):874–875, ???? 1948.

Fisher:1949:BAT

[Fis49a] R. A. Fisher. A biological assay of tuberculins. Biometrics, 5(4): 300–316, December 1949. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/3001513.

Fisher:1949:LPT

[Fis49b] R. A. Fisher. The linkage problem in a tetrasomic wild plant, Lythrum salicaria. Hereditas, 35(S1):225–233, 1949. CO-DEN HEREAY. ISSN 0018-0661 (print), 1601-5223 (electronic).

URL https://onlinelibrary.wiley.com/doi/abs/10.1111/j.

1601-5223.1949.tb03335.x. Proceedings of the 8th International Congress of Genetics (Supplement to Hereditas).

Fisher:1949:PSG

[Fis49c] R. A. Fisher. Papers on the Soviet genetics controversy. Occasional Pamphlet 9, Society for Freedom in Science, ????, 1949.

Fisher:1949:PLT

[Fis49d] R. A. Fisher. A preliminary linkage test with Agouti and Undulated mice. 1. The fifth linkage group. Heredity, 3(2):229– 241, August 1, 1949. CODEN HDTYAT. ISSN 0018-067X (print), 1365-2540 (electronic). URL https://www.nature.com/

articles/hdy194916.

Fisher:1949:RRC

[Fis49e] R. A. Fisher. The report of the Royal Commission on Population. Cambridge Journal, 3(??):32–39, ???? 1949.

Fisher:1949:SCS

[Fis49f] R. A. Fisher. The Sub-Commission on Statistical Sampling of the United Nations. Bulletin de l’Institut International de Statistique, 32(2):207–209, ???? 1949.

Fisher:1949:TSS

[Fis49g] R. A. Fisher. A theoretical system of selection for homostyle Primula. Sankhyā (Indian Journal of Statistics), 9(4):325–342, September 1949. CODEN SNKYA5. ISSN 0036-4452. URL https://www.

jstor.org/stable/25047987.

Fisher:1949:DE

[Fis49h] Ronald Aylmer Fisher. The Design of Experiments. Oliver and Boyd, Edinburgh, UK; London, UK, fifth edition, 1949. xi + 242 pp. LCCN HA29 .F48 1949.

Fisher:1949:TIa

[Fis49i] Ronald Aylmer Fisher. The Theory of Inbreeding. Oliver and Boyd, Edinburgh, UK; London, UK, 1949. viii + 120 pp. LCCN S494 .F56 1965.

Fisher:1949:TIb

[Fis49j] Ronald Aylmer Fisher. The Theory of Inbreeding. Academic Press, New York, USA, 1949. viii + 120 pp. LCCN S494 .F56 1965.

Fisher:1950:CEI

[Fis50a] R. A. Fisher. A class of enumerations of importance in genetics. Proceedings of the Royal Society of London. Series B. Biological sciences, 136(886):509–520, January 10, 1950. CODEN PRLBA4. ISSN 0962-8452 (print), 1471-2954 (electronic). URL https://

royalsocietypublishing.org/doi/10.1098/rspb.1950.0002.

Fisher:1950:CAN

[Fis50b] R. A. Fisher. Creative Aspects of Natural Law. The Fourth Arthur Stanley Eddington Memorial Lecture, 2 November 1950, volume 4 of Arthur Stanley Eddington memorial lectures. Cambridge University Press, Cambridge, UK, 1950. v + 23 pp.

Fisher:1950:GFC

[Fis50c] R. A. Fisher. Gene frequencies in a cline determined by selection and diffusion. Biometrics, 6(4):353–361, December 1950. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/3001780.

Fisher:1950:PM

[Fis50d] R. A. Fisher. Polydactyly in mice. Nature, 165(4193):407, March 11, 1950. CODEN NATUAS. ISSN 0028-0836 (print), 1476-

4687 (electronic). URL https://www.nature.com/articles/
165407a0.

Fisher:1950:SDE

[Fis50e] R. A. Fisher. The significance of deviations from expectation in a Poisson series. Biometrics, 6(1):17–24, March 1950. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/3001420.

Fisher:1950:SMR

[Fis50f] R. A. Fisher. Statistical Methods for Research Workers. Oliver and Boyd, Edinburgh, UK; London, UK, eleventh edition, 1950. xv + 354 pp.

Fisher:1950:CMSa

[Fis50g] Ronald Aylmer Fisher. Contributions to Mathematical Statistics. Wiley publications in statistics. Chapman and Hall, Ltd., London, UK, 1950. 339 (est.) pp. LCCN QA276 .F49; QH 211; QH 230. Index prepared by John Tukey.

Fisher:1950:CMSb

[Fis50h] Ronald Aylmer Fisher. Contributions to Mathematical Statistics. Wiley publications in statistics. John Wiley, New York, NY, USA, 1950. xvi + 658 pp. LCCN QA276 .F49; QH 211; QH 230. Index prepared by John Tukey. Edited by Walter A. Shewhart. Biography … of Professor Fisher by P. C. Mahalanobis … reprinted from Sankhyā (vol. 4, pt. 2, pp. 265–272, December, 1938) (9 pages following preface).

Fisher:1951:AQI

[Fis51a] R. A. Fisher. Answer to Query 91 on interaction of quantity and quality in agricultural field trials. Biometrics, 7(4):433–434, December 1951. CODEN BIOMB6. ISSN 0006-341X (print),

1541-0420 (electronic). URL https://www.jstor.org/stable/
3001657.

Fisher:1951:CFM

[Fis51b] R. A. Fisher. A combinatorial formulation of multiple linkage tests. Nature, 167(4248):520, March 31, 1951. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL https://

www.nature.com/articles/167520a0.

Fisher:1951:R

[Fis51c] R. A. Fisher. Reviews. Heredity, 5(1):149–150, April 1, 1951. CO-DEN HDTYAT. ISSN 0018-067X (print), 1365-2540 (electronic).

URL https://www.nature.com/articles/hdy195112.

Fisher:1951:SCE

[Fis51d] R. A. Fisher. Standard calculations for evaluating a blood-group system. Heredity, 5(1):95–102, April 1, 1951. CODEN HDTYAT. ISSN 0018-067X (print), 1365-2540 (electronic). URL https://

www.nature.com/articles/hdy19515.

Fisher:1951:DE

[Fis51e] Ronald Aylmer Fisher. The Design of Experiments. Oliver and Boyd, Edinburgh, UK; London, UK, sixth edition, 1951. xv + 244 pp. LCCN HA29 .F48 1951.

Fisher:1952:SE

[Fis52a] R. A. Fisher. Sequential experimentation. Biometrics, 8(3):183– 187, September 1952. CODEN BIOMB6. ISSN 0006-341X (print),

1541-0420 (electronic). URL https://www.jstor.org/stable/
3001546.

Fisher:1952:SMG

[Fis52b] R. A. Fisher. Statistical methods in genetics. the Bateson Lecture, 1951. Heredity, 6(1):1–12, April 1, 1952. CODEN HDTYAT. ISSN 0018-067X (print), 1365-2540 (electronic). URL https://

www.nature.com/articles/hdy19521.

Fisher:1953:CLP

[Fis53a] R. A. Fisher. Croonian Lecture: Population genetics. Proceedings of the Royal Society of London. Series B. Biological sciences, 141(905):510–523, September 9, 1953. CODEN PRLBA4. ISSN 0962-8452 (print), 1471-2954 (electronic). URL https://

royalsocietypublishing.org/doi/10.1098/rspb.1953.0058.

Fisher:1953:ES

[Fis53b] R. A. Fisher. The expansion of statistics. Journal of the Royal Statistical Society. Series A (General), 116(1):1–10, 1953. CO-DEN JSSAEF. ISSN 0035-9238. URL https://www.jstor.org/

stable/2980946.

Fisher:1953:SDC

[Fis53c] R. A. Fisher. Sex differences of crossing-over in close linkage. American Naturalist, 87(833):116, March/April 1953. CODEN AMNTA4. ISSN 0003-0147 (print), 1537-5323 (electronic). URL

https://www.jstor.org/stable/2458853.

Fisher:1953:S

[Fis53d] R. A. Fisher. Statistics. Sankhyā (Indian Journal of Statistics), 12 (3):303–304, June 1953. CODEN SNKYA5. ISSN 0036-4452. URL

https://www.jstor.org/stable/25048136.

Fisher:1953:DS

[Fis53e] Ronald Fisher. Dispersion on a sphere. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 217(1130):295–305, May 7, 1953. CODEN PRLAAZ. ISSN 0080-4630 (print), 2053-9169 (electronic). URL https:/

/royalsocietypublishing.org/doi/abs/10.1098/rspa.1953.
0064.

Fisher:1953:LPL

[Fis53f] Ronald Fisher. The linkage of polydactyly with leaden in the house-mouse. Heredity, 7(1):91–95, April 1, 1953. CODEN HDTYAT. ISSN 0018-067X (print), 1365-2540 (electronic). URL

https://www.nature.com/articles/hdy19538.

Fisher:1953:VSH

[Fis53g] Ronald Fisher. The variation in strength of the human blood group P. Heredity, 7(1):81–89, April 1, 1953. CODEN HDTYAT. ISSN 0018-067X (print), 1365-2540 (electronic). URL https://

www.nature.com/articles/hdy19537.

Fisher:1954:CDP

[Fis54a] R. A. Fisher. Contribution to a discussion of a paper on interval estimation by M. A. Creasy. Journal of the Royal Statistical Society. Series B (Methodological), 16(2):212–213, ???? 1954. CO-DEN JSTBAJ. ISSN 0035-9246. URL https://www.jstor.org/

stable/2984045.

Fisher:1954:ESM

[Fis54b] R. A. Fisher. The experimental study of multiple crossing over. In ????, editor, Proceedings of the 9th International Congress on Genetics, pages 227–231. ????, ????, 1954. Supplement Caryologia.

Fisher:1954:SMR

[Fis54c] R. A. Fisher. Statistical Methods for Research Workers. Oliver and Boyd, Edinburgh, UK; London, UK, twelfth edition, 1954. xv + 356 pp.

Fisher:1954:AVV

[Fis54d] Ronald Fisher. The analysis of variance with various binomial transformations. Biometrics, 10(1):130–139, March 1954. CO-DEN BIOMB6. ISSN 0006-341x (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/3001667. See discussion

[Fis54d].

Fisher:1954:DAV

[Fis54e] Ronald Fisher. Discussion of the analysis of variance with various binomial transformations. Biometrics, 10(1):140–151, March 1954. CODEN BIOMB6. ISSN 0006-341x (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/3001668. See [Fis54d].

Fisher:1954:RCT

[Fis54f] Ronald Fisher. Retrospect of the criticisms of the theory of natural selection. In Huxley et al. [HHF54], pages 84–98. LCCN QH367 .H96.

Fisher:1954:FTJ

[Fis54g] Ronald A. Fisher. A fuller theory of “junctions” in inbreeding. Heredity, 8(2):187–197, August 1, 1954. CODEN HDTYAT. ISSN 0018-067X (print), 1365-2540 (electronic). URL https://www.

nature.com/articles/hdy195417.

Fisher:1955:AQE

[Fis55a] R. A. Fisher. Answer to Query 114 on the effect of errors of grouping in an analysis of variance. Biometrics, 11(2):237–238, June 1955. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL https://www.jstor.org/stable/3001800.

Fisher:1955:E

[Fis55b] R. A. Fisher. La estadística. (Spanish) [Statistics]. Trabajos de Estadística, 6(??):33–34, February 1955. ISSN 0210-5675.

Fisher:1955:SC

[Fis55c] R. A. Fisher. Science and Christianity. Friend, 113(42):??, ????

1955. Radio broadcast.

Fisher:1955:R

[Fis55d] R. B. Fisher. Reviews. Heredity, 9(2):278, August 1, 1955. CODEN HDTYAT. ISSN 0018-067X (print), 1365-2540 (electronic). URL

https://www.nature.com/articles/hdy195528.

Fisher:1955:SMS

[Fis55e] Ronald Fisher. Statistical methods and scientific induction. Journal of the Royal Statistical Society. Series B (Methodological), 17(1):69–78, 1955. CODEN JSTBAJ. ISSN 0035-9246.

URL http://links.jstor.org/sici?sici=0035-9246(1955)
17:1<69:SMASI>2.0.CO;2-M&origin=MSN;   https://onlinelibrary.
wiley.com/doi/abs/10.1111/j.2517-6161.1955.tb00180.x.

Fisher:1956:R

[Fis56a] R. A. Fisher. Reviews. Heredity, 10(2):275, August 1, 1956. CO-DEN HDTYAT. ISSN 0018-067X (print), 1365-2540 (electronic).

URL https://www.nature.com/articles/hdy195625.

Fisher:1956:SMW

[Fis56b] R. A. Fisher. Statistische Methoden für die Wissenschaft. (German) [Statistical Methods for Research Workers]. Oliver and Boyd, Edinburgh, UK; London, UK, twelfth edition, 1956. xiv + 359 pp. Translation to German by Dora Lucka of [Fis54c].

Fisher:1956:TSP

[Fis56c] Ronald Fisher. On a test of significance in Pearson’s Biometrika Tables (No. 11). Journal of the Royal Statistical Society. Series B (Methodological), 18(1):56–60, 1956. CODEN JSTBAJ.

ISSN 0035-9246. URL http://links.jstor.org/sici?sici=
0035-9246(1956)18:1<56:OATOSI>2.0.CO;2-R&origin=MSN;
https://onlinelibrary.wiley.com/doi/abs/10.1111/j.2517-
6161.1956.tb00209.x.

Fisher:1956:SMSa

[Fis56d] Sir Ronald Aylmer Fisher. Statistical Methods and Scientific Inference. Oliver and Boyd, Edinburgh, UK; London, UK, 1956. viii + 175 pp.

Fisher:1956:SMSb

[Fis56e] Sir Ronald Aylmer Fisher. Statistical Methods and Scientific Inference. Hafner Publishing Company, New York, NY, USA, 1956. 175 pp. LCCN QA9 .F54 1956.

Fisher:1956:BGP

[Fis57a] R. A. Fisher. Blood groups and population genetics. Acta Genetica et Statistica Medica, 6(4):507–509, 1956–1957. URL https://www.

karger.com/Article/Abstract/150883.

Fisher:1957:CNN

[Fis57b] R. A. Fisher. Comment on the notes by Neyman, Bartlett, and Welch in this Journal (vol. 18, no. 2, 56–60, 1956). Journal of the Royal Statistical Society. Series B (Methodological), 19(1):179, ???? 1957. CODEN JSTBAJ. ISSN 0035-

9246. URL http://www.jstor.org/stable/2984005;   https:
//onlinelibrary.wiley.com/doi/abs/10.1111/j.2517-6161.

1957.tb00255.x. See [Ney56, Bar56, Wel56].

Fisher:1957:DCS

[Fis57c] R. A. Fisher. Dangers of cigarette-smoking. British Medical Journal, 2(5039):43, 297–298, August 3, 1957. CODEN BMJOAE.

ISSN 0007-1447. URL https://www.bmj.com/content/2/5039/
297.3.

Fisher:1957:MHG

[Fis57d] R. A. Fisher. Methods in human genetics. Acta Genetica et Statistica Medica, 7(1):7–10, 1957. URL https://www.karger.com/

Article/Abstract/150915.

Fisher:1957:UP

[Fis57e] R. A. Fisher. The underworld of probability. Sankhyā (Indian Journal of Statistics), 18(3–4):201–210, September 1957. CO-DEN SNKYA5. ISSN 0036-4452. URL https://www.jstor.org/

stable/25048352.

Fisher:1958:DI

[Fis58a] R. A. Fisher. The discontinuous inheritance. The Listener (BBC), 60(??):85–87, ???? 1958. Broadcast of 1 July 1958.

Fisher:1958:GTN

[Fis58b] R. A. Fisher. The Genetical Theory of Natural Selection. Dover books on the biological sciences. Dover, New York, NY, USA, second edition, 1958. ISBN 0-486-60466-7. xiv + 291 pp. LCCN QH366 .F5 1958.

Fisher:1958:LD

[Fis58c] R. A. Fisher. The life of Darwin. Nature, 182(4628):71, July 12, 1958. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (elec-

tronic). URL https://www.nature.com/articles/182071a0.

Fisher:1958:MPN

[Fis58d] R. A. Fisher. Mathematical probability in the natural sciences. In Proceedings of the 18th International Congress on Pharmaceutical Science, page ?? ????, ????, 1958.

Fisher:1958:NP

[Fis58e] R. A. Fisher. The nature of probability. Centennial Review, 2(??): 261–274, ???? 1958.

Fisher:1958:PNSa

[Fis58f] R. A. Fisher. Polymorphism and natural selection. Bulletin de l’Institut International de Statistique, 36(??):284–289, ???? 1958.

Fisher:1958:PNSb

[Fis58g] R. A. Fisher. Polymorphism and natural selection. Journal of Ecology, 46(2):289–293, July 1958. CODEN ECOABT. ISSN 0022-0477 (print), 1365-2745 (electronic). URL https://www.jstor.

org/stable/2257396.

Fisher:1958:SMR

[Fis58h] R. A. Fisher. Statistical Methods for Research Workers. Oliver and Boyd, Edinburgh, UK; London, UK, thirteenth edition, 1958. xv + 356 pp.

Fisher:1958:RCT

[Fis58i] Ronald Fisher. Retrospect of the criticisms of the theory of natural selection. In Huxley et al. [HHF58], pages 84–98. LCCN QH367 .H96.

Fisher:1958:CS

[Fis58j] Ronald A. Fisher. Cancer and smoking. Nature, 182(4635): 596, August 30, 1958. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL https://www.nature.com/

articles/182596a0.

Fisher:1958:LCC

[Fis58k] Ronald A. Fisher. Lung cancer and cigarettes? Nature, 182 (4628):108, July 12, 1958. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL https://www.nature.com/

articles/182108a0.

Fisher:1959:MPNa

[Fis59a] Prof. Sir Ronald A. Fisher. Mathematical probability in the natural sciences. Metrika. International Journal for Theoretical and Applied Statistics, 2(1):1–10, ???? 1959. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926x (electronic). URL http://link.

springer.com/article/10.1007/BF02613720.

Fisher:1959:CCS

[Fis59b] R. A. Fisher. Cigarettes, cancer, and statistics. Centennial Review, 2(??):151–166, ???? 1959.

Fisher:1959:NSG

[Fis59c] R. A. Fisher. Natural selection from the genetical standpoint. Australian Journal of Science, 22(??):16–17, ???? 1959. CODEN AJSIA9. ISSN 0365-3668.

Fisher:1959:SCC

[Fis59d] R. A. Fisher. Smoking: the Cancer Controversy, Some Attempts to Assess the Evidence. Oliver and Boyd, Edinburgh, UK; London, UK, 1959. 47 pp. LCCN QH 252.

Fisher:1959:AEE

[Fis59e] Ronald A. Fisher. An algebraically exact examination of junction formation and transmission in parent–offspring inbreeding. Heredity, 13(2):179–186, May 1, 1959. CODEN HDTYAT. ISSN 0018-067X (print), 1365-2540 (electronic). URL https://www.nature.

com/articles/hdy195921.

Fisher:1959:SMS

[Fis59f] Ronald A. Fisher. Statistical Methods and Scientific Inference. 2nd ed., revised. Hafner Publishing Company, New York, NY, USA, second edition, 1959. viii + 178 pp. LCCN QA9 .F54 1959.

Fisher:1959:MPNb

[Fis59g] Sir Ronald A. Fisher. Mathematical probability in the natural sciences. Technometrics, 1(1):21–29, February 1959. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266307.

Fisher:1960:MPN

[Fis60a] R. A. Fisher. Mathematical probability in the natural sciences. La Scuola in Azione, 20(??):5–19, ???? 1960.

Fisher:1960:SEB

[Fis60b] R. A. Fisher. On some extensions of Bayesian inference proposed by Mr Lindley. Journal of the Royal Statistical Society. Series B (Methodological), 22(2):299–301, ???? 1960. CODEN JST-

BAJ. ISSN 0035-9246. URL http://links.jstor.org/sici?
sici=0035-9246(1960)22:2<299:OSEOBI>2.0.CO;2-U&origin=
MSN;   http://www.jstor.org/stable/2984098;   https://
onlinelibrary.wiley.com/doi/abs/10.1111/j.2517-6161.1960.
tb00374.x.

Fisher:1960:DE

[Fis60c] Sir Ronald Aylmer Fisher. The Design of Experiments. Oliver and Boyd, Edinburgh, UK; London, UK, seventh edition, 1960. xiv + 245 pp.

Fisher:1961:MGS

[Fis61a] R. A. Fisher. A model for the generation of self-sterility alleles. Journal of Theoretical Biology, 1(4):411–414, October 1961. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL https://pubmed.ncbi.nlm.nih.gov/13893275/.

Fisher:1961:PDW

[Fis61b] R. A. Fisher. Possible differentiation in the wild population of Oenothera Organensis. Australian Journal of Biological Sciences, 14(1):76–78, 1961. CODEN AJBSAM. ISSN 0004-9417. URL

https://www.publish.csiro.au/bi/BI9610076.

Fisher:1961:SRS

[Fis61c] Ronald A. Fisher. Sampling the reference set. Sankhyā (Indian Journal of Statistics), Series A. Methods and Techniques, 23(1): 3–8, February 1961. CODEN SANABS. ISSN 0581-572X. URL

https://www.jstor.org/stable/25049128.

Fisher:1961:WMT

[Fis61d] Ronald A. Fisher. The weighted mean of two normal samples with unknown variance ratio. Sankhyā (Indian Journal of Statistics), Series A. Methods and Techniques, 23(2):103–114, May 1961. CO-DEN SANABS. ISSN 0581-572X. URL https://www.jstor.org/

stable/25049139.

Fisher:1960:STR

[Fis61e] Ronald Aylmer Fisher. Scientific thought and the refinement of human reasoning. Journal of the Operations Research Society of Japan, 3:1–10, 1960–1961. CODEN JORJA5. ISSN 0453-4514 (print), 1878-6871 (electronic).

Fisher:1962:DSD

[Fis62a] R. A. Fisher. The detection of a sex difference in recombination values using double heterozygotes. Journal of Theoretical Biology, 3(3):509–513, November 1962. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519362800423.

Fisher:1962:RPD

[Fis62b] R. A. Fisher. Il ruolo del ‘piano degli esperimenti’ nella logica della inferenza scientifica. (Italian) [The role of the ‘experiment plan’ in the logic of scientific inference]. La Scuola in Azione, 9(??):33–42, ???? 1962.

Fisher:1962:PDE

[Fis62c] R. A. Fisher. The place of the design of experiments in the logic of scientific inference. Colloques Internationaux du Centre National de la Recherche Scientifique, ??(??):13–19, ???? 1962. CODEN COINAV. ISSN 0366-7634.

Fisher:1962:SEB

[Fis62d] Ronald Fisher. Some examples of Bayes’ method of the experimental determination of probabilities a priori. Journal of the Royal Statistical Society. Series B (Methodological), 24(1):118–124, 1962. CODEN JSTBAJ. ISSN 0035-

9246. URL http://links.jstor.org/sici?sici=0035-9246(
1962)24:1<118:SEOBMO>2.0.CO;2-D&origin=MSN;   https://
onlinelibrary.wiley.com/doi/abs/10.1111/j.2517-6161.1962.
tb00443.x.

Fisher:1962:ECP

[Fis62e] Ronald A. Fisher. Enumeration and classification in polysomic inheritance. Journal of Theoretical Biology, 2(3):309–311, May 1962. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519362900334.

Fisher:1962:LES

[Fis62f] Ronald A. Fisher. Letter to the Editor: Self-sterility alleles: a reply to Professor D. Lewis. Journal of Theoretical Biology, 3(1): 146–147, July 1962. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519362800113.

Fisher:1962:SDC

[Fis62g] Ronald A. Fisher. The simultaneous distribution of correlation coefficients. Sankhyā (Indian Journal of Statistics), Series A. Methods and Techniques, 24(1):1–8, February 1962. CODEN SAN-

ABS. ISSN 0581-572X. URL https://www.jstor.org/stable/
25049187.

Fisher:1962:CLC

[Fis62h] Sir Ronald A. Fisher. Confidence limits for a cross-product ratio. Australian Journal of Statistics, 4(1):41, April 1962. CODEN AUJSA3. ISSN 0004-9581.

Fisher:1963:SMR

[Fis63] R. A. Fisher. Statistical Methods for Research Workers. Oliver and Boyd, Edinburgh, UK; London, UK, thirteenth edition, 1963. xiii + 356 pp.

Fisher:1964:B

[Fis64a] R. A. Fisher. Biometry. Biometrics, 20(2):261–264, June 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.jstor.org/stable/2528394.

Fisher:1964:QTG

[Fis64b] R. A. Fisher. A quantitative theory of genetic recombination and chiasma formation. Biometrics, 20(2):253–260, June 1964. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2528393.

Fisher:1964:SDE

[Fis64c] R. A. Fisher. The significance of deviations from expectation in a Poisson series. Biometrics, 20(2):265–272, June 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528395.

Fisher:1965:PDE

[Fis65a] R. A. Fisher. The place of the design of experiments in the logic of scientific inference. Sankhyā (Indian Journal of Statistics), Series A. Methods and Techniques, 27(1):33–38, March 1965. CO-DEN SANABS. ISSN 0036-4452. URL https://www.jstor.org/

stable/25049350.

Fisher:1965:TI

[Fis65b] Ronald Aylmer Fisher. The Theory of Inbreeding. Oliver and Boyd, Edinburgh, UK; London, UK, second edition, 1965. viii + 150 pp. LCCN S494 .F56 1965.

Fisher:1966:DE

[Fis66] Sir Ronald Aylmer Fisher. The Design of Experiments. Oliver and Boyd, Edinburgh, UK; London, UK, eighth edition, 1966. xv + 248 pp. LCCN HA29 .F48 1966.

Fisher:1967:SMR

[Fis67] R. A. Fisher. Statistical Methods for Research Workers. Oliver and Boyd, Edinburgh, UK; London, UK, thirteenth edition, 1967. xv + 356 pp.

Fisher:1970:SMRa

[Fis70a] R. A. Fisher. Statistical Methods for Research Workers. Oliver and Boyd, Edinburgh, UK; London, UK, fourteenth edition, 1970. xiii + 362 pp.

Fisher:1970:SMRb

[Fis70b] R. A. Fisher. Statistical Methods for Research Workers. Hafner Publishing Company, New York, NY, USA, fourteenth edition, 1970. xv + 362 pp.

Fisher:1971:DE

[Fis71] Sir Ronald Aylmer Fisher. The Design of Experiments. Hafner Publishing Company, New York, NY, USA, eighth edition, 1971. xv + 248 pp. LCCN HA29 .F48 1971. URL https://www.phil.

vt.edu/dmayo/PhilStatistics/b%20Fisher%20design%20of%20experiments.
pdf.

Fisher:1973:SMS

[Fis73a] Ronald A. Fisher. Statistical Methods and Scientific Inference. Hafner Press and Collier Macmillan Publishers, New York, NY, USA and London, UK, third edition, 1973. ISBN 0-02-844740-9. viii + 182 pp.

Fisher:1973:SMR

[Fis73b] Ronald A. Fisher. Statistical Methods for Research Workers. Hafner Publishing Company, New York, NY, USA, fourteenth edition, 1973. xv + 362 pp.

Fisher:1974:RFS

[Fis74] R. A. Fisher. The Rhesus factor: a study in scientific method. American Scientist, 35(1):195–103, ???? 1974. CODEN AMSCAC. ISSN 0003-0996 (print), 1545-2786 (electronic).

Fisher:1990:DR

[Fis90] R. A. Fisher. On the dominance ratio. Bulletin of Mathematical Biology, 52(1–2):297–318, January 1990. CODEN BMTBAP.

ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://

link.springer.com/article/10.1007/BF02459576.

Fisher:1992:AFE

[Fis92a] R. A. Fisher. The arrangement of field experiments. In Kotz and Johnson [KJ92b], pages 82–91. ISBN 0-387-94039-1 (New York: v. 2: softcover), 0-387-97572-1 (New York: v. 2: hardcover), 3-540-94039-1 (Berlin: v. 2: softcover), 3-540-97572-1 (Berlin: v. 2: hardcover). LCCN QA276 .B68465 1992. URL http://www.loc.gov/

catdir/enhancements/fy0815/93003854-d.html;   http://www.
loc.gov/catdir/enhancements/fy0815/93003854-t.html.

Fisher:1992:MFT

[Fis92b] R. A. Fisher. On the mathematical foundations of theoretical statistics. In Kotz and Johnson [KJ92a], pages 11–44. ISBN 0-387-94037-5 (New York: v. 1: softcover), 0-387-97566-7 (New York: v. 1: hardcover). LCCN QA276 .B68465 1992. URL http://www.

loc.gov/catdir/enhancements/fy0815/93003854-d.html;   http:
//www.loc.gov/catdir/enhancements/fy0815/93003854-t.html.

Fisher:1992:SMR

[Fis92c] R. A. Fisher. Statistical methods for research workers. In Kotz and Johnson [KJ92b]. ISBN 0-387-94039-1 (New York: v. 2: softcover), 0-387-97572-1 (New York: v. 2: hardcover), 3-540-94039-1 (Berlin: v. 2: softcover), 3-540-97572-1 (Berlin: v. 2: hardcover).

LCCN QA276 .B68465 1992. URL http://www.loc.gov/catdir/
enhancements/fy0815/93003854-d.html;   http://www.loc.
gov/catdir/enhancements/fy0815/93003854-t.html.

Fisher:1999:GTN

[Fis99] R. A. Fisher. The Genetical Theory of Natural Selection. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, variorum edition, 1999. ISBN 0-19-850440-3. xxii + xiv + 318 pp. LCCN QH366 .F5 1999; QH 366 F535 1999. Revised reprint of the 1930 original. Edited, with a foreword and notes, by J. H. Bennett.

Fisher:2008:MWB

[Fis08] R. A. Fisher. Has Mendel’s work been rediscovered? In Franklin et al. [F^{+}08], page ?? ISBN 0-8229-4319-0 (hardcover), 0-8229-5986-0 (paperback). LCCN QH428 .E53 2008. URL http:

//www.loc.gov/catdir/enhancements/fy0805/2007041848-b.
html;   http://www.loc.gov/catdir/enhancements/fy0805/

2007041848-d.html;   http://www.loc.gov/catdir/toc/ecip082/
2007041848.html.

Fisher:1932:GIS

[FIT32] R. A. Fisher, F. R. Immer, and Olof Tedlin. The genetical interpretation of statistics of the third degree in the study of quantitative inheritance. Genetics, 17(2):107–124, March 1932. CODEN GENTAE. ISSN 0016-6731 (print), 1943-2631 (electronic). URL

https://www.genetics.org/content/17/2/107.

Fisher:1947:SCH

[FLO47] R. A. Fisher, M. F. Lyon, and A. R. G. Owen. The sex chromosome in the house mouse. Heredity, 1(3):355–365, December 1, 1947. CODEN HDTYAT. ISSN 0018-067X (print), 1365-2540 (elec-

tronic). URL https://www.nature.com/articles/hdy194723.

Fisher:1922:CWR

[FM22] R. A. Fisher and Winifred A. MacKenzie. The correlation of weekly rainfall. Quarterly journal of the Royal Meteorological Society, 48 (203):234–242, ???? 1922. CODEN QJRMAM. ISSN 0035-9009 (print), 1477-870X (electronic). URL https://digital.library.

adelaide.edu.au/dspace/handle/2440/15175.

Fisher:1923:SCV

[FM23] R. A. Fisher and W. A. MacKenzie. Studies in crop variation. II. The manurial response of different potato varieties. Journal of Agricultural Science, 13(3):311–320, ???? 1923. CO-DEN JASIAB. ISSN 0021-8596 (print), 1469-5146 (electronic).

URL https://hdl.handle.net/2440/15179;   https://www.
cambridge.org/core/journals/journal-of-agricultural-science/
article/studies-in-crop-variation-ii-the-manurial-response-
of-different-potato-varieties/326342EA3B72577737653A3D54741D46.

Fisher:1936:LTM

[FM36a] R. A. Fisher and K. Mather. A linkage test with mice. Annals of Eugenics, 7(3):265–280, November 1, 1936. ISSN 2050-1420 (print), 2050-1439 (electronic). URL https://onlinelibrary.

wiley.com/doi/10.1111/j.1469-1809.1936.tb02145.x.

Fisher:1936:VMP

[FM36b] R. A. Fisher and K. Mather. Verification in mice of the possibility of more than fifty per cent recombination. Nature, 137(3461): 362–363, February 1, 1936. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL https://www.nature.com/

articles/137362a0.

Fisher:1940:NLM

[FM40] R. A. Fisher and K. Mather. Non-lethality of the Mid factor in Lythrum salicaria. Nature, 146(3703):521, October 19, 1940. CO-DEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

URL https://www.nature.com/articles/146521a0.

Fisher:1942:PIL

[FM42] R. A. Fisher and K. Mather. Polyploid inheritance in Lythrum salicaria. Nature, 150(3806):430, October 10, 1942. CODEN NAT-UAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL https:

//www.nature.com/articles/150430a0.

Fisher:1943:ISL

[FM43] R. A. Fisher and K. Mather. The inheritance of style length in Lythrum salicaria. Annals of Eugenics, 12(1):1–13, January 1, 1943. ISSN 2050-1420 (print), 2050-1439 (elec-

tronic).  URL https://onlinelibrary.wiley.com/doi/10.
1111/j.1469-1809.1943.tb02307.x.

Fisher:1947:SOL

[FM47] R. A. Fisher and V. C. Martin. Spontaneous occurrence in Lythrum salicaria of plants duplex for the short-style gene. Nature, 160 (4068):541, October 18, 1947. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL https://www.nature.com/

articles/160541a0.

Fisher:1948:GSL

[FM48] R. A. Fisher and V. C. Martin. Genetics of style-length in Oxalis. Nature, 162(4118):533, October 2, 1948. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL https://

www.nature.com/articles/162533a0.

Fisher:1957:STA

[FM57] R. A. Fisher and W. H. McCrea. Space travel and ageing. Discovery, 18(??):56–58, 174–175, ???? 1957.

Fisher:1984:ING

[FNS84] Robert A. Fisher, Michael Martin Nieto, and Vernon D. Sandberg. Impossibility of naively generalizing squeezed coherent states. Physical Review D (Particles and Fields), 29(6):1107–1110, 1984. CODEN PRVDAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-4500 (cd-rom).

Fisher:1924:TMA

[FO24] R. A. Fisher and S. Odén. The theory of the mechanical analysis of sediments by means of the automatic balance. Proceedings of the Royal Society of Edinburgh, 44(??):98–115, ???? 1924. CODEN PRSEAE. ISSN 0080-4541 (print), 2053-5902 (electronic).

Fisher:1946:RGF

[FR46] R. A. Fisher and R. R. Race. Rh gene frequencies in Britain. Nature, 157(3976):48–49, January 12, 1946. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL https://

www.nature.com/articles/157048b0.

Freeman:1983:CRB

[Fre83] G. H. Freeman. Consultancy and research in biometrics — some thoughts inspired by reading the biography of R. A. Fisher. Bulletin in Applied Statistics (BIAS), 10(2):138–157, 1983. CODEN ???? ISSN ????

Fisher:1944:MRR

[FRT44] R. A. Fisher, R. R. Race, and G. L. Taylor. Mutation and the rhesus reaction. Nature, 153(3873):106, January 22, 1944. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL

https://www.nature.com/articles/153106b0.

Fisher:1948:TLG

[FS48] R. A. Fisher and G. D. Snell. A twelfth linkage group of the house mouse. Heredity, 2(2):271–273, September 1, 1948. CODEN HDTYAT. ISSN 0018-067X (print), 1365-2540 (electronic). URL

https://www.nature.com/articles/hdy194816.

Fisher:1928:LFF

[FT28] R. A. Fisher and L. H. C. Tippett. Limiting forms of the frequency distribution of the largest of smallest member of a sample. Proceedings of the Cambridge Philosophical Society. Mathematical and physical sciences, 24(2):180–190, April 1928. CODEN PCPSA4.

ISSN 0008-1981. URL https://www.cambridge.org/core/
journals/mathematical-proceedings-of-the-cambridge-philosophical-
society/article/limiting-forms-of-the-frequency-distribution-
of-the-largest-or-smallest-member-of-a-sample/7BE8DE65FCDFC3ABECFE1054DFB56CB5.

Fisher:1939:BGG

[FT39] R. A. Fisher and G. L. Taylor. Blood groups in Great Britain. British Medical Journal, 2(4111):826, October 21, 1939. CO-DEN BMJOAE. ISSN 0007-1447. URL https://www.jstor.org/

stable/20314371.

Fisher:1940:SIS

[FT40] R. A. Fisher and G. L. Taylor. Scandinavian influence in Scottish ethnology. Nature, 145(3676):590, April 13, 1940. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL

https://www.nature.com/articles/145590a0.

Fienberg:1996:RFC

[FT96] Stephen E. Fienberg and Judith M. Tanur. Reconsidering the fundamental contributions of Fisher and Neyman on experimentation and sampling. International Statistical Review = Revue Internationale de Statistique, 64(3):237–253, December 1996. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL

http://www.jstor.org/stable/1403784.

Fisher:1922:APM

[FTM22] R. A. Fisher, H. G. Thornton, and W. A. MacKenzie. The accuracy of the plating method of estimating the density of bacterial populations. Annals of Applied Biology, 9(3–4): 325–359, November 1922. CODEN AABIAV. ISSN 0003-4746 (print), 1744-7348 (electronic). URL https://hdl.

handle.net/2440/F15176;   https://onlinelibrary.wiley.
com/doi/abs/10.1111/j.1744-7348.1922.tb05962.x.

Fisher:1939:SBG

[FV39] R. A. Fisher and Janet Vaughan. Surnames and blood-groups. Nature, 144(3660):1047–1048, December 23, 1939. CODEN NAT-UAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL https:

//www.nature.com/articles/1441047c0.

Fisher:1930:AFE

[FW30] R. A. Fisher and J. Wishart. The arrangement of field experiments and the statistical reduction of the results. Imperial Bureau of Soil Science Technical Communication, 10(??):??, ???? 1930.

Fisher:1931:DPF

[FW31] R. A. Fisher and J. Wishart. The derivation of the pattern formulae of two-way partitions from those of simpler patterns. Proceedings of the London Mathematical Society. Second Series, 33(1): 195–208, 1931. CODEN PLMTAL. ISSN 0024-6115 (print), 1460-

244x (electronic). URL https://londmathsoc.onlinelibrary.
wiley.com/doi/abs/10.1112/plms/s2-33.1.195.

Fisher:1934:LS

[FY34] R. A. Fisher and F. Yates. The 6 × 6 Latin squares. Mathematical Proceedings of the Cambridge Philosophical Society, 30(4):492–507, October 1934. CODEN MPCPCO. ISSN 0305-0041 (print), 1469-8064 (electronic).

Fisher:1938:STB

[FY38] Ronald Aylmer Fisher and Frank Yates. Statistical Tables for Biological, Agricultural and Medical Research. Oliver and Boyd, Edinburgh, UK; London, UK, 1938. viii + 90 + 1 pp. LCCN HA33 .F53.

Fisher:1943:STB

[FY43] Ronald Aylmer Fisher and Frank Yates. Statistical Tables for Biological, Agricultural, and Medical Research. Oliver and Boyd, Edinburgh, UK; London, UK, second edition, 1943. viii + 98 pp. LCCN HA33 .F53 1943.

Fisher:1944:FFL

[FY44] R. A. Fisher and F. Yates. Über S-fastperiodische Funktionen mit linear unabhängigen Ereporenten. (German) [On S-strongly periodic functions with linear independent erporents]. Norsk Matematisk Tidsskrift, 26(??):33–40, ???? 1944. ISSN 2387-2187.

Fisher:1948:STB

[FY48] Ronald A. Fisher and Frank Yates. Statistical Tables for Biological, Agricultural and Medical Research. Oliver and Boyd, Edinburgh, UK; London, UK, third edition, 1948. viii + 112 pp.

Fisher:1949:STB

[FY49] Ronald Aylmer Fisher and Frank Yates. Statistical Tables for Biological, Agricultural and Medical Research. Hafner Publishing Company, New York, NY, USA, third edition, 1949. viii + 112 pp. LCCN HA33 .F53 1949.

Fisher:1953:STBa

[FY53a] Sir Ronald Aylmer Fisher and Frank Yates. Statistical Tables for Biological, Agricultural and Medical Research. Oliver and Boyd, Edinburgh, UK; London, UK, fourth edition, 1953. xi + 126 pp. LCCN QA276 .F498 1953a.

Fisher:1953:STBb

[FY53b] Sir Ronald Aylmer Fisher and Frank Yates. Statistical Tables for Biological, Agricultural and Medical Research. Hafner Publishing Company, New York, NY, USA, fourth edition, 1953. 126 pp. LCCN QA276 .F498 1953.

Fisher:1957:STBa

[FY57a] Sir Ronald Aylmer Fisher and Frank Yates. Statistical Tables for Biological, Agricultural and Medical Research. Oliver and Boyd, Edinburgh, UK; London, UK, fifth edition, 1957. x + 138 pp. LCCN QA276 .F498 1957a.

Fisher:1957:STBb

[FY57b] Sir Ronald Aylmer Fisher and Frank Yates. Statistical Tables for Biological, Agricultural and Medical Research. Hafner Publishing Company, New York, NY, USA, fifth edition, 1957. x + 138 pp. LCCN QA276 .F498 1957.

Fisher:1963:STB

[FY63] Sir Ronald Aylmer Fisher and Frank Yates. Statistical Tables for Biological, Agricultural and Medical Research. Oliver and Boyd, Edinburgh, UK; London, UK, sixth edition, 1963. ISBN 0-05-000872-2, 0-582-44525-6. x + 146 pp. LCCN QH324 .F52 1963.

Gani:1982:MS

[Gan82] J. M. (Joseph Mark) Gani, editor. The Making of Statisticians. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1982. ISBN 0-387-90684-3 (New York), 3-540-90684-3 (Berlin). viii + 263 pp. LCCN QA276.156 .M34 1982.

Geisser:1992:IFM

[Gei92] Seymour Geisser. Introduction to Fisher (1922) On the Mathematical Foundations of Theoretical Statistics. In Kotz and Johnson [KJ92a], pages 1–10. ISBN 0-387-94037-5 (New York: v. 1: softcover), 0-387-97566-7 (New York: v. 1: hardcover).

LCCN QA276 .B68465 1992. URL http://www.loc.gov/catdir/
enhancements/fy0815/93003854-d.html;   http://www.loc.
gov/catdir/enhancements/fy0815/93003854-t.html.

Granell:2011:MAN

[GGA11] Clara Granell, Sergio Gómez, and Alex Arenas. Mesoscopic analysis of networks: Applications to exploratory analysis and data clustering. Chaos (Woodbury, NY), 21(1):016102:1–016102:9, March 2011. CODEN CHAOEH. ISSN 1054-1500.

Green:2003:NLW

[Gre03] Peter J. Green. Notes on the life and work of R. A. Fisher. Journal of the Royal Statistical Society. Series D (The Statistician), 52(3): 299–301, ???? 2003. CODEN ???? ISSN 0039-0526 (print), 1467-

9884 (electronic). URL http://www.jstor.org/stable/4128204.

Gridgeman:1982:BRR

[Gri82] N. T. Gridgeman. Book review: R. A. Fisher. The life of a scientist: By Joan Fisher Box. New York (John Wiley & Sons), 1978. xii + 512 pp. Illustrated. Historia Mathematica, 9(1):104– 106, February 1982. CODEN HIMADS. ISSN 0315-0860 (print),

1090-249X (electronic). URL http://www.sciencedirect.com/
science/article/pii/0315086082901483.

Grove:1930:RPR

[Gro30] Charles C. Grove. Recent publications: Reviews: Statistical Methods for Research Workers, by R. A. Fisher. American Mathematical Monthly, 37(10):547–550, December 1930. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Gigerenzer:1989:ECH

[GSP^{+}89] Gerd Gigerenzer, Zeno Swijtink, Theodore Porter, Lorraine Daston, John Beatty, and Lorenz Krüger. The Empire of Chance: How Probability Changed Science and Everyday Life. Ideas in context. Cambridge University Press, Cambridge, UK, 1989. ISBN 0-511-72048-3 (e-book), 0-521-33115-3 (hardcover), 0-521-39838-X (paperback), 1-107-38722-1 (e-book), 1-4619-4906-8 (e-book). xvii + 340 pp. LCCN QA273 .E57 1989. URL http://www.loc.

gov/catdir/description/cam023/88016928.htlm;   http://
www.loc.gov/catdir/toc/cam022/88016928.html.

Gupta:1960:OSG

[Gup60] Shanti S. Gupta. Order statistics from the gamma distribution. Technometrics, 2(2):243–262, May 1960. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1266548. See errata [Ano60].

Garay:1999:RAS

[GV99] József Garay and Zoltán Varga. Relative advantage: a substitute for mean fitness in Fisher’s fundamental theorem? Journal of Theoretical Biology, 201(3):215–218, December 7, 1999. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519399910202.

H:1931:BRBa

[H.31] A. B. H. Book review: The Genetical Theory of Natural Selection, by R. A. Fisher. Journal of the Royal Statistical Society, 94(1): 98–100, ???? 1931. ISSN 0952-8385. URL http://www.jstor.

org/stable/2341823.

Hald:1990:HPS

[Hal90] Anders Hald. A history of probability and statistics and their applications before 1750. Wiley series in probability and mathematical statistics. Probability and mathematical statistics. John Wiley, New York, NY, USA, 1990. ISBN 0-471-50230-

8. ISSN 0271-6232. xiii + 586 pp. LCCN QA273.A4 H35

1990. URL http://www.gbv.de/dms/ilmenau/toc/025661744.
PDF;   http://www.loc.gov/catdir/description/wiley033/
89033269.html;   http://www.loc.gov/catdir/toc/onix04/
89033269.html.

Hald:1998:HMS

[Hal98] Anders Hald. A History of Mathematical Statistics from 1750 to 1930. Wiley series in probability and statistics. Texts and references section. John Wiley, New York, NY, USA, 1998. ISBN 0-471-17912-4. xvii + 795 pp. LCCN QA276.15 .H35 1998. URL

http://www.loc.gov/catdir/bios/wiley041/97019513.html;
 http://www.loc.gov/catdir/description/wiley032/97019513.
html;   http://www.loc.gov/catdir/toc/onix02/97019513.
html;   https://pdfs.semanticscholar.org/c92f/a3fee16bd992e7d873b78e206c1764d50395.
pdf.

Hall:2002:RFR

[Hal02] Nancy Louise Schafer Hall. R. A. Fisher and randomized experimental design. Ph.D. dissertation, University of Maryland, College Park, MD, USA, 2002. viii + 214 pp. URL https://search.

proquest.com/pqdtglobal/docview/288070490.

Hall:2007:RFH

[Hal07] Nancy S. Hall. R. A. Fisher and his advocacy of randomization. Journal of the History of Biology, 40(2):295–325, June 2007. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10739-006-9119-z;   http://link.springer.com/content/
pdf/10.1007/s10739-006-9119-z.pdf.

Hall:2010:RFG

[Hal10] Nancy S. Hall. Ronald Fisher and Gertrude Cox: two statistical pioneers sometimes cooperate and sometimes collide. The American Statistician, 64(3):212–220, 2010. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hartl:2008:AWS

[Har08] Daniel L. Hartl. Amiable Wright and suspicious Fisher on Mendel’s “personal equation”. In Franklin et al. [F^{+}08], page ?? ISBN 0-8229-4319-0 (hardcover), 0-8229-5986-0 (paperback). LCCN QH428 .E53 2008. URL http:

//www.loc.gov/catdir/enhancements/fy0805/2007041848-b.
html;   http://www.loc.gov/catdir/enhancements/fy0805/
2007041848-d.html;   http://www.loc.gov/catdir/toc/ecip082/
2007041848.html.

Healy:1967:BRBc

[Hea67] M. J. R. Healy. Book review: The Design of Experiments, by Ronald A. Fisher. Journal of the Royal Statistical Society. Series A (General), 130(2):250, ???? 1967. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.org/stable/2343406.

Healy:1973:BRBc

[Hea73] M. J. R. Healy. Book review: Collected Papers of R. A. Fisher. Volume I, 1912–1924, by R. A. Bennett. Journal of the Royal Statistical Society. Series A (General), 136(3):451–452, ???? 1973. CODEN JSSAEF. ISSN 0035-9238. URL http://www.jstor.

org/stable/2345001.

Healy:2003:RFS

[Hea03] M. J. R. Healy. R. A. Fisher the statistician. Journal of the Royal Statistical Society. Series D (The Statistician), 52(3):303– 310, ???? 2003. CODEN ???? ISSN 0039-0526 (print), 1467-9884

(electronic). URL http://www.jstor.org/stable/4128205.

Huxley:1954:EPE

[HHF54] Julian Huxley, A. C. Hardy, and H. B. Ford, editors. Evolution as a process. [Essays]. Allen and Unwin, London, UK, 1954. 367 pp. LCCN QH367 .H96.

Huxley:1958:EPE

[HHF58] Julian Huxley, A. C. Hardy, and H. B. Ford, editors. Evolution as a process. [Essays]. Allen and Unwin, London, UK, second edition, 1958. 367 pp. LCCN QH367 .H96.

Hinkley:1980:RAE

[Hin80] David V. Hinkley. Randomization analysis of experimental data: The Fisher randomization test comment. Journal of the American Statistical Association, 75(371):582–584, September 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287649.

Hotelling:1927:BRR

[Hot27] Harold Hotelling. Book review: R. A. Fisher, Statistical Methods for Research Workers. Journal of the American Statistical Association, 22(159):411–412, September 1927. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2276824.

Hotelling:1951:IRF

[Hot51] Harold Hotelling. The impact of R. A. Fisher on statistics. Journal of the American Statistical Association, 46(253):35–46, March 1951. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (elec-

tronic). URL http://www.jstor.org/stable/2280091.

Howie:1999:IPH

[How99] David John Henry Howie. Interpretations of probability, 1919– 1939: Harold Jeffreys, R. A. Fisher, and the Bayesian controversy. Ph.D. dissertation in history and sociology of science, University of Pennsylvania, Philadelphia, PA 19104, USA, 1999. vii +

309 pp. URL https://repository.upenn.edu/dissertations/
AAI9937734;   https://search.proquest.com/docview/304517544/

.

I:1926:BRBa

[I.26] L. I. Book review: Statistical Methods for Research Workers, by R. A. Fisher. Journal of the Royal Statistical Society, 89(1):144–145, January 1926. ISSN 0952-8385. URL http://www.jstor.org/

stable/2341488.

I:1929:BRB

[I.29] J. O. I. Book review: Statistical Methods for Research Workers, by R. A. Fisher. Journal of the Royal Statistical Society, 92(1): 101–103, ???? 1929. ISSN 0952-8385. URL http://www.jstor.

org/stable/2341440.

I:1940:BRB

[I.40] J. O. I. Book review: The Statistical Theory of Estimation, by R. A. Fisher. Journal of the Royal Statistical Society, 103(2):250, ????

1940. ISSN 0952-8385. URL http://www.jstor.org/stable/
2980419.

Irwin:1963:SRA

[IBM^{+}63] J. O. Irwin, G. A. Barnard, Kenneth Mather, F. Yates, and M. J. R. Healy. Sir Ronald Aylmer Fisher, 1890–1962. Journal of the Royal Statistical Society. Series A (General), 126(1): 159–178, ???? 1963. CODEN JSSAEF. ISSN 0035-9238. URL

http://www.jstor.org/stable/2982477.

Inman:1994:KPR

[Inm94] Henry F. Inman. Karl Pearson and R. A. Fisher on statistical tests: a 1935 exchange from Nature. The American Statistician, 48(1): 2–11, February 1994. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

J:1979:BRBc

[J.79] J. E. J. Book review: R. A. Fisher, the Life of a Scientist by Joan Fisher Box. Technometrics, 21(1):135, February 1979. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1268594.

Jeffreys:1974:FIP

[Jef74] Harold Jeffreys. Fisher and inverse probability. International Statistical Review = Revue Internationale de Statistique, 42(1):1–3, April 1974. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823

(electronic). URL http://www.jstor.org/stable/1402679.

Johnstone:1987:TSF

[Joh87] D. J. Johnstone. Tests of significance following R. A. Fisher. British Journal for the Philosophy of Science, 38(4):481–499, December 1987. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537

(electronic). URL http://bjps.oxfordjournals.org/content/
38/4/481.full.pdf+html;   http://www.jstor.org/stable/
687354.

Jowett:1956:BRS

[Jow56] G. H. Jowett. Book reviews: Statistical Methods for Research Workers, by Ronald A. Fisher. Applied Statistics, 5(1):68–70, March 1956. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Jowett:1957:BRS

[Jow57] G. H. Jowett. Book reviews: Statistical Methods and Scientific Inference, by Ronald A. Fisher. Applied Statistics, 6(3):226–227, November 1957. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

K:1939:BRBb

[K.39] M. G. K. Book review: Statistical Tables for Biological, Agricultural and Medical Research, by R. A. Fisher; F. Yates. Journal of the Royal Statistical Society, 102(2):298, ???? 1939. ISSN 0952-

8385. URL http://www.jstor.org/stable/2980011.

K:1973:BRBb

[K.73] G. G. K. Book review: Collected Papers of R. A. Fisher, Volume I by J. H. Bennett; R. A. Fisher. Technometrics, 15(2):426–427, May 1973. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1267011.

Kanji:1981:BRB

[Kan81] G. K. Kanji. Book review: R. A. Fisher, The Life of a Scientist, by Joan Fisher Box. Journal of the Royal Statistical Society. Series D (The Statistician), 30(2):157–158, June 1981. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2987574.

Karlin:1992:RFE

[Kar92] Samuel Karlin. R. A. Fisher and evolutionary theory. Statistical Science, 7(1):13–33, February 1992. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177011441.

Kruger:1987:PRVa

[KDH87] Lorenz Krüger, Lorraine J. Daston, and Michael Heidelberger, editors. The Probabilistic Revolution, Vol. I: Ideas in History. MIT Press, Cambridge, MA, USA, 1987. ISBN 0-262-11118-7 (hardcover), 0-262-61062-0 (paperback), 0-262-11125-X (set). xiv + 449 pp. LCCN QA273.A4 P76 1987.

Kempthorne:1980:RAE

[Kem80] Oscar Kempthorne. Randomization analysis of experimental data: The Fisher randomization test comment. Journal of the American Statistical Association, 75(371):584–587, September 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287650.

Kemp:1993:BRB

[Kem93] A. W. Kemp. Book review: Statistical Inference and Analysis: Selected Correspondence of R. A. Fisher, by R. A. Fisher; J. H. Bennett. Journal of the Royal Statistical Society. Series D (The Statistician), 42(1):75–76, ???? 1993. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.jstor.

org/stable/2348120.

Kendall:1948:SMP

[Ken48] David G. Kendall. On some modes of population growth leading to R. A. Fisher’s logarithmic series distribution. Biometrika, 35(1/2): 6–15, May 1948. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2332624.

Kendall:1963:RAF

[Ken63] M. G. Kendall. Ronald Aylmer Fisher, 1890–1962. Biometrika, 50 (1/2):1–15 (1 plate), June 1963. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2333741.

Kendall:1970:RAF

[Ken70] David G. Kendall. Ronald Aylmer Fisher (1890–1962). In Pearson and Kendall [PK70], pages 438–453. ISBN 0-85264-193-1. LCCN QA276.15 .P43.

Kruger:1987:PRVb

[KGM87] Lorraine Krüger, Gerd Gigerenzer, and Mary S. Morgan, editors. The Probabilistic Revolution, Vol. II: Ideas in the Sciences. MIT Press, Cambridge, MA, USA, 1987. ISBN 0-262-11119-5 (hardcover), 0-262-61063-9 (paperback), 0-262-11125-X (set). xvi + 459 pp. LCCN QA273.A4 P76 1987.

Kotz:1992:BSVa

[KJ92a] Samuel Kotz and Norman Lloyd Johnson, editors. Breakthroughs in Statistics: Volume 1. Foundations and Basic Theory. Springer series in statistics. Perspectives in statistics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1992. ISBN 0-387-94037-5 (New York: v. 1: softcover), 0-387-97566-7 (New York: v. 1: hardcover). xli + 631 pp.

LCCN QA276 .B68465 1992. URL http://www.loc.gov/catdir/
enhancements/fy0815/93003854-d.html;   http://www.loc.
gov/catdir/enhancements/fy0815/93003854-t.html.

Kotz:1992:BSVb

[KJ92b] Samuel Kotz and Norman Lloyd Johnson, editors. Breakthroughs in Statistics: Volume 2. Methodology and Distribution. Springer series in statistics. Perspectives in statistics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1992. ISBN 0-387-94039-1 (New York: v. 2: softcover), 0-387-97572-1 (New York: v. 2: hardcover), 3-540-94039-1 (Berlin: v. 2: softcover), 3-540-97572-1 (Berlin: v. 2: hardcover). xxi + 600 pp.

LCCN QA276 .B68465 1992. URL http://www.loc.gov/catdir/
enhancements/fy0815/93003854-d.html;   http://www.loc.
gov/catdir/enhancements/fy0815/93003854-t.html.

Kotz:1997:BSV

[KJ97] Samuel Kotz and Norman Lloyd Johnson, editors. Breakthroughs in Statistics: Volume 3. Springer series in statistics. Perspectives in statistics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1997. ISBN 0-387-94988-7 (hardcover), 0-387-94989-5 (softcover). xxv + 559 pp.

LCCN QA276 .B68465 1997. URL http://www.loc.gov/catdir/
enhancements/fy0815/93003854-d.html;   http://www.loc.
gov/catdir/enhancements/fy0815/93003854-t.html.

Kagan:2013:NPS

[KM13] Abram M. Kagan and Yaakov Malinovsky. On the Nile problem by Sir Ronald Fisher. Electron. J. Stat., 7:1968–1982, 2013.

Kendall:1977:SHS

[KP77] Maurice G. (Maurice George) Kendall and R. L. Plackett, editors. Studies in the History of Statistics and Probability, Volume II: a Series of Papers. MacMillan Publishing Company, New York, NY, USA, 1977. ISBN 0-85264-193-1. 481 pp. LCCN QA276.15 .P431x.

Lane:1980:FJN

[Lan80a] D. A. Lane. Fisher, Jeffreys, and the nature of probability. In R. A. Fisher: An Appreciation [FH80], page ?? ISBN 0-387-90476-

X. LCCN QA276.16 .R18. URL http://www.loc.gov/catdir/
enhancements/fy0814/80000255-d.html;   http://www.loc.
gov/catdir/enhancements/fy0814/80000255-t.html.

Lane:1980:RAE

[Lan80b] David A. Lane. Randomization analysis of experimental data: The Fisher randomization test comment. Journal of the American Statistical Association, 75(371):587–589, September 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287651.

Lehmann:1990:MSV

[Leh90] E. L. Lehmann. Model specification: The views of Fisher and Neyman, and later developments. Statistical Science, 5(2):160– 168, May 1990. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.ss/
1177012164.

Lehmann:1993:FNP

[Leh93] E. L. Lehmann. The Fisher, Neyman–Pearson theories of testing hypotheses: One theory or two? Journal of the American Statistical Association, 88(424):1242–1249, December 1993. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2291263.

Lehmann:2011:FNC

[Leh11] E. L. (Erich Leo) Lehmann. Fisher, Neyman, and the creation of classical statistics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2011. ISBN 1-4419-9499-8 (print), 1-4419-9500-5 (e-book). viii + 115 pp. LCCN QA276.15 .L44 2011. URL http://catdir.loc.gov/

catdir/enhancements/fy1402/2011930669-b.html;   http:/
/catdir.loc.gov/catdir/enhancements/fy1402/2011930669-

d.html;   http://catdir.loc.gov/catdir/enhancements/fy1402/
2011930669-t.html;   http://www.dawsonera.com/depp/reader/
protected/external/AbstractView/S9781441995001.

Lenhard:2006:MSI

[Len06] Johannes Lenhard. Models and statistical inference: the controversy between Fisher and Neyman–Pearson. British Journal for the Philosophy of Science, 57(1):69–91, March 2006. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http://

bjps.oxfordjournals.org/content/57/1/69.full.pdf+html;
 http://www.jstor.org/stable/3541653.

Lindley:1980:RAE

[Lin80] D. V. Lindley. Randomization analysis of experimental data: The Fisher randomization test comment. Journal of the American Statistical Association, 75(371):589–590, September 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287652.

Lindley:1990:FR

[Lin90] D. V. Lindley. Fisher: a retrospective. Chance, 3(1):31–32, Winter 1990. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Ludbrook:2005:RFL

[Lud05] John Ludbrook. R. A. Fisher’s life and death in Australia, 1959– 1962. The American Statistician, 59(2):164–165, May 2005. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Lury:1972:BRBe

[Lur72] D. A. Lury. Book review: Statistical Methods for Research Workers, by R. A. Fisher. Journal of the Royal Statistical Society. Series D (The Statistician), 21(3):229, September 1972. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http:

//www.jstor.org/stable/2986695.

Mahalanobis:1938:PRA

[Mah38] P. C. Mahalanobis. Professor Ronald Aylmer Fisher. Sankhyā (Indian Journal of Statistics), 4(Part 2):265–272, December 1938. CODEN SNKYA5. ISSN 0036-4452.

Mahalanobis:1964:PRA

[Mah64a] P. C. Mahalanobis. Professor Ronald Aylmer Fisher. Biometrics, 20(2):238–252, June 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.org/

stable/2528392.

Mahalanobis:1964:SPM

[Mah64b] P. C. Mahalanobis. Some personal memories of R. A. Fisher. Biometrics, 20(2):368–371, June 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2528404.

Martin:1945:HFA

[Mar45] Laurence Martin. The hereditary and familial aspects of exophthalmic goitre and nodular goitre. Quarterly Journal of Medicine, 14(4):207–219, October 1945. CODEN QTMJAO. ISSN 1460-2393 (print), 1460-2725 (electronic).

Marriott:1980:BRBa

[Mar80] F. H. C. Marriott. Book review: Philosophical Problems of Statistical Inference. Learning from R. A. Fisher, by T. Seidenfeld. Journal of the Royal Statistical Society. Series A (General), 143 (2):195, ???? 1980. CODEN JSSAEF. ISSN 0035-9238. URL

http://www.jstor.org/stable/2981990.

Mather:1951:RFS

[Mat51] Kenneth Mather. R. A. Fisher’s statistical methods for research workers: An appreciation. Journal of the American Statistical Association, 46(253):51–54, March 1951. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

jstor.org/stable/2280093.

Mather:1964:RFW

[Mat64] K. Mather. R. A. Fisher’s work in genetics. Biometrics, 20(2): 330–342, June 1964. CODEN BIOMB6. ISSN 0006-341X (print),

1541-0420 (electronic). URL https://www.jstor.org/stable/
2528401.

Mayo:2004:KLW

[May04] Oliver Mayo. Knibbs Lecture for 2002: To what extent has Fisher’s research program been fulfilled in Australia? Australian & New Zealand Journal of Statistics, 46(4):517–529, December 2004. CO-DEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Mayo:2014:FA

[May14] Oliver Mayo. Fisher in Adelaide. Biometrics, 70(2):266–269, June 2014. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Mendel:1965:EPH

[Men65] Gregor Mendel. Experiments in Plant Hybridisation. Oliver and Boyd, Edinburgh, UK; London, UK, 1965. ix + 95 pp. LCCN QH423 .M5313. Mendel’s original paper, Versuche über Pflanzenhybride, from Verh. naturf. Ver. in Brunn, Abhandlungen, IV 1865 (appeared in 1866), in English translation, with commentary and assessment by Sir Ronald A. Fisher, together with a reprint of W. Bateson’s biographical notice of Mendel. Edited by J. H. Bennett.

Fisher:1945:HFA

[MF45] Laurence Martin and R. A. Fisher. The hereditary and familial aspects of exophthalmic goitre and nodular goitre. Note on a paper by L. Martin. Quarterly Journal of Medicine, 14(4):207–219, October 1945. CODEN QTMJAO. ISSN 1460-2393 (print), 1460-2725

(electronic). URL https://academic.oup.com/qjmed/article/
14/4/207/1569672.

Mosteller:1948:QAA

[MF48] Frederick Mosteller and R. A. Fisher. Questions and answers: Answer to question 14 on combining independent tests of significance. The American Statistician, 2(5):30–31, October 1948. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2681650.

Fisher:1951:HFA

[MF51] Laurence Martin and R. A. Fisher. The hereditary and familial aspects of toxic nodular goitre (secondary thyrotoxicosis). Quarterly Journal of Medicine, 20(2):293–312, July 1951. CODEN QTM-JAO. ISSN 1460-2393 (print), 1460-2725 (electronic).

Mooney:1995:HJM

[Moo95] Susan M. Mooney. H. J. Muller and R. A. Fisher on the evolutionary significance of sex. Journal of the History of Biology, 28(1):133–149, Spring 1995. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (electronic). URL http://link.springer.

com/article/10.1007/BF01061249.

Moore:2007:RFF

[Moo07a] James Moore. R. A. Fisher: a faith fit for eugenics. Studies in History and Philosophy of Biological and Biomedical Sciences, 38 (1):110–135, March 2007. CODEN ???? ISSN 1369-8486 (print),

1879-2499 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S136984860600094X.

Moore:2007:RAF

[Moo07b] James Moore. Ronald Aylmer Fisher (1890–1962). In Rupke [Rup07], page ?? ISBN 3-631-56803-7 (paperback). LCCN BL240.3 .E48 2007.

Morrison:2002:MPP

[Mor02] Margaret Morrison. Modelling populations: Pearson and Fisher on Mendelism and biometry. British Journal for the Philosophy of Science, 53(1):39–68, March 2002. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http://

bjps.oxfordjournals.org/content/53/1/39.full.pdf+html;
 http://www.jstor.org/stable/3541640.

Mehta:1986:AFF

[MP86a] Cyrus R. Mehta and Nitin R. Patel. Algorithm 643: FEXACT: A FORTRAN subroutine for Fisher’s exact test on unordered r × c contingency tables. ACM Transactions on Mathematical Software, 12(2):154–161, June 1986. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL http://www.acm.org/pubs/

citations/journals/toms/1986-12-2/p154-mehta/;   http:
//www.acm.org/pubs/toc/Abstracts/0098-3500/214326.html.

See remark [CaFJ93].

Menta:1986:AFF

[MP86b] Cyrus R. Menta and Nitin R. Patel. Algorithm 643: FEXACT: A FORTRAN subroutine for Fisher’s exact test on unordered r × c contingency tables. ACM Transactions on Mathematical Software, 12(2):154–161, June 1986. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Moran:1966:CRF

[MS66] P. A. P. Moran and C. A. B. Smith. Commentary on R. A. Fisher’s paper on ‘The correlation between relatives on the supposition of Mendelian inheritance’. Eugenics Laboratory Memoirs, XLI:??,

Matthen:2007:PB

[MS07] Mohan Matthen and Christopher Stephens, editors. Philosophy of Biology. Handbook of the philosophy of science. Elsevier, Amsterdam, The Netherlands, 2007. ISBN 0-444-51543-7 (hardcover), 0-08-047124-2 (e-book). xvii + 618 pp. LCCN QH331

.P45 2007. URL http://www.sciencedirect.com/science/
book/9780444515438.

Musso:2012:RBE

[Mus12] Fabio Musso. On the relation between the Eigen model and the asexual Wright–Fisher model. Bulletin of Mathematical Biology, 74(1):103–115, January 2012. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic).

URL http://link.springer.com/article/10.1007/s11538-
011-9666-0;   http://link.springer.com/content/pdf/10.
1007/s11538-011-9666-0.pdf.

Neyman:1934:TDA

[Ney34] Jerzy Neyman. On the two different aspects of the representative method: The method of stratified sampling and the method of purposive selection. Journal of the Royal Statistical Society, 97(4): 558–624, ???? 1934. ISSN 0952-8385. URL http://www.jstor. org/stable/2342192. Discussion by R. A. Fisher on pages 614– 169.

Neyman:1956:NAS

[Ney56] Jerzy Neyman. Note on an article by Sir Ronald Fisher. Journal of the Royal Statistical Society. Series B (Methodological), 18(2): 288–294, ???? 1956. CODEN JSTBAJ. ISSN 0035-9246. URL

http://www.jstor.org/stable/2983716. See comment [Fis57b].

Neyman:1928:UICa

[NP28] J. Neyman and E. S. Pearson. On the use and interpretation of certain test criteria for purposes of statistical inference: Part I. Biometrika, 20A(1/2):175–240, July 1928. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.

jstor.org/stable/2331945.

Norton:1976:NBR

[NP76] Bernard Norton and E. S. Pearson, F.R.S. A note on the background to, and refereeing of, R. A. Fisher’s 1918 paper: “On the correlation between relatives on the supposition of Mendelian inheritance” (Trans. Roy. Soc. Edinburgh 52 (1918/19), 399–434).

Notes and Records of the Royal Society of London, 31(1):151–162, July 1976. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178

(electronic). URL http://www.jstor.org/stable/531555.

ODonald:1990:FCT

[O’D90] Peter O’Donald. Fisher’s contributions to the theory of sexual selection as the basis of recent research. Theoretical Population Biology, 38(3):285–300, ???? 1990. CODEN TLPBAQ. ISSN 0040-5809 (print), 1096-0325 (electronic).

Owen:1962:ALW

[Owe62] A. R. G. Owen. An appreciation of the life and work of Sir Ronald Aylmer Fisher: F.R.S., F.S.S., Sc.D. Journal of the Royal Statistical Society. Series D (The Statistician), 12(4):313–319, ???? 1962. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/2986951.

Pires:2010:SME

[PB10] Ana M. Pires and João A. Branco. A statistical model to explain the Mendel–Fisher controversy. Statistical Science, 25(4):545–565, November 2010. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.ss/
1300108237.

Pearson:1928:THI

[PBFN28] Karl Pearson, Julia Bell, Ronald Aylmer Fisher, and Edward Nettleship, editors. Treasury of Human Inheritance. Volume 2. Anomalies and Diseases of the Eye: Nettleship Memorial Volume. Part 3: Blue Sclerotics and Fragility of Bone, volume 24 of Eugenics laboratory memoirs. Cambridge University Press, Cambridge, UK, 1928. 270–324 + xlii–xlvii pp. LCCN ????

Pearson:1931:THI

[PBFN31] Karl Pearson, Julia Bell, Ronald Aylmer Fisher, and Edward Nettleship, editors. Treasury of human inheritance. Volume 2. Anomalies and diseases of the eye: Nettleship memorial volume. Part 4. Hereditary optic atrophy (Leber’s disease), volume 26 of Eugenics laboratory memoirs. Cambridge University Press, Cambridge, UK, 1931. 326–423 + xlviii–lxiii pp. LCCN ????

Pearson:1951:BRBa

[Pea51] E. S. Pearson. Book review: Contributions to Mathematical Statistics by R. A. Fisher. Biometrika, 38(1/2):257–259, June 1951. CO-DEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic).

URL http://www.jstor.org/stable/2332332.

Pearson:1968:SHP

[Pea68] E. S. Pearson. Studies in the history of probability and statistics. XX. Some early correspondence between W. S. Gosset, R. A. Fisher and Karl Pearson, with notes and comments. Biometrika, 55(3):445–457, November 1968. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2334250.

Pearson:1970:SEC

[Pea70] E. S. Pearson. Some early correspondence between W. S. Gosset, R. A. Fisher and Karl Pearson, with notes and comments. In Pearson and Kendall [PK70], page ?? ISBN 0-85264-193-1. LCCN QA276.15 .P43.

Pearce:1975:BRB

[Pea75] S. C. Pearce. Book review: Collected Papers of R. A. Fisher. (Volume 2), by R. A. Fisher. Journal of the Royal Statistical Society. Series D (The Statistician), 24(2):151, June 1975. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2987675.

Pearce:1979:EDR

[Pea79] S. C. Pearce. Experimental design: R. A. Fisher and some modern rivals. Journal of the Royal Statistical Society. Series D (The Statistician), 28(3):153–161, September 1979. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

jstor.org/stable/2987864.

Pearce:1992:IFS

[Pea92] S. C. Pearce. Introduction to Fisher (1925) Statistical methods for research workers. In Kotz and Johnson [KJ92b], pages 59– 65. ISBN 0-387-94039-1 (New York: v. 2: softcover), 0-387-97572-1 (New York: v. 2: hardcover), 3-540-94039-1 (Berlin: v. 2: softcover), 3-540-97572-1 (Berlin: v. 2: hardcover).

LCCN QA276 .B68465 1992. URL http://www.loc.gov/catdir/
enhancements/fy0815/93003854-d.html;   http://www.loc.
gov/catdir/enhancements/fy0815/93003854-t.html.

Pearson:2001:SHP

[Pea01] E. S. Pearson. Studies in the history of probability and statistics. XX. Some early correspondence between W. S. Gosset, R. A. Fisher and Karl Pearson, with notes and comments. In Titterington and Cox [TC01], pages 317–330. ISBN 0-19-850993-6. LCCN QH301 .T55 2001. URL http://www.loc.gov/

catdir/enhancements/fy0612/2001036127-d.html;   http:/
/www.loc.gov/catdir/toc/fy031/2001036127.html. Reprinted

from Biometrika 55, 445–457 (1968).

Pearson:1994:KPR

[PFI94] Karl Pearson, R. A. Fisher, and Henry F. Inman. Karl Pearson and R. A. Fisher on statistical tests: a 1935 exchange from Nature. The American Statistician, 48(1):2–11, February 1994. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2685077.

Pearson:1925:THI

[PFSB25] Karl Pearson, Ronald Aylmer Fisher, Percy Stocks, and Amy Barrington, editors. Treasury of Human Inheritance. Volume 3. Hereditary Disorders of Bone Development. Part 1 Diaphysial Aclasis (multiple Exostoses), Multiple Enchondromata, Cleido-cranial Dystosis, volume 22 of Eugenics Laboratory memoirs. Cambridge University Press, Cambridge, UK, 1925. v + 182 + vii pp. LCCN ????

Piegorsch:1990:FCG

[Pie90] Walter W. Piegorsch. Fisher’s contributions to genetics and heredity, with special emphasis on the Gregor Mendel controversy. Biometrics, 46(4):915–924, ???? 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Pilpel:2007:SER

[Pil07] Avital Pilpel. Statistics is not enough: revisiting Ronald A. Fisher’s critique (1936) of Mendel’s experimental results (1866). Studies in History and Philosophy of Biological and Biomedical Sciences, 38(3):618–626, September 2007. CO-DEN ???? ISSN 1369-8486 (print), 1879-2499 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1369848607000416.

Pearson:1970:SHS

[PK70] E. S. (Egon Sharpe) Pearson and M. G. (Maurice George) Kendall, editors. Studies in the History of Statistics and Probability. Charles Griffin, London, UK, 1970. ISBN 0-85264-193-1. 10 + 481 pp. LCCN QA276.15 .P43.

Park:2012:EIR

[PL12] Sangun Park and Johan Lim. On the effect of imperfect ranking on the amount of Fisher information in ranked set samples. Communications in Statistics: Theory and Methods, 41(19):3608–3620, 2012. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Plackett:1989:RFH

[Pla89] Robin L. Plackett. [R. A. Fisher on the history of inverse probability]: Comment. Statistical Science, 4(3):256–258, August 1989. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic).

URL http://projecteuclid.org/euclid.ss/1177012489.

Pratt:1976:FER

[Pra76] John W. Pratt. F. Y. Edgeworth and R. A. Fisher on the efficiency of maximum likelihood estimation. Annals of Statistics, 4(3):501–514, May 1976. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176343457.

Preece:1990:RFE

[Pre90] D. A. Preece. R. A. Fisher and experimental design: a review. Biometrics, 46(4):925–935, ???? 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Provine:1992:RFS

[Pro92] William B. Provine. The R. A. Fisher–Sewall Wright controversy. In Sarkar [Sar92], chapter 8, pages 201–229. ISBN 0-7923-1777-7 (hardcover), 94-011-2856-1. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 142. URL https://link.

springer.com/chapter/10.1007/978-94-011-2856-8_8.

Quinn:2021:FR

[Qui21] Barry G. Quinn. Fisher’s g revisited. International Statistical Review = Revue Internationale de Statistique, 89(2):402–419, August 2021. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Race:1964:SNF

[Rac64] R. R. Race. Some notes on Fisher’s contributions to human blood groups. Biometrics, 20(2):361–367, June 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://

www.jstor.org/stable/2528403.

Rao:1964:SRA

[Rao64] C. Radhakrishna Rao. Sir Ronald Aylmer Fisher – the architect of multivariate analysis. Biometrics, 20(2):286–300, June 1964. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL https://www.jstor.org/stable/2528397.

Rao:1992:RFF

[Rao92] C. Radhakrishna Rao. R. A. Fisher: The founder of modern statistics. Statistical Science, 7(1):34–48, February 1992. CODEN STS-CEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.ss/1177011442.

Rao:2000:RFF

[Rao00] C. R. Rao. R. A. Fisher: the founder of modern statistics. In Rao and Székely [RS00], page ?? ISBN 0-8247-9029-4 (hardcover).

LCCN QA276.16 .S844 2000. URL http://www.loc.gov/catdir/
enhancements/fy0647/99057753-d.html.

Rao:2007:LWR

[Rao07] C. R. Rao. Life and work of Ronald Aylmer Fisher. Journal of Statistical Theory and Practice, 1(3–4):489–499, September 2007. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic).

URL http://link.springer.com/article/10.1080/15598608.
2007.10411854.

Rao:2008:LWR

[Rao08] C. R. Rao. Life and work of Ronald Aylmer Fisher. Journal of Statistical Theory and Practice, 2(1):131–141, March 2008. CODEN ???? ISSN 1559-8608 (print), 1559-8616 (electronic).

URL http://link.springer.com/article/10.1080/15598608.
2008.10411866.

Rosales:2017:TNW

[Ros17] Alirio Rosales. Theories that narrate the world: Ronald A. Fisher’s mass selection and Sewall Wright’s shifting balance. Studies in History and Philosophy of Science Part A, 62(??):22–30, April 2017. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0039368117300742.

Rao:2000:SCM

[RS00] C. Radhakrishna (Calyampudi Radhakrishna) Rao and Gábor J. Székely, editors. Statistics for the 21st Century: Methodologies for Applications of the Future [8th Eugence Lukacs Symposium, 1998, Bowling Green State University, Bowling Green, OH, USA], volume 161 of Statistics, textbooks and monographs. Marcel Dekker, Inc., New York, NY, USA, 2000. ISBN 0-8247-9029-4 (hardcover).

LCCN QA276.16 .S844 2000. URL http://www.loc.gov/catdir/
enhancements/fy0647/99057753-d.html.

Rubin:1980:RAE

[Rub80] Donald B. Rubin. Randomization analysis of experimental data: The Fisher randomization test comment. Journal of the American Statistical Association, 75(371):591–593, September 1980. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2287653.

Rupke:2007:ELT

[Rup07] Nicolaas A. Rupke, editor. Eminent Lives in Twentieth-century Science and Religion. Peter Lang, Frankfurt am Main, Germany, 2007. ISBN 3-631-56803-7 (paperback). xiv + 255 pp. LCCN BL240.3 .E48 2007.

Ryan:1980:HDR

[Rya80] K. C. Ryan. Historical development of R. A. Fisher’s fiducial argument. Ph.D. dissertation, University of London, Chelsea College, London, UK, 1980. ?? pp. URL http://catalogue.

libraries.london.ac.uk/record=b3010948~S24;   https://
search.proquest.com/pqdtglobal/docview/301325697.

Salsburg:2001:LTT

[Sal01] David Salsburg. The Lady Tasting Tea: How Statistics Revolutionized Science in the Twentieth Century. W. H. Freeman, New York, NY, USA, 2001. ISBN 0-7167-4106-7 (hardcover), 0-8050-7134-2 (paperback). xi + 340 pp. LCCN Q175 .S2345 2001. URL http:/

/catdir.loc.gov/catdir/enhancements/fy0667/00049523-b.
html;   http://catdir.loc.gov/catdir/enhancements/fy0667/
00049523-d.html.

Sarkar:1992:FEG

[Sar92] Sahotra Sarkar, editor. The Founders of Evolutionary Genetics: a Centenary Reappraisal, volume 142 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1992. ISBN 0-7923-1777-7 (hardcover), 94-011-2856-1. ISSN 0068-0346 (print), 2214-7942 (electronic). 300 pp.

LCCN Q174 .B67 vol. 142. URL http://www.loc.gov/catdir/
enhancements/fy0823/92012823-d.html;   http://www.loc.
gov/catdir/enhancements/fy0823/92012823-t.html;   https:
//link.springer.com/book/10.1007/978-94-011-2856-8.

Sarkar:1995:JBH

[Sar95] Sahotra Sarkar. J. B. S. Haldane and R. A. Fisher’s draft life of Karl Pearson. Notes and Records of the Royal Society of London, 49(1):119–124, July 1995. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.jstor.

org/stable/531888.

Savage:1976:RRF

[Sav76] Leonard J. Savage. On rereading R. A. Fisher. Annals of Statistics, 4(3):441–500, May 1976. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176343456.

Seidenfeld:1992:RFD

[Sei92a] T. Seidenfeld. R. A. Fisher on the design of experiments and statistical estimation. In Sarkar [Sar92], chapter 2, pages 23–36. ISBN 0-7923-1777-7 (hardcover), 94-011-2856-1. ISSN 0068-0346 (print), 2214-7942 (electronic). LCCN Q174 .B67 vol. 142. URL https://

link.springer.com/chapter/10.1007/978-94-011-2856-8_2;
 https://www.cmu.edu/dietrich/philosophy/docs/seidenfeld/
Fisher%20on%20Design.pdf.

Seidenfeld:1992:RFF

[Sei92b] Teddy Seidenfeld. R. A. Fisher’s fiducial argument and Bayes’ Theorem. Statistical Science, 7(3):358–368, August 1992. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1177011232.

Senn:1991:BRS

[Sen91] Stephen Senn. Book review: Statistical inference and analysis: Selected Correspondence of R. A. Fisher. J. H. Bennett (ed), Oxford University Press, Oxford, 1990. No. of pages: xviii + 380. Price: £40.00. Statistics in Medicine, 10(8):1326, August 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Skipper:2000:RFS

[Ski00] Robert Alan Skipper, Jr. The R. A. Fisher–Sewall Wright controversy in philosophical focus: Theory evaluation in population genetics. Ph.D. dissertation, University of Maryland, College Park, MD, USA, 2000. viii + 397 pp. URL https://search.proquest.

com/pqdtglobal/docview/304605614.

Skipper:2007:SRA

[Ski07] Robert A. Skipper, Jr. Sir Ronald Aylmer Fisher. In Matthen and Stephens [MS07], pages 37–48. ISBN 0-444-51543-7 (hardcover), 0-08-047124-2 (e-book). LCCN QH331 .P45 2007. URL http://

www.sciencedirect.com/science/book/9780444515438.

Smith:1966:BRBb

[Smi66] Cedric A. B. Smith. Book review: Experiments in Plant Hybridisation, by Gregor Mendel; Ronald A. Fisher. Journal of the Royal Statistical Society. Series A (General), 129(1):141–142, ???? 1966. CODEN JSSAEF. ISSN 0035-9238. URL http://www.jstor.

org/stable/2343925.

Speed:1992:IFA

[Spe92] T. P. Speed. Introduction to Fisher (1926) The arrangement of field experiments. In Kotz and Johnson [KJ92b], pages 71– 81. ISBN 0-387-94039-1 (New York: v. 2: softcover), 0-387-97572-1 (New York: v. 2: hardcover), 3-540-94039-1 (Berlin: v. 2: softcover), 3-540-97572-1 (Berlin: v. 2: hardcover).

LCCN QA276 .B68465 1992. URL http://www.loc.gov/catdir/
enhancements/fy0815/93003854-d.html;   http://www.loc.
gov/catdir/enhancements/fy0815/93003854-t.html.

Spicer:1959:BRB

[Spi59] C. C. Spicer. Book review: Smoking: The Cancer Controversy. Some Attempts to Assess the Evidence, by Ronald A. Fisher. Journal of the Royal Statistical Society. Series A (General), 122(4): 554–556, ???? 1959. CODEN JSSAEF. ISSN 0035-9238. URL

http://www.jstor.org/stable/2343092.

Sabbaghi:2014:CNF

[SR14] Arman Sabbaghi and Donald B. Rubin. Comments on the Neyman–Fisher controversy and its consequences. Statistical Science, 29(2):267–284, May 2014. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1408368581.

Stigler:1973:SHP

[Sti73] Stephen M. Stigler. Studies in the history of probability and statistics. XXXII. Laplace, Fisher, and the discovery of the concept of sufficiency. Biometrika, 60(3):439–445, December 1973. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2334992.

Stigler:1976:CDR

[Sti76] Stephen M. Stigler. Contribution to discussion of “On rereading R. A. Fisher. Annals of Statistics, 4(3):498–500, May 1976. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176343456.

Stigler:1999:FSS

[Sti99a] Stephen M. Stigler. The foundations of statistics at Stanford. The American Statistician, 53(3):263–266, August 1999. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//www.jstor.org/stable/2686107.

Stigler:1999:STH

[Sti99b] Stephen M. Stigler. Statistics on the Table: the History of Statistical Concepts and Methods. Harvard University Press, Cambridge, MA, USA, 1999. ISBN 0-674-83601-4. ix + 488 pp. LCCN QA276.15 .S755 1999.

Stigler:2005:F

[Sti05] Stephen Stigler. Fisher in 1921. Statistical Science, 20(1):32–49, February 2005. CODEN STSCEP. ISSN 0883-4237 (print), 2168-

8745 (electronic). URL http://projecteuclid.org/euclid.ss/
1118065041.

Stigler:2006:HRF

[Sti06] Stephen M. Stigler. How Ronald Fisher became a mathematical statistician. Mathématiques et Sciences Humaines. Mathematics and Social Sciences, 44(176):23–30, Hiver 2006. ISSN 0987-6936 (print), 1950-6821 (electronic). URL https://journals.

openedition.org/msh/3631.

Street:1990:FCA

[Str90] Deborah J. Street. Fisher’s contributions to agricultural statistics. Biometrics, 46(4):937–945, ???? 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Stute:1989:HSZ

[Stu89] Winfried Stute. Der historische Streit zwischen R. A. Fisher und J. Neyman oder: Ein Sittengemälde aus der Blütezeit der englischen Schule für Statistik. (German) [The historical controversy between R. A. Fisher and J. Neyman or: a portrayal of customs belonging to the blossom time of the English school of statistics]. Mathematische Semesterberichte, 36(1):61–84, 1989. ISSN 0720-728X (print), 1432-1815 (electronic).

Stute:1992:HCB

[Stu92] W. Stute. History of controversies between R. A. Fisher and J. Neyman or a picture of manners in time of the rise of the English school of statistics. Ann. Soc. Math. Pol. Ser. II Wiad. Mat., 29 (2):205–221, ???? 1992.

Soper:1917:DCC

[SYC^{+}17] H. E. Soper, A. W. Young, B. M. Cave, A. Lee, and K. Pearson. On the distribution of the correlation coefficient in small samples. Appendix II to the papers of “Student” and R. A. Fisher. Biometrika, 11(4):328–413, May 1917. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/

stable/2331830.

Tabery:2008:RFL

[Tab08] James Tabery. R. A. Fisher, Lancelot Hogben, and the origin(s) of genotype–environment interaction. Journal of the History of Biology, 41(4):717–761, December 2008. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (electronic).

URL http://link.springer.com/article/10.1007/s10739-
008-9155-y;   http://link.springer.com/content/pdf/10.
1007/s10739-008-9155-y.pdf.

Titterington:2001:BOH

[TC01] D. M. Titterington and D. R. (David Roxbee) Cox, editors. Biometrika: one hundred years. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2001. ISBN 0-19-850993-6. viii + 383 pp. LCCN QH301 .T55 2001. URL http://www.loc.gov/

catdir/enhancements/fy0612/2001036127-d.html;   http:/
/www.loc.gov/catdir/toc/fy031/2001036127.html.

Thornton:1927:EDC

[TF27] H. G. Thornton and R. A. Fisher. On the existence of daily changes in the bacterial numbers in an American soil. Soil Science, 23(4):

253–259, April 1927. URL https://journals.lww.com/soilsci/
Citation/1927/04000/ON_THE_EXISTENCE_OF_DAILY_CHANGES_
IN_THE_BACTERIAL.1.aspx.

Thornton:1934:NBC

[TG34] Henry Gerard Thornton and P. H. H. Gray. The numbers of bacterial cells in field soils, as estimated by the ratio method. Proceedings of the Royal Society of London. Series B. Biological sciences, 115(795):522–543, August 1, 1934. CODEN PRLBA4. ISSN 0962-8452 (print), 1471-2954 (electronic). With an Appendix by Ronald A. Fisher.

Thompson:1990:RFC

[Tho90] E. A. Thompson. R. A. Fisher’s contributions to genetical statistics. Biometrics, 46(4):905–914, ???? 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Tippett:1953:BRS

[Tip53] L. H. C. Tippett. Book reviews: Statistical Tables for Biological, Agricultural and Medical Research, by Ronald A. Fisher and Frank Yates. Applied Statistics, 2(3):203, November 1953. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Tocher:1958:BRBa

[Toc58] K. D. Tocher. Book review: Statistical Tables for Biological, Agricultural and Medical Research, by R. A. Fisher; F. Yates. Journal of the Royal Statistical Society. Series A (General), 121(2):240, ???? 1958. CODEN JSSAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343368.

Tabery:2015:RFL

[TS15] James Tabery and Sahotra Sarkar. R. A. Fisher, Lancelot Hogben, and the ‘competition’ for the Chair of Social Biology at the London School of Economics in 1930: Correcting the legend. Notes and Records of the Royal Society of London, 69(4):437–446, December 20, 2015. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Tsumoto:2002:STR

[Tsu02] Shusaku Tsumoto. Statistical test for rough set approximation based on Fisher’s exact test. Lecture Notes in Computer Science, 2475:381–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/series/0558/bibs/2475/
24750381.htm;   http://link.springer.de/link/service/series/
0558/papers/2475/24750381.pdf.

Tukey:1952:RSMa

[Tuk52] John W. Tukey. Review of Statistical Methods for Research Workers, 11th ed, by R. A. Fisher. Econometrica, 20(??):511–512, ???? 1952. CODEN ECMTA7. ISSN 0012-9682.

Turner:1987:RGD

[Tur87] J. R. G. Turner. Random genetic drift, R. A. Fisher, and the Oxford School of ecological genetics. In Krüger et al. [KDH87], page ???? ISBN 0-262-11118-7 (hardcover), 0-262-61062-0 (paperback), 0-262-11125-X (set). LCCN QA273.A4 P76 1987.

Ullrich:1999:ESB

[Ull99] Peter Ullrich. An Eulerian square before Euler and an experimental design before R. A. Fisher: On the early history of latin squares. Chance, 12(1):22–27, 1999. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

VonPlato:1994:CMP

[Von94] Jan Von Plato. Creating Modern Probability: Its Mathematics, Physics, and Philosophy in Historical Perspective. Cambridge studies in probability, induction, and decision theory. Cambridge University Press, Cambridge, UK, 1994. ISBN 0-521-44403-9 (hardcover). x + 323 pp. LCCN QA273 .V66 1994. URL http:

//www.loc.gov/catdir/description/cam025/93019026.html;
 http://www.loc.gov/catdir/toc/cam025/93019026.html.

Welch:1956:NSC

[Wel56] B. L. Welch. Note on some criticisms made by Sir Ronald Fisher. Journal of the Royal Statistical Society. Series B (Methodological), 18(2):297–302, ???? 1956. CODEN JSTBAJ. ISSN 0035-9246.

URL http://www.jstor.org/stable/2983718. See [Fis57b].

Wetherill:1965:BRBb

[Wet65] G. B. Wetherill. Book review: Statistical Tables for Biological, Agricultural and Medical Research, by R. A. Fisher; F. Yates. Journal of the Royal Statistical Society. Series A (General), 128(1):146, ???? 1965. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343448.

Williams:1964:SEB

[Wil64] C. B. Williams. Some experiences of a biologist with R. A. Fisher and statistics. Biometrics, 20(2):301–306, June 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL

https://www.jstor.org/stable/2528398.

Wishart:1927:DEI

[Wis27] R. A. Fisher J. Wishart. On the distribution of the error of an interpolated value, and on the construction of tables. Proceedings of the Cambridge Philosophical Society. Mathematical and physical sciences, 23(8):912–921, October 1927. CODEN PCPSA4. ISSN

0008-1981. URL https://www.cambridge.org/core/journals/
mathematical-proceedings-of-the-cambridge-philosophical-
society/article/on-the-distribution-of-the-error-of-an-
interpolated-value-and-on-the-construction-of-tables/9D9AE42324CAE8BB685A589EB1443CAF#

.

Watson:1936:DPS

[WTRS^{+}36] D. M. S. Watson, N. W. Timofeeff-Ressovsky, E. J. Salisbury, W. B. Turrill, T. J. Jenkin, R. R. Gates, R. A. Fisher, C. Diver, G. D. H. Carpenter, J. B. S. Haldane, E. W. MacBride, and R. N. Salaman. A discussion on the present state of the theory of natural selection. Proceedings of the Royal Society of London. Series B. Biological sciences, 121(820):43–73, August 1, 1936. CODEN PRLBA4. ISSN 0962-8452 (print), 1471-2954 (elec-

tronic). URL https://royalsocietypublishing.org/doi/10.
1098/rspb.1936.0052.

Y:1939:BRBa

[Y.39] F. Y. Book review: Statistical Methods for Research Workers, by R. A. Fisher. Journal of the Royal Statistical Society, 102(2): 298, ???? 1939. ISSN 0952-8385. URL http://www.jstor.org/

stable/2980010.

Yates:1964:FPR

[Yat64a] F. Yates. Fiducial probability, recognisable sub-sets and Behrens’ test. Biometrics, 20(2):343–360, June 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://

www.jstor.org/stable/2528402.

Yates:1964:SRF

[Yat64b] F. Yates. Sir Ronald Fisher and the design of experiments. Biometrics, 20(2):307–321, June 1964. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL https://www.jstor.

org/stable/2528399.

Yates:1979:BRB

[Yat79] F. Yates. Book review: R. A. Fisher. The Life of a Scientist, by Joan Fisher Box. Journal of the Royal Statistical Society. Series A (General), 142(4):504–506, ???? 1979. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.org/stable/2982557.

Yates:1982:BRB

[Yat82] F. Yates. Book review: R. A. Fisher: An Appreciation, by R. A. Fisher; S. E. Fienberg; D. V. Hinckley. Journal of the Royal Statistical Society. Series A (General), 145(3):368–370, ???? 1982. CODEN JSSAEF. ISSN 0035-9238. URL http://www.jstor.

org/stable/2981872.

Yates:1963:RAF

[YM63] Frank Yates and Kenneth Mather. Ronald Aylmer Fisher 1890–1962. Biographical Memoirs of Fellows of the Royal Society, 9(??):91–129, November 1, 1963. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https://

royalsocietypublishing.org/doi/10.1098/rsbm.1963.0006;
 https://www.jstor.org/stable/769423.

Youden:1951:FRM

[You51] W. J. Youden. The Fisherian revolution in methods of experimentation. Journal of the American Statistical Association, 46 (253):47–50, March 1951. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2280092.

Youden:1962:MSR

[You62] W. J. Youden. Memorial to Sir Ronald Aylmer Fisher: 1890– 1962. Journal of the American Statistical Association, 57(300):

727–728, December 1962. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274x (electronic). URL http://links.jstor.org/

sici?sici=0162-1459(196212)57:300<727:MTSRAF>2.0.CO;2-
P&origin=MSN;   http://www.jstor.org/stable/2281804.

Zabell:1989:RFHa

[Zab89a] Sandy Zabell. R. A. Fisher on the history of inverse probability. Statistical Science, 4(3):247–256, August 1989. CODEN STS-CEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http:

//projecteuclid.org/euclid.ss/1177012488.

Zabell:1989:RFHb

[Zab89b] Sandy Zabell. [R. A. Fisher on the history of inverse probability]: Rejoinder. Statistical Science, 4(3):261–263, August 1989. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL

http://projecteuclid.org/euclid.ss/1177012491.

Zabell:1992:RFF

[Zab92] S. L. Zabell. R. A. Fisher and fiducial argument. Statistical Science, 7(3):369–387, August 1992. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://

projecteuclid.org/euclid.ss/1177011233.

Ziliak:2008:CSS

[ZM08] Stephen Thomas Ziliak and Deirdre N. McCloskey. The Cult of Statistical Significance: How the Standard Error Costs Us Jobs, Justice, and Lives. Economics, cognition, and society. University of Michigan Press, Ann Arbor, MI, USA, 2008. ISBN 0-472-07007-X (cloth), 0-472-05007-9 (paperback). xxiii + 321 pp. LCCN HB137

.Z55 2008. URL http://www.loc.gov/catdir/enhancements/
fy0809/2007035401-d.html.